A Complete Bibliography of Publications in the Journal of Cell Biology: 1955–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 June 2019 Version 1.00

Title word cross-reference

6 [Bie55]. ^{14} [BSB59, Koe58b, TS59]. ^{35} [BVLC57, Dzi56, DBDA57, EU57, JC57]. ^{45} [BVLC57, Joh58]. _{4} [Cau57, Gal59b, LDK55, Rud55]. α [Car57]. β [Str56]. d [Car57]. n [Car57].

-acetyl- [Car57]. -chitin [Car57]. -Collidine [BL59]. -glucuronidase [Str56]. -Glyceraldehyde-3-Phosphate [NS59]. -Mercaptopurine [Bie55]. -Sulfate [Dzi56, JC57].

45 [SB59b].

5 [ENM58, SD58]. 5-Hydroxyuracil [SD58]. 5-Nucleotidase [ENM58].

6-Diaminopurine [SD55].

Aberrations [Kih56, Kih57, KMS57, Kih59a]. Ability [DM55b]. Abnormal [Bin59]. abramus [Yam58]. Absence [WZ59, ZW59]. Absorptive [Cla59]. Accompanying [LVH58]. According [BS56, SB56]. Accumulation [FS59, KSB59]. Acetabularia [SP58f]. acetyl [Car57]. Acetylcholinesterase [FK59]. Acid [BG55b, Gri56, JK56, Koe58b, LSW^{+}57, LR56b, MM55, MKT59, Str56, BO59, BS59, CvTF58, DO58, Gut57, HL59, Koe58a, LS55, LL55, LBVS59, MKT58, Pau56, PR57a, Pre59, RB58, Sie55, dLAC59]. Acid-Soluble [MM55]. Acids [FSF56, Woo58, Ber56]. Acinar [Mun58]. Acoustic [De 56d]. Acrididae [BTD55]. Acroblasts [BTDA56]. Acth [FSF56]. Actin [SB56]. Actin-Rich [SB56]. Action [DOMtTAoJCWW57, FSF56, SD55]. Active [Ben56a]. Activities [ENM58, NS59, WPE^{+}58]. Activity [BP57, BR56a, Gri56, OP57, BP59a, FH59, Qui58, SD58, SM58, SNM58, WMK59b]. Acylamido [BS58a]. Added [Hol59]. Adenine [Koe58b]. Adenine-8-C [Koe58b]. Adenocarcinoma [Faw56a]. Adenosinetriphosphatase [ENM58, SNM58]. Adenovirus [BMG^{+}57, Eps59, GM59c, MGR^{+}57]. Adenoviruses [LB57]. Adepidermal [SS59]. Adhesion [Bac55]. Adipose [BM58b, LC58, NF58]. Adrenal [DF57, FSF56, LH57, Bel58, RPLL58, SM58]. Adrenocortex [DS58]. Adrenocortical [Lev56a]. Adult [Koe58b, Rob55, BMC59]. Affecting [Kih56]. After [BVLC57, De 56d, EU57, FD56, HHP57, MRC56, Str57a, AvB58, DF57, Gal59b, WPE^{+}58]. Afterload [KC56]. Agar [Wen59b]. Age [PH58]. Agent [LMM^{+}59, MLM^{+}59]. Agents [Kih59a]. Aggregates [BMG^{+}57]. Aging [TC59]. Aid [NM56, WK56, WM59]. Albino [Wei57, WL55, Lad58]. Albumin [HDG57]. Albumin-Secreting [HDG57]. Aldolase [NS59]. Alkali [DM55b]. Alkaline [ED55a, ED55b, BB59]. Allium [Laf58a, Laf58b]. Allomyces [RM57]. Alpha [BS58a]. Alpharadiography [Bél59]. Alterations [Gut57, LR56b, Rob58a]. Alveolar [Kar56c, Kar58]. Amblystoma [Por56, Hay58]. Ameba [AR59, Coh56, AR58, PB58]. Ameboid [AR59, AR58]. americana [Hes58]. Amino [HL59]. Amoeba [Coh57b, GKF56, Pap56a, Pap56b, PR57a, PS58, Pre59]. among [LB58]. Amounts [Woo58]. Amphibian [Gal56, WF56, Wis57]. Anaerobic [Mer59c]. Analysis [HMF57, KHD56, Str58, TB56, WS56b, SD58]. Anastomoses [Tha59]. Angle [FmF57]. Animal [JK57, MM57, MPM56, Wat59]. Animals [YI57, YT58]. Annelida [CCP57]. Annulate [Swi56, Mer59a]. Antagonistic [Bie55]. Anterior [LB58, AR58, LB59]. Anthers [FS59, Ste58]. Anti [RWPP55]. Anti-Collagen [RWPP55]. Antibodies [HNS59]. Antibody [Wen59b, Lat59]. Antigen [Wen59b]. Antimetabolite [SD55, SD58]. Antimyosin [HMF57]. Aoyama [LC56]. Apc [BMG^{+}57, LP56]. Apoferritin [Ric59]. Apparatus [CS56, LC56, Lac57, Wen59b, FW57, Wen59a]. Appearing [TI55]. Applications [FM56]. aquilinum [Man59]. Araldite [GG58]. Arbacia [MM55]. Architecture [CER57, Kar56c]. Area [ET58]. Arenarum [BF56]. Areticular [Smi59]. Arginase [WPE^{+}58]. Arrangement [SB58]. Array [Hux57]. Arteries [MR57b, Buc58]. Artifacts [TI55, Wat57]. Ascites [AP57, LOK56, FERS59, KSB59, Lov57]. Ascitic [Eps57b]. Ascorbic [JK56]. Ase [DOMtTAoJCWW57, WWC55]. Aspects [DBDA57, Rot56, SBF58, BBG59a, BBG59b]. Assays [BSB59]. Associated [Faw56a, Bin59]. Association [CD59]. Asterias [KJTD55]. ATP [WWC55]. ATP-Ase [WWC55]. ATPase [WM58]. Atrium [FS58a]. Attachment [Woo59]. aurelia [Dip58, HG58]. Autoradiograms [Dzi56]. Autoradiographic [BSB59, EU57, JC57, Joh58, RV59, HL59, Koe58a, TC59]. Autoradiographs [SB59b]. Autoradiography [BVLC57]. Avena [BCCMS57]. Avian [PR57b]. Avoiding [DM59]. Axial [Mos58]. Axoplasm [Koe55, TD56, Via58].

B [Mat56]. Bacillus [Han57a, FJ58, FJY58, TY59a, TY59b]. Bacterial [YFJ59a, YFJ59b, YFJ59c, Cha59]. bactéries [SRK59]. Bacteriophage [Bac55, SRK59]. Bacterium [WZ59, ZW59]. Band [BP59a]. Barrier [DW55a, WL55, vBC55]. Basal [Pea56, SS59]. Base [RMW57]. Basement [WF56, vBRC56]. Basic [RW59]. Basis [BL59]. Basophile [Fin58]. Basophilic [Reb56a, Reb56b, Rut58]. Bat [Yam58]. Be [Cos57, Orn56a]. Beam [MMR56]. Bearing [AR59]. Beef [LB59]. Behavior [Gro56, Cha59]. Behrens [ED55b]. berghei [RT59]. Between [BR56a, CD59, KSB59, Mos58, Tha59]. Bicoecidae [Rob56b]. Biochemical [BB57, HMM57, KHD56, LR56b, LVH58]. Biological [FM56, SM58, Ste57, Mer58, MLM^{+}59]. Biopsy [HSH^{+}59]. Biosynthesis [JS57, SJ57, WG57]. Bipolar [Lad58]. Birefringence [Tho57]. Bladder [Yam55a]. Blastema [Hay58]. Blood [DW55a, WL55, Kar59, TG58]. Blood-Brain [DW55a]. Blood-Ocular [WL55]. Blowfly [Han56]. Blue [HM59]. Bodies [MBMR55, FJY58, NH57, Smi59]. Body [AR57, Han57a, Han57b, STC57, Yam59, Ban59, Ros59, Yam58]. Bombix [RV59]. Bone [RC56, SR57]. Bones [BVLC57]. Border [SBF58]. Born [BM58a]. Both [Lit59]. Bound [BS58a]. Bovine [SM58]. Bracken [Man59]. Brain [DW55a]. Breaking [Cam56]. Breast [How56]. Brei [And56]. Bridges [FIS59, Tha59]. Bromphenol [HM59]. Brown [LC58, NF58]. Brush [SBF58]. Buds [dL58a]. Buffered [LDK55]. Buffering [BL59]. Bufo [BF56]. Bull [LS55].

C [BSB59, Koe58b, TS59]. Ca [BVLC57, Joh58]. Calcium [DBDA57, SB59b]. Calcium-45 [DBDA57, SB59b]. Calf [HL58b]. Calliphora [Han56]. Capable [WLP59]. Capacitance [Rot57c]. Capacity [DOMtTAoJCWW57]. Capillaries [MR57b, WM59]. Capillary [Kar56c]. Capsular [KP55]. Capsule [HSH^{+}59]. Carbon [Wat56, DW58, Spe59, SP58g, Wat55b, dH58]. Carboxyl [BS58a]. Carcinoma [LMM^{+}59]. Cardiac [MR57a]. Carotid [Ros59]. Cartilage [RC56, Hay58, SR58]. Cat [DW56, BP59b, Ros59]. cathepsin [Str56]. Cavity [Odo56]. Cecum [KJTD55]. Cell [BCCMS57, CH58, Coh57a, DM55b, DWM^{+}55, DOMtTAoJCWW57, Gri56, KR58, LB58, LB59, MLKD56, Nao57, Nov56, Sac55, SL57, dHF58, Bad59, BM58b, Gre58, Lad58, Lov57, Mun58, Nov59, RW59, Rot57b, SNM58, Wat59, WPE^{+}58, WLP59]. Cell-Free [dHF58]. Cellophane [RPST58]. Cells [AP57, BTD55, BTDA56, Bél59, BR56a, BG55a, BMG^{+}57, DM56, DM57, ED55a, Eps57b, Eps57a, Faw56a, FD56, GM59c, HHPvR56, HHP57, HDG57, Hes58, HMM56, HMM57, Hol59, JK56, KGF56, Koe58b, Lac57, Mol59, MR57a, Pal56b, Por56, RR59, Ros57, RB56, RMW57, Sid57, Str57a, Sul55, TB56, Str56, AvB58, Cla59, CD59, DW58, DM55a, ENM58, FI58, FIS59, FERS59, GM59a, GT59, Hay58, HA58, KWDV58, KSB59, Laf58b, LD59, LB57, MQ59, Mil58, MGR^{+}57, NH57, Obl58, SB58, SNM58, Str57b, Tha59, TG58, TCS59, WMK59a, Yam58, dLAC59]. Cellular [BR56a, Cla57, DGW58, BV59, Cha59, KAWI59, Lat59]. Cellulose [SB58]. Central [Har56, Lus56a, Lus56b, Pal56c, vBC55, DGW58]. Centrifugal [KHD56]. Centrifugation [DM59, HL58b, SM58]. cepa [Laf58b, Laf58a]. Cephalopods [Wol58]. Cerebellum [BM58a]. Cerebral [Gra59, PG59]. Certain [Lev56b, WPE^{+}58]. Chaetopterus [Mer59b]. Chain [Coh55, Hes58]. Chambers [RPST58]. Chang [Hol59]. Change [Gri56, BMC59]. Changes [BO59, De 56d, DF57, Lev56a, NS59, Reb57, Str57a, TPD56, DS58, KWDV58, LD59, LVH58, RSK59, Via58]. Characteristics [Gre58, Mer59b]. Chelating [Kih59a]. Chemical [Alf56, Coh57a, HK56, HL56, KJTD55, Koe55, LS57, YFJ59a, YFJ59b, YFJ59c, FJ58, FJY58, HL57, HL58b, Pau56, SJS59]. Chemically [Hol59, Bin59]. Chemicals [Kih56, Tay59]. Chick [JC57, WK56, Joh58, PP59]. Chicken [Bin59]. chitin [Car57]. Chlamydomonas [SP57]. Chlorophyll [FH59]. Chloroplast [SP57]. Chloroplasts [HMM55]. Cholesterol [Fei56]. Cholinesterase [LO59]. Choroid [Cas59, MP56]. Chromatin [MKT58]. Chromatographic [Woo57]. Chromium [LF56]. Chromosomal [MKT59, Mos56a, Kih59a]. Chromosome [Gay56, Kih56, Kih57, KMS57, Mos58, WARL58]. Chromosome-Nuclear [Gay56]. Chromosomes [De 56a, Faw56b, KD56, Mos58, Ris56, LR58, Mer59c]. Cicada [ERdH58b]. Cilia [Mil58]. Ciliary [Rot56]. Ciliated [Gal59a]. Cinematographic [Ros57]. cinereus [Mos58]. Class [GMC55]. Clear [CH58]. Cleavage [KD58]. Cleaving [JK57]. Clefts [Rob58b]. Close [BR56a]. Cloudy [FS56]. Cnidoblasts [SF59]. Cockroach [Hes58, MB59]. coelicolor [GH59]. Coenzyme [Bie55, CZ59a]. Coleoptiles [BCCMS57]. coli [Bac55, CvTF58, KR58, WZ59, ZW59]. Collagen [Coh55, GMC55, Gro56, JS57, PP59, RWPP55, SR57, SJ57, WG57, Bea56]. Collagenous [WL56]. Colleterial [MB59]. Collidine [BL59]. Colloidal [HHP57, Cla59]. Colorimetric [Str58]. Colorless [Rob56b]. Columnar [Cla59]. Comb [BSB59]. Common [Lac57]. Comparative [DF56a, ERdSSVF56]. Comparison [FD56, HK56, HL58b, MR58]. Compensatory [BO59]. complanata [BM58b]. Complex [BG55a, DF56a, Mos58, Lit59]. Component [GH59, Laf58a, Pal55a]. Components [Gré57, HMM57, MB55, SP55, Tho57, TS59, HN59]. Composition [Koe55, LR56b, HL58b, dLAC59]. Compound [GP57, Mil57]. Compounds [FS59, LF56]. Concentration [Bél59, Gri56, Sid57, Str56, Ste58]. Concept [Bad59]. Concepts [Ben56a]. Concerning [WL56]. Cones [DF56b, DL58b]. Configuration [Coh55]. Conjugating [ET58]. Connective [Orn56c, WK56, WLP59]. Considérations [SRK59]. Consistency [AR59]. Constituents [HL59]. Contact [ET58]. Contagiosum [Ban59]. Containing [Rob56b, Tho57, FH59, KRS58]. Content [LR56a, GL59]. Continuous [WLP59]. Contractile [PB58, Rud58, STC59]. Contraction [KC56]. Contrast [Ros57, FI58]. Contribution [Mer58]. Control [WMK59b]. Convoluted [RMW57]. Cornea [Jak56, She56]. Corneal [HL59]. Cornified [MO55, MB55]. Corpuscle [PQ57, Qui58]. Correlated [BMG^{+}57, Mos56b]. Correlation [LH57, MLM^{+}59]. Cortex [FSF56, Gra59, PG59, SM58]. Cortical [Kem56a, Pit56, SP55, Bel58]. Corticotropin [FSF56]. Cortisone [LMS55, LR56b, LR56a, LBVS59]. coupes [SRK59]. Coupling [LO59]. Coverslip [NM56]. Cow [LH57]. Crab [HC58]. crassa [ST59]. Crayfish [Mos56a, Rut58, Rut59]. Cricket [BTDA56]. Crosby [HSH^{+}59]. Cross [Han56, Hux57]. Cross-Striated [Hux57]. Cross-Striation [Han56]. cruentum [BV59]. Crystal [RR59, SR57, Ste56, Car57]. Crystal-Collagen [SR57]. Crystalline [BMG^{+}57, Yam59, LP56, MGR^{+}57, Ric58]. Crystals [LB57, Str59b]. Cultivated [MM57, MPM56]. Cultivation [Hol59, WPE^{+}58]. Culture [Bie55, GMGL57, RWPP55, DW58, GM59a, Lat59, RPST58, WLP59]. Cultured [DM55a, DO58, HA58, KWDV58]. Cultures [Ros57, BM58a, Fri59, LMM^{+}59, MQ59, Pau56]. Culturing [RPST58]. Cupferron [Kih59b]. Cyanide [Kih59b]. Cycle [Gri56, KWDV58, NS59]. Cytochemical [CS56, JK56, LOK56, Mos56b, PCV58, SW56, WS56a, Bad59, HSP58, KD58, SHP58, SP58c, SP58d, SP58e, SP59a, Str59a]. Cytochemistry [Mon55, Mon57]. Cytological [CT59a, CT59b, FJ58, HDG57, HD58, Lit59, LOK56, SD55, LDK55, FK59, LVH58, SD58]. Cytology [Mat58, Rot57a]. Cytophotometric [LBVS59]. Cytophysiology [ERdSSVF56]. Cytoplasm [AR59, GH59, HHPvR56, Kem56a, Pal55a, PS58]. Cytoplasmic [Cle56, Gay56, GM59a, Gri59, HMM57, KHD56, PR57a, Bin59, Cha59, FI58, KR58, PH55, PH58, Pre59, SB59a, dLAC59].

D [NS59]. dans [BBG59a, BBG59b]. Defects [Rob58b]. Defined [Hol59]. Definite [BR56a]. Degeneration [MRC56, BM58a, Via58]. Dehydrogenase [BP57, WM55, NS59]. Dehydrogenases [CZ59a]. Delineate [MBR56]. Demonstrable [WM55]. Demonstrated [MBMR55]. Demonstration [BP57, WL55, BS58a, CZ59a, HL59, NWS58b, WM59]. Dendritic [CH58]. Denervated [NP58]. Dense [Ban59]. Densitometry [Hal55]. Density [Mer58]. Deoxyribonucleic [LSW^{+}57, CvTF58, LBVS59, Pau56, RB58].

Deoxyribosenucleic [BS59]. Deoxyribosidic [FS59]. Dependent [CZ59a]. Depolarization [IH57]. Deposition [WL55, vBC55]. Deposits [Ric58]. Derivatives [Mil58]. Derived [WB55]. Dermis [PP59]. Descemet [Jak56]. Designed [Eps58]. Desmo [HN59]. Desmoenzymes [NPS56]. Desoxyribonuclease [OP57, Str56]. Desoxyribonucleic [BG55b, MKT59]. Desoxyribonucleoprotein [BG55a]. Desoxyribose [BO59, LS55]. Destroying [DOMtTAoJCWW57]. Destruction [DWM^{+}55]. Detail [Ste57]. Details [Gel57]. Detergents [WM58]. Determination [Bél59, EU57, Pau56]. Determined [HK56]. Developing [Gre58]. Development [BR56b, MHRM56, SP57, MFW59, SF59]. Developments [Bor56a, Bor56b]. Device [Pea59]. Diaminopurine [SD55]. Diamond [FM56, Eps58]. Diaphorase [NWS58a, NWS58b]. Diaphorases [CZ59a]. Diazo [LO59]. Dictyosomes [BTDA56]. Differences [BG55a]. Differential [dF59, HL58b, SM58]. differentialis [Rot57b]. Differentiating [Hay58]. Differentiation [Alf56, BF56, Cla57, Kem56a, FIS59, Fei56, LVH58, Str59a]. Differents [BBG59a, BBG59b]. Diffraction [BS56, FmF57, SB56, Fin59]. Dilute [DM55b]. Dimethyl [Str58]. Dioxide [SP58g]. Diphosphopyridine [NWS58b]. Disc [Mui57]. Dispar [MBR56]. Dissection [AF59]. Distal [RC56, Mil58]. Distinct [MKT58]. Distinctive [Sie55]. Distribution [AP57, DW55b, Doy55, ED55a, FHM56, JK57, LB58, PM56, WM55, dLAC58]. Dividing [BG55a, Koe58b]. Division [PM56, Cha59, FS59]. DNA [LS55, DOMtTAoJCWW57, Hal56, HL58a, HBW^{+}59, KRS58, Koe58b, LSW^{+}57, Lit59, LR56a, PM56, SP58b, Woo56]. Dog [Sul55]. Dollar [Mer59a]. Double [Fre56, Hux57, GH58]. Dpn [Coh56]. Dpnh [BB57]. Dried [Ger56, Fin58]. Droplet [Str57a]. Droplets [Str57b, Str56]. Drosophila [Gay56, MKT59, RR59]. Dry [HK56]. Ductuli [LY58]. Duplication [Gal59a]. During [Alf56, GB57, Gri56, KC56, Reb57, Ste56, TPD56, Woo58, DR58, KD58, NS59, RW59, RH59, Via58, Cha59, KWDV58, Ste58]. Dwarf [LS55]. Dynamics [Tay59].

Ear [SR58]. Early [CCP57, Via58]. Earthworm [DB55, Han57b, Ham59]. Eel [Luf56a]. Effect [Coh56, Hol59, KvH59, LR56a, OP57, WL56, WM58, HL59, HF59b, Kih59a, Mer59c, SS56, SP58b, SP58f]. Effects [BVLC57, Bie55, Cau57, FM57, Hod56a, LMS55, MMR56, RWPP55, FERS59, HD58, Kih59b, TC59]. Efferentes [LY58]. Egg [Rot57c, Str57a, PCV58]. Eggs [MM55, KD58, Mer59a, Mer59b]. Ehrlich [AP57, FERS59, KSB59]. Eisenia [Ham59]. Electric [Luf56a]. Electron [AP57, Afz59, AB56, AvB58, BP57, BTD55, BTDA56, BP59b, BM57a, BM57b, BM57c, BMG^{+}57, BF56, Cha59, CI59, Cla57, CH58, Coh57b, CCP57, DF56b, De 56c, De 56a, De 56d, DW55a, DW55b, DW56, ERdH58a, FW57, Faw56a, FmF57, Fin59, Fri59, Ger56, Gra59, GKF56, Hal55, HNS59, HHP57, Har56, HC59, KRS58, Kem56b, KGF56, KC56, KHD56, LC56, LW59, Luf56b, LH57, Lus56a, Lus56b, MP56, MRC56, MR57a, MG57, MBMR55, MHRM56, MMR56, MBR56, MGR^{+}57, Mos56b, Mui57, NM56, NBdD56, PQ57, PG59, Pit56, PR57b, Reb56a, Reb56b, RIF58, Ric58, Ris56, RC56, RPLL58, Ros59, Rot57a, Sel55, ST59, She56, SR57, SD59, SL57, STC57, Ste57, TPD56, TD56, UNG57, WG59, Wat57, Wei57, WK55, YI57, dHF58, dHC59, dL57, dL58a, vBC55]. Electron [BSLM59, Ber56, BS59, CMG58, DW58, Elb59, ENM58, FI58, Fin58, GG58, GT59, GH58, Hal56, HL58a, HM59, HSH^{+}59, HA58, LY58, LR58, Lat59, LO59, LF56, MB59, Mer58, MQ59, PK59a, PK59b, Pea59, Rud58, SR58, Sho58, SB59a, SP59b, Spe59, Sto59, TG58, TY59c, Wat55b, Wat58a, Wat58b, Woo59, WH59, YT58]. Elementary [Ban59, STC59]. Elements [FSF56]. Elicited [DO58]. Elongating [SB58]. Embedded [MMR56, RVH58]. Embedding [Bor56a, Bor56b, MG57, NM56, GG58]. Embryo [JC57, WK56, Fri59, Joh58, PP59]. Embryogeny [Woo56]. Embryology [Mui57]. Embryonic [MQ59]. Embryos [JSS59]. Emperipolesis [SD59]. Employing [DW55a]. Endings [DF57]. Endocrine [Lev56b]. Endoplasmic [ERdSSVF56, Pal55b, Pal56a, PP57, CD59, RPLL58, WMK59a]. Endothelium [Buc58, MC58]. Enteropneusta [CCP57]. Entwicklung [MFW59]. Enucleated [SP58f]. Enucleation [Coh56]. Envelope [Gal59b, Pap56a, Wat55a, Wat59]. Environment [Hod56a]. Enzymatic [BP59a]. Enzyme [Gri56, RH59]. Enzymes [EtTAoJH57, JK57, WL56, HN59, HSP58, SHP58]. Epicotyl [TS59]. Epidermal [BCCMS57, Gri56, Mat56]. Epidermis [CH58, Sel55, CI59, HC59, Obl58]. Epiphyseal [RC56]. Epithelia [Pea56]. Epithelial [Lac57, GT59]. Epithelium [MO55, MB55, She56, Yam55a, HSH^{+}59, HL59, Woo59, Yam58]. Ergastoplasm [BR56a]. Erythroblasts [BBG57]. Erythrocytes [BBG57]. Escherichia [Bac55, CvTF58, KR58, WZ59, ZW59]. Essential [SP58g]. Esterase [NH57, UHBdD56]. Esterase-Rich [NH57]. Esters [Fei56]. Estimation [NPS56, SB59a]. Ethylenediaminetetraacetic [DO58]. Euchromatin [dF59]. Euplotes [Rot56, Rot57a, Gal59a]. Evaluating [She57]. Evaluation [HF59b]. Evidence [AB56, BG55a, MKT58, FW57, Lit59]. Examined [MBMR55, Wat57, FI58]. Exceptional [LS55]. Exchanges [WARL58]. Excised [WMK59b]. Exhibiting [FIS59]. Exocrine [Pal56b]. Experimental [BS58b, KC59, Nov59]. Experimentally [Kih57, KMS57]. Experiments [WL56, LF56]. Extended [Mer59c]. Extract [MH55]. Extracted [FHM56]. Eye [Mil57, Yam59, EW59]. Eyes [GP57].

faba [Laf58b, Mer59c, SD58, SD55, Laf58a]. Fact [Orn57]. Factor [PH55]. Factors [CZ59a, DM55b, Kih56, WARL58, PH58]. Family [Rob56b]. Fasciola [GT59]. Fasted [ED55b]. Fate [Mer59a]. Features [DB55, HL58a]. Fed [ED55b]. Feeding [BVLC57]. Feline [Koe58b]. Femur [RC56]. fer [BBG59a, BBG59b]. Fern [Man59]. Ferritin [Ric58]. Fertilized [PCV58]. Fetal [RPLL58]. Feulgen [Hir57, LDK55, Nao57, Rud55, Woo57].

Feulgen-Stained [Rud55]. Fibers [Hes58, Koe55, KC59, Rob55, Ham59, Rob57, Rob58a, WG59]. Fibres [Han57b]. Fibril [Orn56c]. Fibrillogenesis [WK56]. Fibrils [BS56, WL56, Bea56, Gra59]. Fibrinogen [HS59]. Fibroblasts [BMC59, DO58, PP59, SP58g]. Fiction [Orn57]. Figures [RIF58, Sto59]. Filaments [Hux57, SB56]. Film [WH59]. Films [Wat56, BS59, DW58, Spe59, Wat55b, dH58]. Findings [LS55]. Fine [BF55, EW59, ET58, Eps57b, Eps57a, Eps59, Faw56b, FM56, GB57, HG58, Hes58, Hod56b, Jak56, KWDV58, KAWI59, KD56, Luf56a, MR57b, Nov56, PP55, Pap56a, RJ58, Rot56, RP55, Rud56, SP55, Swi56, TY59a, TY59b, Yam55a, Yam55b, BM58b, Buc58, Dip58, Edw59, FS58a, Ham59, Hay58, HC58, Lad58, NF58, Obl58, PB58, Rut59, SS59, STC59, TY59c, Tri58, WH59]. First [KD58]. Fixation [Gal59b, Mol59, NPS56, TI55, LDK55, Cau57, FS58b]. Fixative [Luf56b, Rud55, Afz59]. Fixatives [BL59]. Fixed [DM56, DM57, NH57, RVH58]. Fixers [LF56]. Flagellate [Gri59]. Flagellates [Rob56b]. flavicolis [Ans58]. Flight [Han56, Sac55, ERdH58b, SS56]. Flow [Ben56a]. Fluids [Eps57b]. Fluorescent [HMF57]. Fluoride [BVLC57]. foetida [Ham59]. Foliata [BB59]. Follicle [BM57a, BM57b, BM57c, TCS59]. Forbesi [KJTD55]. Form [DM55b, DOMtTAoJCWW57, Tho57, Lit59]. Formation [Cle56, Lus56b, RWPP55, Uzm56, UNG57, YFJ59a, YFJ59b, YFJ59c, BM58a, FW57, Laf58b, PP59, Str59b, Tay59, TSW58, WSW58]. Formations [Bin59]. Found [KGF56, Laf58a, Pea56, Yam59]. Fowl [Fri59]. Fraction [MM55, WS56b]. Fractionation [HL56, HL57, HL58b, LB59]. Fractions [Hol59, LB58, NBdD56, Str57a, PG59, SM58, dLAC59]. Fragments [SP58f]. Free [dHF58]. Freeze [FS58b]. Freezing [Ste56]. Frequencies [Kih59a]. Frog [DB55, Kem56a, TD56, Yam59, Rob57]. Frozen [Ger56, Ste57, Fin58]. Frozen-Dried [Ger56]. Function [LH57, RJ58]. Functional [RR59, Qui58]. Further [HN59, Rud56, Wat59].

Gall [Yam55a]. Gametogenesis [Woo56]. Gamma [OP57]. Ganglion [De 56d, DM56, DM57, Hes58, AvB58]. Gastrointestinal [Dzi56]. Gecarcinus [HC58]. Gelatin [Coh55]. Gels [DM55b, DOMtTAoJCWW57]. General [Kar56c]. Germ [BTDA56]. German [MFW59]. Gey [Ros57]. Giant [Koe55, Ham59]. Giving [GH58]. Gland [CD59, HC58, MB59, Tri58]. Glands [MKT59, LB59, RV59, Str59b]. Glass [Cam56, She57, Eps58, GL59, TO59, Wei59]. Globoid [HHPvR56]. Glomerulus [PR57b, Yam55b]. Glossiphonia [BM58b]. glucosamine [Car57]. glucuronidase [Str56]. Glutathione [LS59]. Glyceraldehyde [NS59]. Glycerol [FHM56]. Glycerophosphatase [BB59]. Glycogen [TI55]. Gold [HHP57]. Golgi [CS56, Cle56, DF56a, FW57, LC56, Lac57, WMK59a]. Gradient [AR58]. Grain [RMFW59, Woo58, WH59]. Grana [HMM55]. Granule [Ban59, FW57]. Granules [Ben56c, Cas59, FD56, Gal56, Pal56b]. Grasshopper [BTD55, DR58, GB58, TPD56]. Green [RB58]. Grid [Elb59]. Groups [BS58a, FIS59, KD58]. Growing [HHPvR56, Kem56b]. Grown [MQ59]. Growth [Coh57b, Gri56, Ste56, Woo58, BM58a, FJ58, KWDV58, LVH58, RW59, RSK59, WLP59]. Guinea [De 56d, LY58, SP58c, SP58d, SP58e, SP59a]. Gut [GT59].

Hair [BM57a, BM57b, BM57c, Gri56]. Hamsters [DS58]. HCl [SP58b]. Heart [Kis56, Mui57, MQ59]. Heating [Wat57]. Heavy [Wat58a, Wat58b]. HeLa [BMG^{+}57, GM59c, HHPvR56, HHP57, LB57, Ros57]. Helical [Coh55, Pap56b]. Helix [TG58]. Hematoencephalic [vBC55]. Hemipteran [Ans58]. Hemosiderin [Cas59, Ric58]. Hensel [BF56]. hepatica [GT59]. Hepatocytes [LR56a, LBVS59]. Hepatoma [LW59]. herculeus [GD59]. Hetero [dF59]. Hetero- [dF59]. Heteromorphic [BMC59]. Hexagonus [CCP57]. High [GL59]. Histochemical [BP57, BS58a, BMG^{+}57, KJTD55, MRC56, Sul55, TI55, TC59, CZ59a, Fei56, HN59, MGR^{+}57, NWS58a, NWS58b, NP58, WM59]. Histochemically [WM55]. Histology [Mon55, Mon57]. Histone [BG55b, WZ59, ZW59]. Histones [Ans58]. Holder [Elb59, Eps58]. Homogenate [DWM^{+}55, PG59]. Homogenates [KHD56]. Homogenizers [And56]. Horseradish [Str58]. Houses [Rob56b]. Human [BM57a, BM57b, BM57c, CH58, Hol59, KGF56, MO55, MB55, Mon55, Mon57, LDK55, Bad59, CI59, DM55a, GM59a, HSH^{+}59, HC59, Obl58, RPLL58, dHC59]. Humor [WB55]. Hydra [CT59a, CT59b, SF59, Woo59]. Hydroides [CCP57]. Hydrolases [LB58]. Hydrolysis [Woo57, SP58b]. Hydrolytic [HN59]. Hydronephrosis [Nov59]. Hydroxyuracil [SD58]. Hypertonic [Rob58a]. Hypertrophy [BO59]. Hypotonic [Rob58a].

Ice [Ste56]. ichthyocol [GSM57]. Identification [CS56, Str59a]. Ii [OP57, Jak56]. III [KP55, Mat58]. Illustrated [Hal56]. Imbedding [DM59, GH58]. Immature [DBDA57, She56]. Immersion [FM57]. Immunochemical [FHM56, Wen59a]. Impregnation [WK56, CMG58]. Improved [BR56b, Elb59]. Inanition [Coh57b]. Inclusions [Kem56a, KGF56, LW59, KAWI59]. Incorporation [PS58, TS59, HL59, PR57a]. Index [AR58, AF59]. Indirect [Han56]. Induced [Kih57, KMS57, Lev56a, LR56b, dHF58, Bin59, Gut57, LD59, LS59, WARL58]. Infant [RC56]. Infected [GM59c, Mat58, LB57, MGR^{+}57]. infectées [SRK59]. Infections [KAWI59]. Inferences [Ted59]. Influence [DWM^{+}55, PH58]. Influencing [DM55b, CZ59a]. Infolded [Pea56]. Infusionum [RP55, Rud56, Rud58]. Ingestion [HHP57, Cla59]. Inhibition [Tay59]. Inhibitors [Kih59a]. Injected [Str58, Str57b]. Injection [EU57, Str57a, AvB58]. Inner [Yam59]. Innervation [Edw59]. Inoculation [FD56]. Inosine [Sie55]. Insect [MBMR55, Sac55, Smi56, Edw59, SS56].

Insects [GP57]. Insoluble [HL59]. Intact [DWM^{+}55]. Integrated [BB57]. Interaction [Orn56b, WB55]. Intercalated [Mui57]. Intercellular [Tha59, Woo59, FIS59]. Interference [Sid57]. Interferometric [HK56]. Internal [PERM56, Ric59]. Internodal [Gre58]. Interpretations [WK55]. Interrelations [Gay56]. Interrelationship [Obl58]. Intervals [KWDV58]. Intestinal [HSH^{+}59, PK59a, PK59b]. Intestine [Cla59]. Intracellular [EtTAoJH57, ED55a, Faw56a, UHBdD56, Woo58, dLAC58, FK59, RIF58, Str59b]. Intracisternal [Pal56b]. Intracytoplasmic [AP57]. Intranuclear [BMG^{+}57, KGF56, LW59, Bin59, LB57]. Intraperitoneal [FD56, Str57a]. Intrasomatic [Bac55]. Intravenously [Str58, Str57b]. Intravitam [WL55]. Invertebrate [Reb56a, Reb56b]. Investigation [MO55]. Iodoacetate [LD59]. Ion [Ben56a, Koe55]. Ionic [FM57]. Iron [Ber56, BBG57, Rob56b]. Iron-Containing [Rob56b]. Irradiation [OP57]. Isolated [DM55b, DOMtTAoJCWW57, ED55b, HMM57, LH57, Nov56, OP57, WWC55, Str56, Fri59, LC58, Mer59b, PG59, SJS59, WM58, dLAC59]. Isolating [DWM^{+}55]. Isolation [BR56b, MLKD56, WS56b, JSS59]. Isometric [KC56]. Isotonic [KC56].

Junction [Rob56a, dHC59]. Junctions [ERdH58a, ERdH58b].

Kappa [Dip58, HG58]. Keratin [Mat56, Mat56]. Keratins [Mat56]. Kidney [RMW57, Str57a, vBRC56, Str56, NH57, SBF58, SNM58]. Kidneys [Cla57, Str57b]. Killer [Dip58]. Kinetin [Gut57]. Kitten [BM58a]. KMnO [Gal59b]. Knife [FM56, MMR56, Eps58, Wei59]. Knives [Cam56, GL59, She57, Eps58, TO59]. Kowalevskii [CCP57].

L [GT59, KSB59]. L. [HMM55]. Lacking [LL55]. Lamella [WF56]. Lamellae [HMM55, Swi56, Mer59a]. Lamellar [HMM56, Tho57, Rut58]. Lamellated [Yam58]. Lampbrush [LR58]. Land [HC58]. Lanterman [Rob58b]. Large [RM57, Buc58]. Larva [RR59]. Larvae [Por56, Hay58]. lateralis [HC58]. Laterosporus [Han57a, FJY58]. Lattice [MBMR55]. Lattices [Ric58]. Laubach [Han57a]. Layer [Hod56a]. Leaf [Mat58]. Leech [BM58b]. Left [MB59]. Leg [ERdH58a]. Lengths [Spi56]. Lens [IH57, WG59]. Lesions [BSLM59]. Leucine [TS59]. Leucine-C [TS59]. Leukemia [dHF58]. Level [Coh56]. Light [IH57, Mos56b, TPD56, Wei57, GT59]. lignorum [Str59b]. Like [AP57, Faw56a]. Lilium [FS59, NS59, Ste58]. Limbs [Hay58]. Limiting [WARL58]. Limnoria [Str59b]. Limpet [Lac57]. Limulus [Mil57]. Line [RC56]. Lining [TG58]. Linn [TG58]. Lipide [KSB59]. Lipides [KJTD55]. Lipogenesis [KC59]. Lipoprotein [Hod56a]. Liposomes [Bel58]. Liver [BR56a, BR56b, ED55a, KGF56, KHD56, LW59, LMS55, LR56b, MLKD56, NBdD56, Nov56, PS56a, RB56, ENM58, Fin58, KvH59, LRS59, LC58, LL55, NH57, PH58, Sie55, UHBdD56, WM58, dLAC58, dLAC59]. Livers [ED55b, FS56]. lividus [PCV58]. Living [DM56, DM57, TB56, LB57].

Lizard [Smi59]. Lobster [MH55]. Localisations [PCV58]. Localization [EtTAoJH57, JK56, BSLM59, CvTF58, FK59, HSP58, LO59, NWS58a, SHP58, UHBdD56]. Locust [GB57]. Long [Hol59, WPE^{+}58]. Long-Term [Hol59]. longiflorum [FS59, NS59]. Lotmar [Coh58]. Low [FmF57]. Low-Angle [FmF57]. Lower [Mer58]. Loxa [Ans58]. Lucidum [BP59b]. LuckÉ [Faw56a]. Lumbricus [GD59]. Lung [Kar56c, Kar56a, Kar56b, Kar58]. Lymphatic [Doy55, HL56]. Lymphocytes [Tro57]. Lymphopoiesis [Tro57]. Lyo [NPS56, HN59]. Lyo- [NPS56]. Lysosome [Ben56c, NBdD56]. Lysosome-Rich [NBdD56].

M [BP59a]. Macromolecular [MBMR55]. Macromolecules [BS59, Hal56, HL58a]. Macronuclear [Gal59a]. Macronucleus [RP55, Rud56]. Macrophage [Kar58]. Macrophages [FD56]. Maintenance [BM58a]. Maize [WMK59a]. Making [TO59]. Malaria [RT59]. Male [BTDA56]. Malononitrile [AvB58]. Malpighian [BTD55]. Mammalian [BF55, CER57, DW55b, MR57a, Sel55, Koe58a]. Mammary [LMM^{+}59, MLM^{+}59]. Man [Wei57]. Mass [Bél59]. Masson [CH58]. Mast [SL57, WLP59]. Material [AB56]. Materials [RMFW59, Cla59]. Matrix [BVLC57, RC56]. Matter [Odo56]. Maturation [GB57]. Mature [She56]. Maturing [Mer59a]. May [Lit59]. mays [HMM55, LVH58]. Means [Wen59b, WL55]. Measles [KAWI59]. Measurement [AF59]. Measuring [RC57]. Mechanism [AR59, HMM56, DGW58]. Mechanisms [Ben56a]. Medium [BR56b, FM57, Hol59, GG58]. Medulla [DF57]. megaterium [FJ58]. Meiosis [Ans58, Ste58]. Meiosporangia [RM57]. Meiotic [De 56a, Faw56b, Mos58, Rut59]. Melanin [FD56]. Melanocyte [CH58, CI59]. melanogaster [RR59]. Melanoplus [dF59, Rot57b]. Member [GMC55]. Membrane [AB56, Ben56a, Fre56, Gay56, GKF56, Jak56, Rot57c, WS56b, Cha59, ENM58, Gri59, KR58, Rot57b, SS59, SB59a, Ted59]. Membrane-Cytoplasmic [Gay56]. Membranes [GM59c, Pea56, vBRC56, FI58, Gri59, RH59, RIF58, Rob57, SNM58]. Membranous [GH59]. Mercaptopurine [Bie55]. Mercuric [HM59]. Meristematic [Laf58b]. Mesothelial [FD56]. Metabolically [MKT58]. Metabolism [DBDA57, LR56b, MKT59, Sac55, CZ59b, Gut57]. Metachromasy [BS58b]. Metals [FM56, Wat58a, Wat58b]. Metaphyses [DBDA57]. Methacrylate [Bor56a, Bor56b, MG57, MMR56, RVH58]. Methacrylate-Embedded [MMR56]. Methionine [EU57]. Methionine-S [EU57]. Method [Bél59, DWM^{+}55, Hal56, She57, BSLM59, DM59, Fei56, GH58, HA58, LO59, Men58a, Men58b, Men58c, NWS58b, RC57, RVH58, Str59a, TO59, Wei59]. Methods [FS58b, HK56]. Methyl [RB58]. Mice [Cla57, FD56, Lus56b, Cla59]. Microanatomy [Man59]. Microbodies [RB56]. Microfibrils [SB58]. Micrographs [AB56, WK55, Hal56, RIF58]. Microkinetospheres [Ros57]. Microphotometric [BG55b]. Microscope [AP57, BP57, BTD55, BTDA56, BF56, Cla57, CH58, CCP57, DF56b, De 56c, De 56d, DW55b, Faw56a, FmF57, IH57, LW59, Lus56b, MR57a, MBMR55, MHRM56, MBR56, Mos56b, Mui57, Pit56, Ris56, Rot57a, Sel55, She56, SR57, STC57, TPD56, TD56, UNG57, Wat57, Wei57, dHF58, vBC55, Cha59, CI59, Elb59, Fin58, Fin59, Fri59, GT59, Hal56, HL58a, HC59, KRS58, LR58, MQ59, RPLL58, Rud58, SR58, Sho58, SP59b, Sto59, TG58, TY59c, WG59, dHC59]. Microscopic [BMG^{+}57, Coh57b, De 56a, DW55a, DW56, KHD56, LH57, Lus56a, MRC56, dL57, AvB58, FW57, HNS59, LY58, LO59, MGR^{+}57, PK59a, PK59b, Ros59, RC57, dL58a]. Microscopical [KC56]. Microscopy [BM57a, BM57b, BM57c, Ger56, GKF56, HHP57, Har56, Kem56b, KGF56, LC56, Luf56b, MP56, MG57, NM56, NBdD56, PQ57, PR57b, Reb56a, Reb56b, RC56, ST59, SD59, Sid57, SL57, Ste57, YI57, Afz59, BSLM59, Ber56, BP59b, BS59, CMG58, DW58, ERdH58a, ENM58, FI58, GG58, Gra59, GH58, HM59, HSH^{+}59, HA58, Lat59, LF56, MB59, Mer58, Pea59, PG59, Ric58, Spe59, Wat55b, Wat58a, Wat58b, Woo59, WH59, YT58]. Microsomal [TBV56, TS59, RH59, Sie55]. Microsomes [PS56a, PS56b]. microspectrophotometric [LDK55]. Microspectrophotometry [Nao57, MR58, Men58a, Men58b, Men58c]. Microspores [NS59]. Microsporocyte [Ste58]. Microstructure [GP57]. Microtome [MMR56]. Microtomes [Eps58]. Microtomy [Wei59]. Midgut [Str59b]. Milk [LMM^{+}59]. Mitochondria [BR56a, Bie55, BR56b, BP59c, Dem56, DWM^{+}55, FS56, Fre56, Har56, How56, Lev56a, LC58, LMS55, OP57, PB59, PERM56, SR56, SW56, TPD56, TB56, WM55, WS56a, WWC55, Bel58, CD59, LRS59, LL55, SS56, Str59a, WM58]. Mitochondrial [BP56, DS58, DWM^{+}55, DOMtTAoJCWW57, LS59, Low56, Orn56b, RB56, WS56b, Ted59]. Mitosis [Bie55, Ans58]. Mitotic [PM56, Wen59b, NS59, Wen59a, WMK59b]. Mode [HS59, Laf58b]. Model [Gro56]. Modification [ED55b]. Molecular [LSW^{+}57]. Molecule [HS59]. Molecules [Bea56]. Mollusc [Wol58]. Molluscum [Ban59]. mori [RV59]. Morphogenesis [De 56b, Gro56, Hod56a, HMM56]. Morphological [BB57, CS56, FJY58, HL58a, TPD56, YFJ59a, YFJ59b, YFJ59c]. Morphology [AP57, DB55, De 56a, Mil57, WWC55]. Mortis [Loc59]. Mosaic [Mat58]. Mother [Gré57]. Mother-Of-Pearl [Gré57]. Mounts [NM56]. Mouse [Eps57b, Kar56c, Kar56a, Kar56b, LW59, She56, UNG57, Yam55a, Yam55b, dHF58, Kar58, LMM^{+}59, MLM^{+}59, Mun58, NF58, SBF58]. Movement [AR59, AR58]. Movements [TB56]. Mucopolysaccharides [GMGL57, TC59]. Mucoproteins [Sul55]. Mucosa [Dzi56, dL57]. Multimicronucleatum [SP55]. Multipurpose [RPST58]. Multiterminal [Edw59]. Muscle [Ben56b, CER57, ERdSSVF56, FHM56, Han57b, Hod55, Hod56b, How56, Hux57, KC56, Loc59, MRC56, MR57a, Por56, Sac55, Spi56, Edw59, MQ59, SS56, Sho58, Tha59]. Muscles [SB56, WM55, ERdH58b, NP58]. Musculature [Kar59]. Mycelia [ST59]. Myelin [FmF57, FM57, Fin57, Lus56b, Uzm56, BM58a, RIF58, Rob58b, Sto59].

Myelinated [Rob55]. Myelination [DGW58]. Myelinogenesis [Rob55]. Myofibrils [Han56]. Myogenesis [HMF57]. Myoneural [Rob56a]. Myosin [FHM56].

N [Str58]. N-Dimethyl-p-phenylenediamine [Str58]. Nature [Ben56c, RM57, HF59b, RB58]. Negative [Ric59]. Nematocysts [CT59a, CT59b]. Neoplastic [WLP59]. Nerve [CER57, De 56d, DF57, FmF57, FM57, Fin57, Hes58, Koe55, MH55, Rob55, Sul55, TD56, dL57, Ham59, Rob57, Rob58a]. Nerves [Uzm56, UNG57, ERdH58a]. Nervous [Har56, Lus56a, Lus56b, Pal56c, vBC55, DGW58]. Neuraxis [Koe58b, Koe58a]. Neuroglia [Hes58]. Neuroglial [Gra59]. Neuromuscular [ERdH58b, ERdH58a, dHC59]. Neuronal [FK59]. Neurons [Lac57, PP55, Smi59]. Neurospora [ST59]. neutral [GSM57]. Newborn [Cla57, RC56, NF58]. Newly [PM56]. Newt [SS59]. Nissl [DM56, DM57, Smi59]. Nitella [Gre58]. Nitrate [DW55a, DW55b]. Nodes [UNG57]. Non [BG55a, Koe58b]. Non-Dividing [BG55a, Koe58b]. Normal [Bac55, She56, Str56, BBG57, DS58, NP58, RW59, SBF58, SR58]. Note [Cam56, DM55a, And56]. Noted [Pea56]. Novel [BS59]. Nucleal [Hir57, Woo57]. Nuclear [Alf56, AB56, BG55b, Gal59b, Gay56, GM59a, GM59c, GKF56, HK56, Koe58b, Orn56b, Pap56a, Pre59, Reb57, Wat55a, FS59, GM59b, Gri59, Gut57, Wat59]. Nuclear-Cytoplasmic [GM59a]. Nucleated [SP58f]. Nuclei [DM55b, DWM^{+}55, DOMtTAoJCWW57, ED55b, GB57, Mos56b, Nao57, Mer59b, RW59, SB59b]. Nucleic [FSF56, Woo58, BO59, Ber56, Koe58a, LS55]. Nucleoli [Laf58a, MLKD56, JSS59, MKT58, SJS59]. Nucleolonemata [DM55a]. Nucleolus [Laf58b, Lov57]. Nucleoprotein [RSK59, TBV56]. Nucleoproteins [PH55]. Nucleotidase [ENM58, dLAC58]. Nucleotide [NWS58a, NWS58b]. Nucleotides [JK57]. Nucleus [DR58, De 56a, Gal56, GB58, BO59, Pap56b]. Number [WARL58]. Nutrient [SP58g]. Nutrition [MM57, MPM56].

Objects [RC57]. Observation [Hal56]. Observations [Bac55, BV59, Coh57a, Coh57b, DF56b, De 56c, De 56a, DW56, Eps59, Faw56a, FI58, FS58a, Hod56a, LH57, Lus56a, Man59, MRC56, Rob56b, RP55, Rud56, WK56, dL57, vBRC56, AvB58, BB59, BM58a, Cha59, CI59, Dip58, Fri59, HNS59, Ham59, HC58, MGR^{+}57, RV59, Rob57, RPLL58, Ros59, SR58, Wat59, dL58a]. Observed [BVLC57, MHRM56, Ros57, Yam58]. Obtained [Afz59, HSH^{+}59, SM58]. Obtrigonoidea [Pit56]. Occurrence [FIS59, LB57, Ric58, Sie55]. Occurring [MGR^{+}57]. Octopus [Wol58]. Ocular [WL55]. Olfactory [dL57]. Ommatidia [Mil57]. Oocyte [AR57, Gal56]. Oocytes [Kem56a, Kem56b, Reb56a, Reb56b, STC57, TCS59, Wis57]. Opalina [Pit56]. Optical [Coh55, Mun58]. Organ [Luf56a]. Organic [BVLC57, Gré57]. Organisation [Eps57a]. organisme [BBG59a, BBG59b]. Organization [DR58, De 56c, PERM56, DL58b]. Organs [HL56]. Origin [Bel58, Mer59a, TCS59]. Orotic [Koe58b]. Orotic-6-C [Koe58b]. Orthoptera [BTD55]. Osmiophilia [Orn57]. osmium [LDK55, Mer58, NH57, RVH58]. OsO [Cau57, LDK55, Rud55]. Osteones [AF59]. other [TS59, BS59, PH58]. Otocyst [Fri59]. Ovarian [Bad59]. Oviduct [HDG57]. Ovular [TCS59]. Ovum [SP59b]. Ox [Loc59, WB55]. Oxidase [BB57]. Oxidation [LL55]. Oxidative [FS56, HSP58, SS56, SHP58]. Oxides [Mer58].

p [Str58]. Pacinian [PQ57, Qui58]. Packing [LP56]. Pairing [Mos58]. Pancreas [Pal56b, RH59, SP58c, SP58d, SP58e, SP59a, WM59]. Pancreatic [PS56b, Mun58]. Papilla [BB59]. Paracentrotus [PCV58]. Parachromatin [Lov57]. Paracrystalline [MBMR55]. Paramecium [SP55, Dip58, HG58]. Paramyosin [BS56, LS57]. Parasite [RT59]. Parasporal [Han57a, FJY58]. Parathyroid [Tri58]. Parenchyma [SB58]. Parenchymal [KGF56]. Parietina [AR57]. Particle [WB55]. Particles [AP57, Faw56a, Hal55, HMM57, RM57, Smi56, TBV56, TS59, BBG57, FH59]. Particular [Koe55, LC56]. Particulate [Laf58a, Odo56, Pal55a]. Particulates [KHD56, Nov56]. Partitioning [Bac55]. Passage [AB56]. Patella [Lac57, Rot56, Rot57a]. Pathological [Mat58, BBG57, RW59]. Pattern [Koe55]. Patterns [Loc59, BM58a]. Pea [TBV56, TS59, JSS59, SJS59]. Pearl [Gré57]. Pecten [Mil58]. Peculiar [Yam58]. Pellicle [ET58]. Pellucida [TCS59]. Penetration [CCP57]. Peptidase [Doy55]. Peptone [Hol59]. Perfused [LH57]. Periosteum [TC59]. Peripheral [FM57, Fin57, Rob55, Uzm56, ERdH58a, Rob57]. Periplaneta [Hes58]. Peritoneal [Odo56]. Permanent [SP58g]. Permanganate [Luf56b, Mol59]. Permeability [Mer59b, Ted59]. Permeable [GH58]. Peroxidase [Str58, Str59a]. pH [GSM57]. Phagosomes [Str59a]. Phagotrophy [RT59]. Phase [BR56a, Ros57, FI58]. phenylenediamine [Str58]. Phosphatase [ED55a, ED55b, Str56]. Phosphatases [WM59]. Phosphate [NS59]. Phospholipide [RH59, Fin59, RIF58]. Phosphorylation [FS56, LL55, SS56]. Photochemical [FH59]. Photographic [RC57]. Photometric [HF59b]. Photoreceptor [Sid57, WCT57]. Physical [Hod56a, MLM^{+}59]. Physiological [Coh57a]. Physiologically [Lev56a]. Picken [Coh58]. Pig [De 56d, LB59, LY58, SP58c, SP58d, SP58e, SP59a]. Pigeon [How56]. Pigment [Yam58]. Pigs [BVLC57]. Pinocytic [Ros57]. pipiens [AvB58, Kem56b]. Pipistrellus [Yam58]. Pituitary [LB58, LB59]. Placenta [DW56]. Plant [HMM56, HMM57, Mol59, RW59, RSK59]. Plants [Kih57, KMS57]. Plasma [Pea56, ENM58]. Plasmodium [RT59]. Plasms [KRS58]. Platelets [Wis57]. Plethodon [Mos58]. Plexus [Cas59, MP56]. Pneumococcus [KP55]. Polarization [Mun58, TD56]. Polio [KWDV58].

Pollen [RMFW59, Woo58, SB59b]. poly [Car57]. poly- [Car57]. Polyhedral [MBMR55]. Polymerization [HS59]. Polymerized [LL55]. Polymethacrylates [DM59]. polymorphus [RJ58]. Polypeptide [Coh55]. polyps [LDK55]. Polysaccharides [KP55]. Polyvinylpyrrollidone [Nov56]. Polyvinylpyrrollidone-Sucrose [Nov56]. pomatia [TG58]. Pores [AB56]. Porphyridium [BV59]. Porthetria [MBR56]. Positive [WH59]. Possible [HMM56]. Posterior [AR58]. Potassium [Kih59b]. Precursor [PR57a]. Prefixation [HD58]. Pregnancy [PH58]. Pregnant [Sho58]. Preliminary [LH57, RWPP55, dHF58, Bel58, Dip58, MGR^{+}57]. Preparation [BB57, Ger56, And56, Spe59, Wei59]. Preparations [Rud55, Sie55]. Presence [Ric58]. Preservation [Nov56]. Primary [BCCMS57, GM59b]. Principles [FS58b]. Problem [RB56]. Problems [MG57, Qui58]. Procedures [HL58b, HF59b]. Produced [GMGL57, KWDV58, Kih59a, Rob58a]. Production [Kih56, LMM^{+}59]. Proliferating [BG55a]. Properties [DWM^{+}55, HMM57, LS57, LOK56, MLKD56, GSM57, MLM^{+}59, Mun58, SJS59, Ted59]. Prophase [Faw56b, Mos58]. Proplastiden [MFW59]. Proplastids [MFW59]. Protein [Woo58, BS58a, KSB59, Koe58a, Lit59, MGR^{+}57, SP58f, Str57b]. Proteins [Alf56, MH55, Cla59, RW59]. proteus [Pap56b, PR57a, PS58, Pre59, Coh57b, GKF56, Pap56a]. Protoplasmic [Tha59]. Protoplasts [FJ58]. Protozoa [SR56]. Protozoan [Rot57a, Gal59a]. Proximal [RMW57, Nov59]. Pseudobranch [CD59]. Pteridium [Man59]. Pulmonate [Reb57]. Pumping [Ben56a]. Purified [BB57, MH55, GSM57]. Purkinje [KC59]. Pyloric [KJTD55]. Pyridine [JK57]. pyriformis [ET58].

Quantitative [NPS56, SB59a]. quantitatives [SRK59].

Rabbit [Mui57, DL58b, HNS59, HL57, LD59, SR58, Sho58, TCS59, dL58a]. Radiation [FERS59, RC57, WARL58]. Radiomimetic [Kih59b]. Ram [SB59a]. Rana [AvB58, Kem56b]. Ranvier [UNG57]. Rapid [Ste56, Str59a, BSLM59]. Rapidly [BG55a, HHPvR56]. Rat [Cle56, DW55a, ED55a, Gri56, KGF56, LMS55, LR56b, LR56a, NBdD56, Odo56, RMW57, Str57a, Str58, Wei57, WL55, vBRC56, BO59, BM58a, Fin58, KvH59, Lad58, LRS59, LC58, LBVS59, NP58, NF58, NH57, PH58, Sie55, SP59b, Str57b, Str59a, UHBdD56, WM58, dLAC58, dLAC59]. Rats [Dzi56, DBDA57, ED55b, Lus56b, Str56, Cla59]. Ray [BS56, FmF57, Luf56a, SB56, Fin59]. Rays [Kih59a]. Re [Tro57]. Re-utilization [Tro57]. Reaction [Hir57, Nao57, Woo57, BB59, Lat59, RB58]. Reactions [Wen59b]. Reconstitution [WL56]. rectal [LDK55]. Red [NP58]. Reduction [Wat57]. Reduplication [GM59c]. Reevaluation [BS58b]. Reference [ERdSSVF56, Koe55, LC56]. Refractive [AR58, AF59]. Refractometry [Sid57]. Regenerating [Hay58]. Regeneration [RB56, KvH59]. Related [Rob58b]. Relation [HMM57, MKT59, Mos58, Rob55, FS59, KSB59]. Relationship [BR56a, Gal56]. Relationships [SR57, TCS59, GM59a]. Renal [Faw56a, PR57b, Yam55b, BO59]. Report [RWPP55, dHF58, Bel58]. Reptilian [Rob56a, EW59]. Respiratory [Kih59a]. Responses [DO58]. Results [Afz59]. Rete [LY58]. Retention [SP58b]. Reticular [Smi59]. Reticulum [Ben56b, ERdSSVF56, Pal55b, Pal56a, Por56, PP57, CD59, RPLL58, SS59, WMK59a]. Retina [EW59, Lad58, Mil58, TY59c, Yam58]. Retinal [DF56b, De 56c, De 56b, Wol58, DL58b]. Revealed [Ric59, YI57, ENM58, Woo59, YT58]. Reversal [LRS59]. Rhodanese [WPE^{+}58]. Rhodospirillum [FH59, HF59a]. Ri [BMG^{+}57, LP56]. Ri-Apc [BMG^{+}57]. ribonuclease [Str56, SP58f]. Ribonucleic [Gri56, LR56b, MKT59, dLAC59, Gut57, LL55, MKT58, PR57a, Pre59]. Ribonucleoprotein [RM57]. Ribonucleoproteins [BMC59, PH58]. Rich [NBdD56, NH57, SB56]. Rigor [Loc59]. RNA [Gal56, GM59b, Lit59]. Rod [Yam59, Lad58]. Rods [DF56b, De 56c, De 56b]. Role [Cle56, ERdSSVF56, Fin57, SP58g]. Rooster [BSB59]. Root [JK56, WMK59a]. Roots [SD55, LVH58, SD58, WMK59b]. Rotation [Coh55]. Rous [Eps57a]. Route [RMFW59]. rubrum [FH59, HF59a].

S [Dzi56, DBDA57, EU57, JC57, BL59, BVLC57]. Saccoglossus [CCP57]. Salamander [Mos58]. Salivary [MKT59]. Salmon [Alf56]. salt [GSM57]. Same [DWM^{+}55, Lit59]. Sampled [KWDV58]. Sand [Mer59a]. Sarcoma [Eps57b, Bin59]. Sarcomere [Spi56]. Sarcoplasmic [Ben56b, Por56]. Sarcosomes [Kis56]. Satellites [Ros57]. Saturated [Fin59]. Scanning [MR58]. Scattered [RC57]. Schmidt [Rob58b]. Schwarzschild [HF59b]. Sciatic [UNG57]. Sea [Rot57c, Wen59b, KD58, Wen59a]. Secale [dF59]. Secreting [HDG57]. Secretion [BP56, RH59]. Secretory [FW57, WM59]. Section [Cos57, De 56d, Orn56a, Wei59]. Sectioning [FM56, Gel57, GL59, HA58, TO59]. Sections [MBMR55, MMR56, MBR56, NPS56, Sel55, Wat57, WK55, GH58, Pea58, Pea59, RVH58, SP58a, Wat55b, Wat58a, Wat58b]. Sedimentation [KP55, Mat56]. Seedlings [TBV56]. Seen [Dzi56, HL58a]. Segment [Yam59]. Segregation [Str57b]. Sense [Mil58]. Sepia [Wol58]. Serial [BM58a, MBR56, WK55]. Serum [RWPP55]. Several [DW55b, HN59]. Sex [PH58]. Shape [HS59]. Shaped [Rob56b]. Shearing [Rob58b]. Sheath [FmF57, Rob58b]. Sheaths [Hes58]. Sheet [WH59]. Shell [TSW58, WSW58]. Should [Cos57, Orn56a]. Showing [FS56]. Siemens [Elb59]. Significance [RR59, AR58, HN59, HF59b]. Silica [GL59]. Silk [RV59]. Silver [CMG58, DW55a, DW55b, WK56, WL55, vBC55]. Similarity [HBW^{+}59]. Simple [Bél59]. Simulating [RIF58]. Single [WK55, KWDV58]. Sinus [HC58]. Sinuses [Wei57]. Site [GM59b]. Sites [BP57, BP56]. Six [SP58g]. Size [HS59]. Skeletal [KC56, WM55]. Skin [MO55, MB55, Mon55, Mon57, Sel55, WF56, SS59]. Slices [CZ59b]. Small [Gal56, MR57b, Pal55a, Cla59, TG58]. Smooth [CER57, Han57b, Sho58, Tha59]. Snail [Reb57]. Sodium [EU57]. Solids [Sid57]. solubility [GSM57]. Soluble [MM55, FS59, Sie55]. Solute [MC58]. Solutions [WB55, GSM57, Rob58a]. Somatic [KP55]. Some [DB55, DBDA57, Ham59, Hod56a, HMM57, HC58, LS57, LF56, MMR56, Reb56a, Reb56b, RWPP55, STC57, SJS59, Wen59a]. Sources [HBW^{+}59]. Space [Lev56b]. Special [ERdSSVF56]. Species [GB58]. Specific [WM59]. Specified [KWDV58]. Specimen [Wat56]. Specimens [Ste57, Bad59, Mer58]. Sperm [CCP57, GB57, Afz59, SB59a]. Spermatid [BF56, Cle56, GB58, Rot57b]. Spermatocyte [Rut58]. Spermatocytes [Faw56b, Mos56a]. Spermatogenesis [Alf56, TPD56, YI57, YT58]. Spermatozoid [Man59]. Spermiogenesis [DR58, GD59, Reb57]. Spider [AR57, STC57]. Spinal [DM56, DM57, AvB58]. Spindle [Rut59, Tay59]. Splanchnic [DF57]. Spleen [EtTAoJH57]. Splenic [Wei57]. Spore [YFJ59a, YFJ59b, YFJ59c]. Spores [FJY58]. Squid [Koe55]. Stability [Wat56]. Stabilizing [PH55]. Stage [Gri56]. Stages [CCP57]. Stain [DW55a, DW55b, Ber56, HM59]. Stained [Hal55, Nao57, Rud55]. Staining [Ric59, Wat58a, Wat58b, Pea59, RVH58]. Stains [SB59a]. Starfish [KJTD55]. Starvation [Coh57a, PH58]. Stem [Bad59]. Stentor [RJ58]. Sterile [LS55]. Steroid [BP56]. Stimulation [DF57]. Stock [Dip58]. Strain [Ros57, KSB59]. Strains [SP58g, WPE^{+}58]. Strength [FM57]. Streptomyces [GH59]. Striated [ERdSSVF56, Hod56b, Hux57, Spi56, Kar59]. Striation [Han56, Loc59]. Strip [RPST58]. Striped [Ben56b]. Stroma [HL59]. Structural [Eps57a, Gre58, LS57, Rob58a, Ste57, Mer59b]. Structure [BCCMS57, BS56, BP56, BF55, Dem56, ET58, Eps57b, Eps59, Faw56b, FM56, FM57, Fin57, FH59, GB57, Han57b, Hes58, Hod55, Hod56b, HMM57, How56, Jak56, KD56, Laf58b, LSW^{+}57, Low56, Luf56a, LH57, MO55, MR57b, MHRM56, MBR56, MFW59, Nov56, PP55, Pap56a, Ric59, Rot56, RP55, Rud56, Sac55, SP57, SP55, SB56, Sid57, Smi56, Swi56, Wei57, Yam55a, Yam55b, Car57, Bea56, BM58b, Buc58, Dip58, EW59, Edw59, FS58a, Ham59, HG58, HBW^{+}59, Hay58, HF59a, HC58, KWDV58, KAWI59, Lad58, LP56, NF58, Obl58, PB58, RJ58, Rut59, SS59, STC59, Ted59, TY59a, TY59b, TY59c, Tri58, Wol58]. Structures [HHPvR56, Pit56, Reb56a, Reb56b, WCT57, BSLM59, BV59, Mos56a, Pap56b, RT59, WMK59a]. Struktur [MFW59]. Studied [Cla57, DW55b, Lus56b, Sid57, vBC55, Lat59, TY59c]. Studies [BSB59, BTD55, BTDA56, BF55, CH58, CCP57, ED55b, FmF57, Han56, Hod55, IH57, JS57, Jak56, JC57, KP55, LC56, Mat56, MC58, Mos56b, Pal55b, PP57, SD55, SR57, SR58, SW56, SJ57, Sul55, TPD56, TSW58, UNG57, WSW58, WS56a, Wen59b, WMK59b, YFJ59a, YFJ59b, YFJ59c, LDK55, CZ59b, CT59a, CT59b, Fin59, FJ58, FJY58, HN59, HC59, Joh58, MGR^{+}57, TC59, Wen59a]. Study [AP57, BB57, BG55b, BMG^{+}57, BF56, DF56a, De 56d, DW55a, FM56, FHM56, HDG57, KJTD55, KC56, LW59, MB55, MR57a, Mui57, Pit56, Ris56, Rot57a, Sel55, She56, STC57, TD56, Wei57, Woo57, dHF58, Bad59, Fin58, GT59, HA58, KD58, KRS58, Koe58a, LY58, LR58, LBVS59, MQ59, NP58, PK59a, PK59b, Rud58, Sho58, SP58c, SP58d, SP58e, SP59a, SP59b, Sto59, TG58, WG59, dHC59]. Subcellular [FSF56, TS59, PG59]. Subendothelial [Lev56b]. Submicroscopic [DR58, DB55, De 56c, De 56a, De 56d, DF57, DL58b, LD59]. Substance [DM56, DM57, Coh58]. Substances [Fin58]. Substitution [FS58b]. Substrate [WM59]. subtilis [TY59a, TY59b]. Succinate [BB57]. Succinic [WM55]. Suckling [Cla59]. Sucrose [Nov56]. Suggesting [FW57]. Suggestion [Ben56c]. Sulfate [Dzi56, DBDA57, EU57, JC57]. Sulfate-S [DBDA57, EU57]. Sulfhydryl [KD58, Ste58]. Supply [CER57]. Support [Wat55b]. Supporting [BS59]. Surface [DO58, TCS59]. Surfaces [IH57]. Suspensions [DM59]. SV [HF59b]. Swelling [FS56, LS59, LRS59, WM58]. Swine [HL57]. Sympathetic [Smi59]. Synapse [De 56d, Lad58]. Synapses [DB55, DF56b, Pal56c]. Synaptic [DF56b]. Synchronous [FIS59]. Syntheses [BG55b]. Synthesis [MKT59, GM59b, KSB59, Pre59, SP58f]. Synthesized [PM56]. Synthetic [Fin59, WM58]. System [BB57, Har56, Lus56a, Lus56b, Pal56c, vBC55, DGW58, STC59]. Systems [BP57, Fre56, Hod56a, HMM56, Tho57, Rut58].

T2 [Bac55]. Tail [Afz59]. Take [Eps58]. Tapetum [BP59b]. Taste [dL58a]. Taurine [Coh58]. Technique [AF59, ED55b, LC56, Ric59, RPST58, dH58]. Techniques [Gel57, Mos56b, Ros57]. Teeth [BVLC57]. Tegenaria [AR57]. Teleost [CD59]. Temperature [SS56]. Ten [LOK56]. Term [Hol59]. Testicular [FI58]. Testis [BF55, LY58]. Tetrahymena [ET58]. tetroxide [LDK55]. Their [Gal56, Hod56a, Pea56, Str59a, HN59]. Theoretical [BS58b]. Thin [Cos57, Gel57, MMR56, Orn56a, Pea58, SP58a, Sel55, Wat57, GL59, HA58, TO59, Wei59]. Thioacetamide [KvH59]. Third [EW59]. Three [Tho57]. Thymidine [dF59, PS58, SP58b]. Thymine [MM55]. Thymus [HL57, HL58b]. Thyroxine [LRS59]. Tip [JK56]. Tissue [Bie55, CER57, Doy55, Ger56, GMGL57, NPS56, Orn56c, RWPP55, Ros57, SL57, WK56, And56, BSLM59, CZ59b, Cau57, DW58, GM59a, Lat59, LC58, MQ59, NF58, Pau56, Pea59, RPST58, RVH58, Wat55b, Wat58a, Wat58b, WLP59]. Tissues [Bél59, DW55b, EU57, FM56, JK57, Lev56b, MM57, MPM56, Ste56, Str58, Str59a]. Toad [BF56]. Tobacco [Mat58]. Tokophrya [RP55, Rud56, Rud58]. Tongue [WM55]. Topographical [BR56a]. Topography [Gré57]. Torpedo [Luf56a]. Tourniquet [MRC56]. Tract [Dzi56]. Tradescantia [Woo56, Woo58]. Transfer [Odo56, RMFW59, MC58]. Transplantable [Bin59]. Transport [Ben56a, Pea56]. Treated [LBVS59]. Treatment [HD58]. Treatments [Mer59c]. Trichonympha [Gri59]. Trillium [Ste58]. Triphosphopyridine [NWS58a]. Tritiated [dF59]. Triturus [SS59]. Tube [SB59b]. Tubule [RMW57, Nov59, SNM58]. Tubules [BTD55]. Tumor [AP57, Bad59, FERS59, Lov57, MLM^{+}59, RSK59]. Tumors [LOK56]. Tumour [Eps57a]. Turnover [Dzi56, Koe58a, RH59]. Turtle [FS58a]. Two [ED55b, GB58, MH55, Rob56b, MKT58, MR58, Men58a, Men58b, Men58c, PG59, RT59, SB59a]. Tymbal [ERdH58b]. Type [Eps59, KP55, MGR^{+}57]. Types [ED55b, MKT58].

ultraminces [SRK59]. Ultrastructure [AR57, Fre56, HMM55, Kar56c, Kar56a, Kar56b, Rob55, Rob56a, Spi56, Kar58, PCV58, Rob57, Rob58b, SBF58, Wis57]. Ultrathin [FM56, MBMR55, GH58]. Ultraviolet [FERS59, RC57]. Ungerminated [JSS59]. Units [Gro56]. Unsegmented [PCV58]. Unusual [MGR^{+}57]. Uptake [BVLC57, Koe58b, dF59, Odo56, Str58, SB59a]. Urchin [Rot57c, Wen59b, KD58, Wen59a]. Use [DW55b, HMF57, MBR56, Rud55, WH59, BS59, DW58, HM59, Wat55b]. Using [Mos56b]. Uterus [Sho58]. Utilization [JC57, Tro57, Joh58].

Vacuole [PB58, Rud58]. Variations [Dem56, Ste58]. Variety [HBW^{+}59]. Various [Spi56, Str58, Str59a]. Varying [Cau57]. Vascular [MC58]. Vase [Rob56b]. Vase-Shaped [Rob56b]. Vehicle [Cau57]. Vertebrate [Faw56b, Rob55]. Vertebrates [Uzm56]. Vesicles [Cle56, DF56b]. Vesiculation [Ben56a]. Vessels [Kar59, TG58]. Vicia [Kih59a, Laf58a, Laf58b, Mer59c, SD58, SD55]. Villiger [HF59b]. Villus [PK59a, PK59b]. Viral [MBMR55]. Virus [AP57, BMG^{+}57, Faw56a, Hal55, Mat58, MBR56, Smi56, KWDV58, KAWI59, LP56]. Virus-Like [AP57, Faw56a]. Viruses [MHRM56]. Viscosity [KP55]. Visual [LD59]. Vital [DW55a, DW55b]. Vitelline [AR57, STC57]. Vitreous [WB55]. Vitro [BVLC57, MM57, MPM56, WL56, CZ59b, Fri59, WPE^{+}58]. Vitrosin [GMC55]. Vp [Ros57]. Vulgata [Lac57].

Wall [BCCMS57, Han57b, RMFW59, KR58]. Wallerian [Via58]. Walls [Kar56c, Gre58, SB58]. Wasp [ERdH58a]. Water [Fin57, Pea56, GH58]. Wavelength [MR58, Men58a, Men58b, Men58c]. Weights [HK56]. White [Str57a, NP58]. Within [Gri56, FW57]. Without [NPS56].

X [FmF57, SB56, BS56, Fin59, Kih59a]. X-Ray [FmF57, SB56, BS56, Fin59].

Yolk [Wis57].

Zea [HMM55, LVH58]. Zona [TCS59]. Zone [Cle56].

References

Anderson:1956:EEM

[AB56] Everett Anderson and H. W. Beams. Evidence from electron micrographs for the passage of material through pores of the nuclear membrane. Journal of Cell Biology, 2(4):439–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/439.

Ascenzi:1959:TDM

[AF59] Antonio Ascenzi and Claudine Fabry. Technique for Dissection and Measurement of Refractive Index of Osteones. Journal of Cell Biology, 6(1):139–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/1/139.

Afzelius:1959:EMS

[Afz59] Björn Afzelius. Electron Microscopy of the Sperm Tail Results Obtained with a New Fixative. Journal of Cell Biology, 5(2):269–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/2/269.

Alfert:1956:CDN

[Alf56] Max Alfert. Chemical differentiation of nuclear proteins during spermatogenesis in the Salmon. Journal of Cell Biology, 2(2):109–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/2/109.

Anderson:1956:NHT

[And56] Norman G. Anderson. A Note on “Homogenizers” for Tissue Brei Preparation. Journal of Cell Biology, 2(2): 219–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/2/219.

Ansley:1958:HMM

[Ans58] Hudson R. Ansley. Histones of Mitosis and Meiosis in Loxa flavicolis (Hemipteran). Journal of Cell Biology, 4 (1):59–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/1/59.

Adams:1957:EMS

[AP57] W. Robert Adams and Alfred M. Prince. An electron microscope study of the morphology and distribution of the intracytoplasmic “virus-like” particles of Ehrlich ascites tumor cells. Journal of Cell Biology, 3 (2):161–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/2/161.

Andre:1957:UVB

[AR57] Jean André and Charles Rouiller. The ultrastructure of the vitelline body in the oocyte of the spider tegenaria parietina. Journal of Cell Biology, 3(6):977– ??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/977.

Allen:1958:APG

[AR58] R. D. Allen and J. D. Roslansky. An Anterior–Posterior Gradient of Refractive Index in the Ameba and Its Significance in Ameboid Movement. Journal of Cell Biology, 4(5):517–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/517.

Allen:1959:CAC

[AR59] Robert D. Allen and John D. Roslansky. The consistency of ameba cytoplasm and its bearing on the mechanism of ameboid movement. Journal of Cell Biology, 6(3): 437–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/437.

Anderson:1958:EMO

[AvB58] Everett Anderson and V. L. van Breemen. Electron Microscopic Observations on Spinal Ganglion Cells of Rana pipiens after Injection of Malononitrile. Journal of Cell Biology, 4(1):83–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/1/83.

Backus:1955:API

[Bac55] Robert C. Backus. Adhesion partitioning: Intrasomatic observations on normal Escherichia coli and T2 bacteriophage. Journal of Cell Biology, 1(2):99–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/2/99.

Bader:1959:CSS

[Bad59] Saul Bader. A Cytochemical Study of the Stem Cell Concept in Specimens of a Human Ovarian Tumor. Journal of Cell Biology, 5(2):217–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
217.

Banfield:1959:DGE

[Ban59] William G. Banfield. Dense Granule in the Elementary Body of Molluscum Contagiosum. Journal of Cell Biology, 5(3):513–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/3/513.

Ball:1957:IMB

[BB57] Eric G. Ball and Russell J. Barrnett. An integrated morphological and biochemical study of a purified preparation of the succinate and dpnh oxidase system. Journal of Cell Biology, 3(6):1023–??, November 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/6/
1023.

Baradi:1959:NOA

[BB59] Abdel F. Baradi and Geoffrey H. Bourne. New Observations on the Alkaline Glycerophosphatase Reaction in the Papilla Foliata. Journal of Cell Biology, 5(1): 173–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/173.

Bessis:1957:IPN

[BBG57] Marcel C. Bessis and Janine Breton-Gorius. Iron Particles in Normal Erythroblasts and Normal and Pathological Erythrocytes. Journal of Cell Biology, 3(3): 503–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/3/503.

Bessis:1959:DADa

[BBG59a] Marcel Bessis and Janine Breton-Gorius. Differents aspects du fer dans l’organisme. Journal of Cell Biology, 6(2):231–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/231.

Bessis:1959:DADb

[BBG59b] Marcel Bessis and Janine Breton-Gorius. Differents aspects du fer dans l’organisme. Journal of Cell Biology, 6(2):237–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/237.

Bayley:1957:SPE

[BCCMS57] S. T. Bayley, J. R. Colvin, F. P. Cooper, and Cecily A. Martin-Smith. The structure of the primary epidermal cell wall of avena coleoptiles. Journal of Cell Biology, 3 (2):171–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/2/171.

Bear:1956:SCM

[Bea56] Richard S. Bear. The Structure of Collagen Molecules and Fibrils. Journal of Cell Biology, 2(3):363–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/3/363.

Belt:1958:OAC

[Bel58] W. Duane Belt. The Origin of Adrenal Cortical Mitochondria and Liposomes: a Preliminary Report. Journal of Cell Biology, 4(3):337–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/3/
337.

Belanger:1959:ASM

[Bél59] Leonard F. Bélanger. Alpharadiography: a simple method for determination of mass concentration in cells and tissues. Journal of Cell Biology, 6(2):197–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/2/197.

Bennett:1956:CMF

[Ben56a] H. Stanley Bennett. The concepts of membrane flow and membrane vesiculation as mechanisms for active transport and ion pumping. Journal of Cell Biology, 2(4):99–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/99.

Bennett:1956:SRS

[Ben56b] H. Stanley Bennett. The sarcoplasmic reticulum of striped muscle. Journal of Cell Biology, 2(4):171–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/171.

Bennett:1956:SNL

[Ben56c] H. Stanley Bennett. A suggestion as to the nature of the lysosome granules. Journal of Cell Biology, 2 (4):185–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/185.

Bernstein:1956:ISN

[Ber56] Maurice H. Bernstein. Iron as a Stain for Nucleic Acids in Electron Microscopy. Journal of Cell Biology, 2(5): 633–??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/5/633.

Burgos:1955:SFS

[BF55] Mario H. Burgos and Don W. Fawcett. Studies on the fine structure of the mammalian testis. Journal of Cell Biology, 1(4):287–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/4/287.

Burgos:1956:EMS

[BF56] Mario H. Burgos and Don W. Fawcett. An electron microscope study of spermatid differentiation in the toad, Bufo Arenarum Hensel. Journal of Cell Biology, 2(3):

223–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/3/223.

Bloch:1955:EDD

[BG55a] David P. Bloch and Gabriel C. Godman. Evidence of differences in the desoxyribonucleoprotein complex of rapidly proliferating and non-dividing cells. Journal of Cell Biology, 1(6):531–??, November 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/6/
531.

Bloch:1955:MSS

[BG55b] David P. Bloch and Gabriel C. Godman. A microphotometric study of the syntheses of desoxyribonucleic acid and nuclear histone. Journal of Cell Biology, 1(1): 17–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/1/17.

Biesele:1955:AEM

[Bie55] John J. Biesele. Antagonistic effects of 6-mercaptopurine and coenzyme a on mitochondria and mitosis in tissue culture. Journal of Cell Biology, 1(2):119–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/1/2/119.

Binggeli:1959:AIC

[Bin59] Marlene Friedlaender Binggeli. Abnormal Intranuclear and Cytoplasmic Formations Associated with a Chemically Induced, Transplantable Chicken Sarcoma. Journal of Cell Biology, 5(1):143–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/
143.

Bennett:1959:CBB

[BL59] H. Stanley Bennett and John H. Luft. s -Collidine as a Basis for Buffering Fixatives. Journal of Cell Biology, 6 (1):113–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/113.

Birbeck:1957:EMHa

[BM57a] M. S. C. Birbeck and E. H. Mercer. The electron microscopy of the human hair follicle. Journal of Cell Biology, 3(2):203–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/2/203.

Birbeck:1957:EMHb

[BM57b] M. S. C. Birbeck and E. H. Mercer. The electron microscopy of the human hair follicle. Journal of Cell Biology, 3(2):215–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/2/215.

Birbeck:1957:EMHc

[BM57c] M. S. C. Birbeck and E. H. Mercer. The electron microscopy of the human hair follicle. Journal of Cell Biology, 3(2):223–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/2/223.

Bornstein:1958:SOP

[BM58a] Murray B. Bornstein and Margaret R. Murray. Serial Observations on Patterns of Growth, Myelin Formation, Maintenance and Degeneration in Cultures of New–Born Rat and Kitten Cerebellum. Journal of Cell Biology, 4 (5):499–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/499.

Bradbury:1958:FSA

[BM58b] S. Bradbury and G. A. Meek. The Fine Structure of the Adipose Cell of the Leech Glossiphonia complanata. Journal of Cell Biology, 4(5):603–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/5/603.

Benitez:1959:HCA

[BMC59] Helena H. Benitez, Margaret R. Murray, and Erwin Chargaff. Heteromorphic Change of Adult Fibroblasts by Ribonucleoproteins. Journal of Cell Biology, 5(1):

25–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/25.

Bloch:1957:CHE

[BMG^{+}57] David P. Bloch, Councilman Morgan, Gabriel C. Godman, Calderon Howe, and Harry M. Rose. A correlated histochemical and electron microscopic study of the intranuclear crystalline aggregates of adenovirus (riapc virus) in HeLa cells. Journal of Cell Biology, 3(1): 1–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/1.

Becker:1959:CDN

[BO59] Norwin H. Becker and Kazuo Ogawa. Changes in Desoxyribose Nucleic Acid per Nucleus in Renal Compensatory Hypertrophy in the Rat. Journal of Cell Biology, 6(2):295–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/295.

Borysko:1956:RDMa

[Bor56a] Emil Borysko. Recent developments in methacrylate embedding. Journal of Cell Biology, 2(4):3–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/3.

Borysko:1956:RDMb

[Bor56b] Emil Borysko. Recent developments in methacrylate embedding. Journal of Cell Biology, 2(4):15–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/15.

Belt:1956:MSS

[BP56] W. Duane Belt and Daniel C. Pease. Mitochondrial structure in sites of steroid secretion. Journal of Cell Biology, 2(4):369–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/369.

Barrnett:1957:HDS

[BP57] Russell J. Barrnett and George E. Palade. Histochemical demonstration of the sites of activity of dehydrogenase systems with the electron microscope. Journal of Cell Biology, 3(4):577–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/4/577.

Barrnett:1959:EAM

[BP59a] Russell J. Barrnett and George E. Palade. Enzymatic Activity in the M Band. Journal of Cell Biology, 6(2): 163–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/2/163.

Bernstein:1959:EMT

[BP59b] Maurice H. Bernstein and Daniel C. Pease. Electron Microscopy of the Tapetum Lucidum of the Cat. Journal of Cell Biology, 5(1):35–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/
35.

Brandt:1959:M

[BP59c] Philip W. Brandt and George D. Pappas. Mitochondria. Journal of Cell Biology, 6(1):91–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/1/91.

Bernhard:1956:CTR

[BR56a] W. Bernhard and C. Rouiller. Close topographical relationship between mitochondria and ergastoplasm of liver cells in a definite phase of cellular activity. Journal of Cell Biology, 2(4):73–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/73.

Birbeck:1956:DIM

[BR56b] M. S. C. Birbeck and E. Reid. Development of an improved medium for the isolation of liver mitochondria. Journal of Cell Biology, 2(5):609–??, September 1956.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/5/609.

Bear:1956:SPF

[BS56] Richard S. Bear and Cecily Cannan Selby. The structure of paramyosin fibrils according to X-ray diffraction. Journal of Cell Biology, 2(1):55–??, January 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/1/
55.

Barrnett:1958:HDP

[BS58a] Russell J. Barrnett and Arnold M. Seligman. Histochemical Demonstration of Protein–Bound Alpha–Acylamido Carboxyl Groups. Journal of Cell Biology, 4(2):169–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/2/169.

Bergeron:1958:MET

[BS58b] John A. Bergeron and Marcus Singer. Metachromasy: an experimental and theoretical reevaluation. Journal of Cell Biology, 4(4):433–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/4/433.

Birbeck:1959:UNS

[BS59] M. S. C. Birbeck and K. A. Stacey. The Use of Novel Supporting Films for the Electron Microscopy of Deoxyribosenucleic Acid and Other Macromolecules. Journal of Cell Biology, 5(1):167–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/
167.

Balazs:1959:CAA

[BSB59] Endre A. Balazs, John A. Szirmai, and Gudrun Bergendahl. C^{14} assays and autoradiographic studies on the rooster comb. Journal of Cell Biology, 5(2):319–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/2/319.

Bencosme:1959:RML

[BSLM59] Sergio A. Bencosme, Robert S. Stone, Harrison Latta, and Sidney C. Madden. A Rapid Method for Localization of Tissue Structures or Lesions for Electron Microscopy. Journal of Cell Biology, 5(3):508–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/3/508.

Beams:1955:EMS

[BTD55] H. W. Beams, T. N. Tahmisian, and R. L. Devine. Electron microscope studies on the cells of the malpighian tubules of the grasshopper (Orthoptera, Acrididae). Journal of Cell Biology, 1(3):197–??, May 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/3/
197.

Beams:1956:EMS

[BTDA56] H. W. Beams, T. N. Tahmisian, R. L. Devine, and Everett Anderson. Electron microscope studies on the dictyosomes and acroblasts in the male germ cells of the Cricket. Journal of Cell Biology, 2(4):123–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/123.

Buck:1958:FSE

[Buc58] R. C. Buck. The Fine Structure of Endothelium of Large Arteries. Journal of Cell Biology, 4(2):187–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/2/187.

Brody:1959:OCS

[BV59] Marcia Brody and Albert E. Vatter. Observations on Cellular Structures of Porphyridium cruentum. Journal of Cell Biology, 5(2):289–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/

Belanger:1957:EFF

[BVLC57] L. F. Bélanger, W. J. Visek, W. E. Lotz, and C. L. Comar. The effects of fluoride feeding on the organic matrix of bones and teeth of pigs as observed by autoradiography after in vitro uptake of Ca^{45} and S^{35}. Journal of Cell Biology, 3(4):559–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/4/559.

Cameron:1956:NBG

[Cam56] D. A. Cameron. A note on breaking glass knives. Journal of Cell Biology, 2(4):57–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
57.

Carlstrom:1957:CSC

[Car57] Diego Carlström. The crystal structure of α-chitin (polyn-acetyl-d–glucosamine). Journal of Cell Biology, 3(5): 669–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/669.

Case:1959:HGC

[Cas59] Norman M. Case. Hemosiderin granules in the choroid plexus. Journal of Cell Biology, 6(3):527–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/3/527.

Caulfield:1957:EVV

[Cau57] James B. Caulfield. Effects of Varying the Vehicle for OsO_{4} in Tissue Fixation. Journal of Cell Biology, 3(5): 827–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/827.

Colwin:1957:EMS

[CCP57] Arthur L. Colwin, Laura Hunter Colwin, and Delbert E. Philpott. Electron microscope studies of early stages of sperm penetration in hydroides hexagonus (annelida) and saccoglossus kowalevskii (enteropneusta). Journal of Cell Biology, 3(3):489–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/3/489.

Copeland:1959:ABM

[CD59] D. E. Copeland and A. J. Dalton. An Association between Mitochondria and the Endoplasmic Reticulum in Cells of the Pseudobranch Gland of a Teleost. Journal of Cell Biology, 5(3):393–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/3/393.

Caesar:1957:ANS

[CER57] Rudolf Caesar, George A. Edwards, and Helmut Ruska. Architecture and nerve supply of mammalian smooth muscle tissue. Journal of Cell Biology, 3(6):867–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/867.

Clark:1958:EMS

[CH58] Wallace H. Clark and Richard G. Hibbs. Electron microscope studies of the human epidermis the clear cell of Masson (dendritic cell or melanocyte). Journal of Cell Biology, 4(6):679–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/679.

Chapman:1959:EMO

[Cha59] George B. Chapman. Electron Microscope Observations on the Behavior of the Bacterial Cytoplasmic Membrane During Cellular Division. Journal of Cell Biology, 6(2): 221–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/2/221.

Charles:1959:EMO

[CI59] Arwyn Charles and John T. Ingram. Electron Microscope Observations of the Melanocyte of the Human Epidermis. Journal of Cell Biology, 6(1):41–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/1/41.

Clark:1957:CDK

[Cla57] Sam L. Clark. Cellular differentiation in the kidneys of newborn mice studied with the electron microscope. Journal of Cell Biology, 3(3):349–??, May 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/3/
349.

Clark:1959:IPC

[Cla59] Sam L. Clark. The Ingestion of Proteins and Colloidal Materials by Columnar Absorptive Cells of the Small Intestine in Suckling Rats and Mice. Journal of Cell Biology, 5(1):41–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/1/41.

Clermont:1956:GZR

[Cle56] Yves Clermont. The Golgi zone of the rat spermatid and its role in the formation of cytoplasmic vesicles. Journal of Cell Biology, 2(4):119–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
119.

Churg:1958:SIE

[CMG58] J. Churg, W. Mautner, and E. Grishman. Silver Impregnation for Electron Microscopy. Journal of Cell Biology, 4(6):841–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/6/841.

Cohen:1955:ORH

[Coh55] Carolyn Cohen. Optical rotation and helical polypeptide chain configuration in collagen and gelatin. Journal of Cell Biology, 1(3):203–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/1/3/203.

Cohen:1956:EED

[Coh56] Adolph I. Cohen. The effect of enucleation on the dpn level of ameba. Journal of Cell Biology, 2(1):

15–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/1/15.

Cohen:1957:CSP

[Coh57a] Adolph I. Cohen. The cell in starvation: Physiological and chemical observations. Journal of Cell Biology, 3 (6):923–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/6/923.

Cohen:1957:EMO

[Coh57b] Adolph I. Cohen. Electron microscopic observations of Amoeba Proteus in growth and inanition. Journal of Cell Biology, 3(6):859–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/6/859.

Cohen:1958:LPS

[Coh58] Carolyn Cohen. The Lotmar–Picken Substance: Taurine. Journal of Cell Biology, 4(4):489–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/4/489.

Cosslett:1957:HTS

[Cos57] V. E. Cosslett. How thin should a section be? Journal of Cell Biology, 3(5):815–??, September 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/5/
815.

Chu:1956:MCI

[CS56] C. H. U. Chu and C. A. Swinyard. Morphological and cytochemical identification of the Golgi apparatus. Journal of Cell Biology, 2(3):263–??, May 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/3/
263.

Chapman:1959:CSNa

[CT59a] George B. Chapman and Lewis G. Tilney. Cytological Studies of the Nematocysts of Hydra. Journal of Cell Biology, 5(1):69–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/1/69.

Chapman:1959:CSNb

[CT59b] George B. Chapman and Lewis G. Tilney. Cytological Studies of the Nematocysts of Hydra. Journal of Cell Biology, 5(1):79–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/1/79.

Caro:1958:LDA

[CvTF58] Lucien G. Caro, Robert P. van Tubergen, and Frederick Forro. The Localization of Deoxyribonucleic Acid in Escherichia coli. Journal of Cell Biology, 4(4): 491–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/4/491.

Cascarano:1959:FIH

[CZ59a] Joseph Cascarano and Benjamin W. Zweifach. Factors Influencing the Histochemical Demonstration of Coenzyme–Dependent Dehydrogenases and Diaphorases. Journal of Cell Biology, 5(2):309–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
309.

Cascarano:1959:VSM

[CZ59b] Joseph Cascarano and Benjamin W. Zweifach. In Vitro Studies of the Metabolism of Tissue Slices. Journal of Cell Biology, 6(2):215–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/2/
215.

DeRobertis:1955:SFS

[DB55] Eduardo D. P. De Robertis and H. Stanley Bennett. Some features of the submicroscopic morphology of synapses in frog and earthworm. Journal of Cell Biology, 1(1):47–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/1/47.

Dziewiatkowski:1957:SAM

[DBDA57] Dominic D. Dziewiatkowski, Felix Bronner, Nicola Di Ferrante, and Reginald M. Archibald. Some aspects of the metabolism of sulfate-s^{35} and calcium-45 in the metaphyses of immature rats. Journal of Cell Biology, 3 (2):151–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/2/151.

DeRobertis:1956:EMOc

[De 56a] E. De Robertis. Electron microscopic observations on the submicroscopic morphology of the meiotic nucleus and chromosomes. Journal of Cell Biology, 2(6):785– ??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/6/785.

DeRobertis:1956:MRR

[De 56b] E. De Robertis. Morphogenesis of the retinal rods. Journal of Cell Biology, 2(4):209–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
209.

DeRobertis:1956:EMOb

[De 56c] Eduardo De Robertis. Electron microscope observations on the submicroscopic organization of the retinal rods. Journal of Cell Biology, 2(3):319–??, May 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/3/
319.

DeRobertis:1956:SCS

[De 56d] Eduardo De Robertis. Submicroscopic changes of the synapse after nerve section in the acoustic ganglion of the guinea pig. An electron microscope study. Journal of Cell Biology, 2(5):503–??, September 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/5/

Dempsey:1956:VSM

[Dem56] E. W. Dempsey. Variations in the structure of mitochondria. Journal of Cell Biology, 2(4):305–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/305.

Dalton:1956:CSG

[DF56a] A. J. Dalton and Marie D. Felix. A comparative study of the Golgi complex. Journal of Cell Biology, 2(4): 79–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/79.

DeRobertis:1956:EMOa

[DF56b] Eduardo De Robertis and Carlos M. Franchi. Electron microscope observations on synaptic vesicles in synapses of the retinal rods and cones. Journal of Cell Biology, 2(3):307–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/3/307.

DeRobertis:1957:SCN

[DF57] Eduardo De Robertis and Alberto Vaz Ferreira. Submicroscopic changes of the nerve endings in the adrenal medulla after stimulation of the splanchnic nerve. Journal of Cell Biology, 3(4):611–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/4/
611.

Lima-de-Faria:1959:DUT

[dF59] A. Lima de Faria. Differential uptake of tritiated thymidine into hetero- and euchromatin in Melanoplus and Secale. Journal of Cell Biology, 6(3):457–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/3/457.

DeRobertis:1958:CMM

[DGW58] Eduardo De Robertis, Hersch M. Gerschenfeld, and Flora Wald. Cellular Mechanism of Myelination in the Central Nervous System. Journal of Cell Biology, 4(5): 651–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/651.

deHarven:1958:NTC

[dH58] Étienne de Harven. A New Technique for Carbon Films. Journal of Cell Biology, 4(1):133–??, January 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/1/
133.

deHarven:1959:EMS

[dHC59] Etienne de Harven and Christian Coërs. Electron Microscope Study of the Human Neuromuscular Junction. Journal of Cell Biology, 6(1):7–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
7.

deHarven:1958:EMS

[dHF58] Étienne de Harven and Charlotte Friend. Electron microscope study of a cell-free induced leukemia of the mouse: a preliminary report. Journal of Cell Biology, 4 (2):151–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/2/151.

Dippell:1958:FSK

[Dip58] Ruth V. Dippell. The Fine Structure of Kappa in Killer Stock 51 of Paramecium aurelia. Preliminary Observations. Journal of Cell Biology, 4(1):125–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/1/125.

deLorenzo:1957:EMO

[dL57] A. J. de Lorenzo. Electron microscopic observations of the olfactory mucosa and olfactory nerve. Journal of Cell Biology, 3(6):839–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/6/839.

deLorenzo:1958:EMO

[dL58a] A. J. de Lorenzo. Electron Microscopic Observations on the Taste Buds of the Rabbit. Journal of Cell Biology, 4 (2):143–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/2/143.

DeRobertis:1958:SOR

[DL58b] Eduardo De Robertis and Arnaldo Lasansky. Submicroscopic Organization of Retinal Cones of the Rabbit. Journal of Cell Biology, 4(6):743–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/6/743.

deLamirande:1958:IDN

[dLAC58] Gaston de Lamirande, Claude Allard, and Antonio Cantero. Intracellular Distribution of 5’ Nucleotidase in Rat Liver. Journal of Cell Biology, 4(4):373–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/4/373.

deLamirande:1959:RAC

[dLAC59] Gaston de Lamirande, Claude Allard, and Antonio Cantero. Ribonucleic Acid Composition in Cytoplasmic Fractions Isolated from Rat Liver Cells. Journal of Cell Biology, 6(2):291–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/2/291.

Denues:1955:NNH

[DM55a] A. R. T. Denues and Frances C. Mottram. Note on Nucleolonemata in Human Cultured Cells. Journal of Cell Biology, 1(2):185–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/1/2/185.

Dounce:1955:FIA

[DM55b] Alexander L. Dounce and Kenneth J. Monty. Factors influencing the ability of isolated cell nuclei to form gels in dilute alkali. Journal of Cell Biology, 1(2):

155–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/2/155.

Deitch:1956:NSL

[DM56] Arline D. Deitch and Margaret R. Murray. The nissl substance of living and fixed spinal ganglion cells. Journal of Cell Biology, 2(4):433–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/433.

Deitch:1957:NSL

[DM57] Arline D. Deitch and Montrose J. Moses. The nissl substance of living and fixed spinal ganglion cells. Journal of Cell Biology, 3(3):449–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/3/449.

Danon:1959:MAC

[DM59] David Danon and Yehuda Marikovsky. A Method for Avoiding Centrifugation in the Imbedding of Suspensions in Polymethacrylates. Journal of Cell Biology, 6 (2):298–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/2/298.

Dornfeld:1958:SRC

[DO58] Ernst J. Dornfeld and Alfred Owczarzak. Surface Responses in Cultured Fibroblasts Elicited by Ethylenediaminetetraacetic Acid. Journal of Cell Biology, 4(3): 243–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/243.

Dounce:1957:AMD

[DOMtTAoJCWW57] Alexander L. Dounce, Marguerite P. O’Connell, Kenneth J. Monty, and With the Technical Assistance of Joan C. W. West. Action of mitochondrial DNA ase i in destroying the capacity of isolated cell nuclei to form gels. Journal of Cell Biology, 3(5):649–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/5/649.

Doyle:1955:PDL

[Doy55] William L. Doyle. Peptidase distribution in lymphatic tissue. Journal of Cell Biology, 1(3):221–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/3/221.

Dass:1958:SON

[DR58] C. M. S. Dass and Hans Ris. Submicroscopic organization of the nucleus during spermiogenesis in the grasshopper. Journal of Cell Biology, 4(1):129–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/1/129.

DeRobertis:1958:MCA

[DS58] E. De Robertis and D. Sabatini. Mitochondrial Changes in the Adrenocortex of Normal Hamsters. Journal of Cell Biology, 4(5):667–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
667.

Dempsey:1955:EMS

[DW55a] Edward W. Dempsey and George B. Wislocki. An electron microscopic study of the blood-brain barrier in the rat, employing Silver nitrate as a vital stain. Journal of Cell Biology, 1(3):245–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/1/3/245.

Dempsey:1955:USN

[DW55b] Edward W. Dempsey and George B. Wislocki. The use of Silver nitrate as a vital stain, and its distribution in several mammalian tissues as studied with the electron microscope. Journal of Cell Biology, 1(2): 111–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/2/111.

Dempsey:1956:EMO

[DW56] E. W. Dempsey and G. B. Wislocki. Electron microscopic observations on the placenta of the cat. Journal of Cell Biology, 2(6):743–??, November 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/6/
743.

Davies:1958:UCF

[DW58] Maurice C. Davies and R. Wallace. The Use of Carbon Films in the Culture of Tissue Cells for Electron Microscopy. Journal of Cell Biology, 4(2):231–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/2/231.

Dounce:1955:MII

[DWM^{+}55] Alexander L. Dounce, Robert F. Witter, Kenneth J. Monty, Sidney Pate, and Mary A. Cottone. A method for isolating intact mitochondria and nuclei from the same homogenate, and the influence of mitochondrial destruction on the properties of cell nuclei. Journal of Cell Biology, 1(2):139–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/2/139.

Dziewiatkowski:1956:TSM

[Dzi56] Dominic D. Dziewiatkowski. Turnover of s^{35}-sulfate in the mucosa of the gastrointestinal tract of rats as seen in autoradiograms. Journal of Cell Biology, 2(1): 29–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/1/29.

Emery:1955:IDA

[ED55a] Arthur J. Emery and Alexander L. Dounce. Intracellular distribution of alkaline phosphatase in rat liver cells. Journal of Cell Biology, 1(4):315–??, July 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/4/
315.

Emery:1955:STT

[ED55b] Arthur J. Emery and Alexander L. Dounce. Studies of two types of alkaline phosphatase in nuclei isolated from the livers of Fed and fasted rats by a modification of the Behrens technique. Journal of Cell Biology, 1(4):331–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/4/331.

Edwards:1959:FSM

[Edw59] George A. Edwards. The Fine Structure of a Multiterminal Innervation of an Insect Muscle. Journal of Cell Biology, 5(2):241–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/2/241.

Elbers:1959:IGH

[Elb59] P. F. Elbers. An Improved Grid–Holder for the Siemens Electron Microscope. Journal of Cell Biology, 6(1): 114–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/114.

Essner:1958:ANA

[ENM58] Edward Essner, Alex B. Novikoff, and Bertha Masek. Adenosinetriphosphatase and 5-Nucleotidase Activities in the Plasma Membrane of Liver Cells as Revealed by Electron Microscopy. Journal of Cell Biology, 4(6):711– ??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/711.

Epstein:1957:FSO

[Eps57a] M. A. Epstein. The fine structural organisation of rous tumour cells. Journal of Cell Biology, 3(6):851– ??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/851.

Epstein:1957:FSC

[Eps57b] M. A. Epstein. The fine structure of the cells in mouse sarcoma 37 ascitic fluids. Journal of Cell Biology, 3 (4):567–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/4/567.

Epstein:1958:DKH

[Eps58] M. A. Epstein. A Diamond Knife Holder for Microtomes Designed to Take Glass Knives. Journal of Cell Biology, 4(5):539–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/539.

Epstein:1959:OFS

[Eps59] M. A. Epstein. Observations on the fine structure of type 5 adenovirus. Journal of Cell Biology, 6(3):523– ??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/523.

Edwards:1958:EMP

[ERdH58a] George A. Edwards, Helmut Ruska, and Étienne de Harven. Electron Microscopy of Peripheral Nerves and Neuromuscular Junctions in the Wasp Leg. Journal of Cell Biology, 4(1):107–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/1/107.

Edwards:1958:NJF

[ERdH58b] George A. Edwards, Helmut Ruska, and Étienne de Harven. Neuromuscular Junctions in Flight and Tymbal Muscles of the Cicada. Journal of Cell Biology, 4(3): 251–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/251.

Edwards:1956:CCS

[ERdSSVF56] George A. Edwards, Helmut Ruska, Persio de Souza Santos, and Aristides Vallejo-Freire. Comparative cytophysiology of striated muscle with special reference to the role of the endoplasmic reticulum. Journal of Cell Biology, 2(4):143–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/143.

Elliott:1958:FSP

[ET58] Alfred M. Elliott and John W. Tremor. The fine structure of the pellicle in the contact area of conjugating Tetrahymena pyriformis. Journal of Cell Biology, 4(6): 839–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/839.

Eichel:1957:ILE

[EtTAoJH57] Herbert J. Eichel and With the Technical Assistance of Jeanne Hagan. Intracellular localization of enzymes in spleen. Journal of Cell Biology, 3(3):397–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/3/397.

Edwards:1957:ADT

[EU57] Leon C. Edwards and K. N. Udupa. Autoradiographic determination of s^{35} in tissues after injection of methionine-s^{35} and sodium sulfate-s^{35}. Journal of Cell Biology, 3(5):757–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/5/
757.

Eakin:1959:FSR

[EW59] Richard M. Eakin and Jane A. Westfall. Fine Structure of the Retina in the Reptilian Third Eye. Journal of Cell Biology, 6(1):133–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/1/133.

Fawcett:1956:EMO

[Faw56a] Don W. Fawcett. Electron microscope observations on intracellular virus-like particles associated with the cells of the LuckÉ renal adenocarcinoma. Journal of Cell Biology, 2(6):725–??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/6/725.

Fawcett:1956:FSC

[Faw56b] Don W. Fawcett. The fine structure of chromosomes in the meiotic prophase of vertebrate spermatocytes. Journal of Cell Biology, 2(4):403–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/

Felix:1956:CMC

[FD56] Marie D. Felix and A. J. Dalton. A comparison of mesothelial cells and macrophages in mice after the intraperitoneal inoculation of melanin granules. Journal of Cell Biology, 2(4):109–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/109.

Feigin:1956:MHD

[Fei56] Irwin Feigin. A Method for the Histochemical Differentiation of Cholesterol and Its Esters. Journal of Cell Biology, 2(2):213–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/2/213.

Freed:1959:URE

[FERS59] Jerome J. Freed, James L. Engle, George T. Rudkin, and Jack Schultz. Ultraviolet Radiation Effects on Ehrlich Ascites Tumor Cells. Journal of Cell Biology, 5(2): 205–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/2/205.

Frenkel:1959:SPA

[FH59] Albert W. Frenkel and Donald D. Hickman. Structure and Photochemical Activity of Chlorophyll–Containing Particles from Rhodospirillum rubrum. Journal of Cell Biology, 6(2):285–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/2/285.

Finck:1956:ISD

[FHM56] Henry Finck, Howard Holtzer, and John M. Marshall. An immunochemical study of the distribution of myosin in glycerol extracted muscle. Journal of Cell Biology, 2 (4):175–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/175.

Fawcett:1958:OCM

[FI58] Don W. Fawcett and Susumu Ito. Observations on the Cytoplasmic Membranes of Testicular Cells, Examined by Phase Contrast and Electron Microscopy. Journal of Cell Biology, 4(2):135–??, March 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/2/
135.

Finean:1957:RWS

[Fin57] J. B. Finean. The role of water in the structure of peripheral nerve myelin. Journal of Cell Biology, 3(1): 95–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/95.

Finck:1958:EMS

[Fin58] Henry Finck. An Electron Microscope Study of Basophile Substances of Frozen–Dried Rat Liver. Journal of Cell Biology, 4(3):291–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/3/291.

Finean:1959:EMX

[Fin59] J. B. Finean. Electron Microscope and X-ray Diffraction Studies of a Saturated Synthetic Phospholipide. Journal of Cell Biology, 6(1):123–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
123.

Fawcett:1959:OIB

[FIS59] Don W. Fawcett, Susumu Ito, and David Slautterback. The Occurrence of Intercellular Bridges in Groups of Cells Exhibiting Synchronous Differentiation. Journal of Cell Biology, 5(3):453–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/3/453.

Fitz-James:1958:CCS

[FJ58] Philip C. Fitz-James. Cytological and Chemical Studies of the Growth of Protoplasts of Bacillus megaterium. Journal of Cell Biology, 4(3):257–??, May 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/3/

Fitz-James:1958:MCS

[FJY58] Philip C. Fitz-James and I. Elizabeth Young. Morphological and Chemical Studies of the Spores and Parasporal Bodies of Bacillus laterosporus. Journal of Cell Biology, 4(5):639–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/5/639.

Fukuda:1959:CLI

[FK59] Tetsuo Fukuda and George B. Koelle. The Cytological Localization of Intracellular Neuronal Acetylcholinesterase. Journal of Cell Biology, 5(3):433–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/3/433.

Fernandez-Moran:1956:ADK

[FM56] H. Fernández-Morán. Applications of a diamond knife for ultrathin sectioning to the study of the fine structure of biological tissues and metals. Journal of Cell Biology, 2(4):29–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/29.

Finean:1957:EIS

[FM57] J. B. Finean and P. F. Millington. Effects of ionic strength of immersion medium on the structure of peripheral nerve myelin. Journal of Cell Biology, 3(1): 89–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/89.

Fernandez-moran:1957:EML

[FmF57] H. Fernández-morán and J. B. Finean. Electron microscope and low-angle x-ray diffraction studies of the nerve myelin sheath. Journal of Cell Biology, 3(5):725– ??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/725.

Freeman:1956:UDM

[Fre56] James A. Freeman. The ultrastructure of the double membrane systems of mitochondria. Journal of Cell Biology, 2(4):353–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/353.

Friedmann:1959:EMO

[Fri59] I. Friedmann. Electron Microscope Observations on in Vitro Cultures of the Isolated Fowl Embryo Otocyst. Journal of Cell Biology, 5(2):263–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
263.

Fonnesu:1956:OPM

[FS56] Alberto Fonnesu and Clara Severi. Oxidative phosphorylation in mitochondria from livers showing cloudy swelling. Journal of Cell Biology, 2(3):293–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/3/293.

Fawcett:1958:OFS

[FS58a] Don W. Fawcett and Cecily C. Selby. Observations on the Fine Structure of the Turtle Atrium. Journal of Cell Biology, 4(1):63–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/1/63.

Feder:1958:MPF

[FS58b] Ned Feder and Richard L. Sidman. Methods and Principles of Fixation by Freeze–Substitution. Journal of Cell Biology, 4(5):593–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
593.

Foster:1959:ASD

[FS59] T. S. Foster and Herbert Stern. The Accumulation of Soluble Deoxyribosidic Compounds in Relation to Nuclear Division in Anthers of Lilium longiflorum. Journal of Cell Biology, 5(2):187–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/

Fiala:1956:ACA

[FSF56] Silvio Fiala, Edith E. Sproul, and Anna Fiala. The action of corticotropin (acth) on nucleic acids and subcellular elements of the adrenal cortex. Journal of Cell Biology, 2(2):115–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/2/115.

Farquhar:1957:EME

[FW57] Marilyn G. Farquhar and S. Robert Wellings. Electron Microscopic Evidence Suggesting Secretory Granule Formation within the Golgi Apparatus. Journal of Cell Biology, 3(2):319–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/2/319.

Gall:1956:SGA

[Gal56] Joseph G. Gall. Small granules in the amphibian oocyte nucleus and their relationship to RNA. Journal of Cell Biology, 2(4):393–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/393.

Gall:1959:MDC

[Gal59a] Joseph G. Gall. Macronuclear Duplication in the Ciliated Protozoan Euplotes. Journal of Cell Biology, 5(2): 295–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/2/295.

Gall:1959:NEA

[Gal59b] Joseph G. Gall. The nuclear envelope after KMnO_{4} fixation. Journal of Cell Biology, 6(1):115–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/1/115.

Gay:1956:CNM

[Gay56] Helen Gay. Chromosome-nuclear membrane-cytoplasmic interrelations in Drosophila. Journal of Cell Biology, 2 (4):407–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/407.

Gibbons:1957:FSN

[GB57] I. R. Gibbons and J. R. G. Bradfield. The fine structure of nuclei during sperm maturation in the locust. Journal of Cell Biology, 3(2):133–??, March 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/2/
133.

Gall:1958:SNT

[GB58] Joseph G. Gall and Lois B. Bjork. The spermatid nucleus in two species of grasshopper. Journal of Cell Biology, 4(4):479–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/4/479.

Gatenby:1959:SLH

[GD59] J. Brontë Gatenby and A. J. Dalton. Spermiogenesis in Lumbricus herculeus. Journal of Cell Biology, 6(1): 45–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/45.

Gelber:1957:TSD

[Gel57] Dorothy Gelber. Thin sectioning: Details of techniques. Journal of Cell Biology, 3(2):311–??, March 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/2/
311.

Gersh:1956:PFD

[Ger56] Isidore Gersh. The preparation of frozen-dried tissue for electron microscopy. Journal of Cell Biology, 2 (4):37–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/37.

Glauert:1958:AEM

[GG58] Audrey M. Glauert and R. H. Glauert. Araldite as an Embedding Medium for Electron Microscopy. Journal of Cell Biology, 4(2):191–??, March 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/2/

Grigg:1958:DIM

[GH58] G. W. Grigg and H. Hoffman. A Double Imbedding Method Giving Water–Permeable Ultrathin Sections for Electron Microscopy. Journal of Cell Biology, 4(3): 331–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/331.

Glauert:1959:MCC

[GH59] Audrey M. Glauert and D. A. Hopwood. A membranous component of the cytoplasm in Streptomyces coelicolor. Journal of Cell Biology, 6(3):515–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/3/515.

Greider:1956:EMN

[GKF56] Marie H. Greider, Wencel J. Kostir, and Walter J. Frajola. Electron microscopy of the nuclear membrane of Amoeba Proteus. Journal of Cell Biology, 2(4): 445–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/445.

Gavin:1959:KHS

[GL59] Mary Ann Gavin and Bolivar J. Lloyd. Knives of High Silica Content Glass for Thin–Sectioning. Journal of Cell Biology, 5(3):507–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/3/507.

Goldstein:1959:NCR

[GM59a] Lester Goldstein and Julie Micou. Nuclear-cytoplasmic Relationships in Human Cells in Tissue Culture. Journal of Cell Biology, 6(1):1–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
1.

Goldstein:1959:PSN

[GM59b] Lester Goldstein and Julie Micou. On the Primary Site of Nuclear RNA Synthesis. Journal of Cell Biology, 6(2):

301–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/2/301.

Gregg:1959:RNM

[GM59c] Michael B. Gregg and Councilman Morgan. Reduplication of nuclear membranes in HeLa cells infected with adenovirus. Journal of Cell Biology, 6(3):539– ??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/539.

Gross:1955:VMC

[GMC55] Jerome Gross, A. Gedeon Matoltsy, and Carolyn Cohen. Vitrosin: a member of the collagen class. Journal of Cell Biology, 1(3):215–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/3/215.

Grossfeld:1957:MPT

[GMGL57] Henry Grossfeld, Karl Meyer, Gabriel Godman, and Alfred Linker. Mucopolysaccharides produced in tissue culture. Journal of Cell Biology, 3(3):391–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/3/391.

Goldsmith:1957:MCE

[GP57] Timothy H. Goldsmith and Delbert E. Philpott. The microstructure of the compound eyes of insects. Journal of Cell Biology, 3(3):429–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/3/429.

Gray:1959:EMN

[Gra59] E. G. Gray. Electron Microscopy of Neuroglial Fibrils of the Cerebral Cortex. Journal of Cell Biology, 6(1): 121–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/121.

Gregoire:1957:TOC

[Gré57] Charles Grégoire. Topography of the organic components in mother-of-pearl. Journal of Cell Biology, 3(5): 797–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/797.

Green:1958:SCD

[Gre58] Paul B. Green. Structural Characteristics of Developing Nitella Internodal Cell Walls. Journal of Cell Biology, 4(5):505–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/505.

Griesemer:1956:CEA

[Gri56] Robert D. Griesemer. Change in enzyme activity and ribonucleic acid concentration within the epidermal cell of the rat during the growth stage of the hair cycle. Journal of Cell Biology, 2(5):523–??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/5/523.

Grimstone:1959:CMN

[Gri59] A. V. Grimstone. Cytoplasmic Membranes and the Nuclear Membrane in the Flagellate Trichonympha. Journal of Cell Biology, 6(3):369–??, December 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/3/
369.

Gross:1956:BCU

[Gro56] Jerome Gross. The behavior of collagen units as a model in morphogenesis. Journal of Cell Biology, 2 (4):261–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/261.

Gallop:1957:SPP

[GSM57] Paul M. Gallop, Sam Seifter, and Edward Meilman. The solubility and properties of a purified ichthyocol in salt solutions of neutral pH. Journal of Cell Biology, 3 (4):545–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/4/545.

Gresson:1959:LEM

[GT59] R. A. R. Gresson and L. T. Threadgold. A Light and Electron Microscope Study of the Epithelial Cells of the Gut of Fasciola hepatica L. Journal of Cell Biology, 6(2): 157–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/2/157.

Guttman:1957:ANR

[Gut57] Ruth Guttman. Alterations in Nuclear Ribonucleic Acid Metabolism Induced by Kinetin. Journal of Cell Biology, 3(1):129–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/1/129.

Howatson:1958:MSC

[HA58] Allan F. Howatson and June D. Almeida. A Method for the Study of Cultured Cells by Thin Sectioning and Electron Microscopy. Journal of Cell Biology, 4(1): 115–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/1/115.

Hall:1955:EDS

[Hal55] Cecil E. Hall. Electron densitometry of stained virus particles. Journal of Cell Biology, 1(1):1–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/1/1.

Hall:1956:MOM

[Hal56] Cecil E. Hall. Method for the Observation of Macromolecules with the Electron Microscope Illustrated with Micrographs of DNA. Journal of Cell Biology, 2(5):625– ??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/5/625.

Hama:1959:SOF

[Ham59] Kiyoshi Hama. Some Observations on the Fine Structure of the Giant Nerve Fibers of the Earthworm, Eisenia foetida. Journal of Cell Biology, 6(1):61–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/1/61.

Hanson:1956:SCS

[Han56] Jean Hanson. Studies on the cross-striation of the indirect flight myofibrils of the blowfly Calliphora. Journal of Cell Biology, 2(6):691–??, November 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/6/
691.

Hannay:1957:PBB

[Han57a] Christopher L. Hannay. The parasporal body of bacillus laterosporus laubach. Journal of Cell Biology, 3(6): 1001–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/1001.

Hanson:1957:SSM

[Han57b] Jean Hanson. The structure of the smooth muscle fibres in the body wall of the earthworm. Journal of Cell Biology, 3(1):111–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/1/111.

Hartmann:1956:EMM

[Har56] J. Francis Hartmann. Electron microscopy of mitochondria in the central nervous system. Journal of Cell Biology, 2(4):375–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/375.

Hay:1958:FSB

[Hay58] Elizabeth D. Hay. The Fine Structure of Blastema Cells and Differentiating Cartilage Cells in Regenerating Limbs of Amblystoma Larvae. Journal of Cell Biology, 4(5):583–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/583.

Hamilton:1959:SSD

[HBW^{+}59] L. D. Hamilton, R. K. Barclay, M. H. F. Wilkins, G. L. Brown, H. R. Wilson, D. A. Marvin, H. Ephrussi-Taylor, and N. S. Simmons. Similarity of the Structure of DNA from a Variety of Sources. Journal of Cell Biology, 5 (3):397–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/3/397.

Hodge:1958:SOF

[HC58] Mary H. Hodge and George B. Chapman. Some Observations on the Fine Structure of the Sinus Gland of a Land Crab, Gecarcinus lateralis. Journal of Cell Biology, 4(5):571–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/571.

Hibbs:1959:EMS

[HC59] Richard G. Hibbs and Wallace H. Clark. Electron Microscope Studies of the Human Epidermis. Journal of Cell Biology, 6(1):71–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/1/71.

Hungerford:1958:CEP

[HD58] David A. Hungerford and Marie DiBerardino. Cytological Effects of Prefixation Treatment. Journal of Cell Biology, 4(4):391–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/4/391.

Hendler:1957:CSA

[HDG57] Richard W. Hendler, Albert J. Dalton, and George G. Glenner. A cytological study of the albumin-secreting cells of the hen oviduct. Journal of Cell Biology, 3 (3):325–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/3/325.

Hess:1958:FSN

[Hes58] Arthur Hess. The fine structure of nerve cells and fibers, neuroglia, and sheaths of the ganglion chain in the cockroach (Periplaneta americana). Journal of Cell Biology, 4(6):731–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/6/731.

Hickman:1959:SRR

[HF59a] Donald D. Hickman and Albert W. Frenkel. The Structure of Rhodospirillum rubrum. Journal of Cell Biology, 6(2):277–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/277.

Howling:1959:NSE

[HF59b] D. H. Howling and P. J. Fitzgerald. The Nature, Significance, and Evaluation of the Schwarzschild–Villiger (SV) Effect in Photometric Procedures. Journal of Cell Biology, 6(3):313–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/3/313.

Hamilton:1958:FSK

[HG58] L. D. Hamilton and M. E. Gettner. Fine Structure of Kappa in Paramecium aurelia. Journal of Cell Biology, 4(1):122–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/1/122.

Harford:1957:EMH

[HHP57] Carl G. Harford, Alice Hamlin, and Esther Parker. Electron microscopy of HeLa cells after the ingestion of colloidal gold. Journal of Cell Biology, 3(5):749– ??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/749.

Harford:1956:GSC

[HHPvR56] Carl G. Harford, Alice Hamlin, Esther Parker, and Theodore van Ravenswaay. Globoid structures in the cytoplasm of rapidly growing HeLa cells. Journal of Cell Biology, 2(4):347–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/347.

Hiraoka:1957:FNR

[Hir57] Tosisuke Hiraoka. Feulgen nucleal reaction. Journal of Cell Biology, 3(4):525–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/4/525.

Hale:1956:CND

[HK56] Arthur J. Hale and Ernest R. M. Kay. A comparison of nuclear dry weights determined by chemical and by interferometric methods. Journal of Cell Biology, 2(2): 147–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/2/147.

Hess:1956:CFL

[HL56] Eugene L. Hess and Saima E. Lagg. The chemical fractionation of lymphatic organs. Journal of Cell Biology, 2(5):557–??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/5/557.

Hess:1957:CFR

[HL57] Eugene L. Hess and Saima E. Lagg. The Chemical Fractionation of Rabbit and Swine Thymus. Journal of Cell Biology, 3(5):812–??, September 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/5/
812.

Hall:1958:MFD

[HL58a] Cecil E. Hall and Michael Litt. Morphological Features of DNA Macromolecules as Seen with the Electron Microscope. Journal of Cell Biology, 4(1):1–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/1/1.

Hess:1958:CTC

[HL58b] Eugene L. Hess and Saima E. Lagg. Calf Thymus Composition. A Comparison of Differential Centrifugation and Chemical Fractionation Procedures. Journal of Cell Biology, 4(6):717–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/717.

Herrmann:1959:ADE

[HL59] Heinz Herrmann and David S. Love. An Autoradiographic Demonstration of the Effect of the Corneal Epithelium on Amino Acid Incorporation into the Insoluble Constituents of the Corneal Stroma. Journal of Cell Biology, 6(1):135–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/1/135.

Harris:1959:UMB

[HM59] Patricia Harris and Daniel Mazia. The Use of Mercuric Bromphenol Blue as a Stain for Electron Microscopy. Journal of Cell Biology, 5(2):343–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
343.

Holtzer:1957:AMU

[HMF57] Howard Holtzer, John M. Marshall, and Henry Finck. An analysis of myogenesis by the use of fluorescent antimyosin. Journal of Cell Biology, 3(5):705–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/5/705.

Hodge:1955:ULG

[HMM55] A. J. Hodge, J. D. McLean, and F. V. Mercer. Ultrastructure of the lamellae and grana in the chloroplasts of Zea mays L. Journal of Cell Biology, 1(6): 605–??, November 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/6/605.

Hodge:1956:PMM

[HMM56] A. J. Hodge, J. D. McLean, and F. V. Mercer. A possible mechanism for the morphogenesis of lamellar systems in plant cells. Journal of Cell Biology, 2(5):597– ??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/5/597.

Hodge:1957:SSC

[HMM57] A. J. Hodge, E. M. Martin, and R. K. Morton. The structure of some cytoplasmic components of plant cells in relation to the biochemical properties of isolated particles. Journal of Cell Biology, 3(1):61–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/1/61.

Hannibal:1959:FSL

[HN59] Mark J. Hannibal and Marvin M. Nachlas. Further Studies on the Lyo and Desmo Components of Several Hydrolytic Enzymes and Their Histochemical Significance. Journal of Cell Biology, 5(2):279–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
279.

Hall:1959:EMO

[HNS59] C. E. Hall, A. Nisonoff, and H. S. Slayter. Electron Microscopic Observations of Rabbit Antibodies. Journal of Cell Biology, 6(3):407–??, December 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/3/
407.

Hodge:1955:SSM

[Hod55] A. J. Hodge. Studies on the structure of muscle. Journal of Cell Biology, 1(4):361–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/4/
361.

Hodge:1956:EPE

[Hod56a] A. J. Hodge. Effects of the physical environment on some lipoprotein layer systems and observations on their morphogenesis. Journal of Cell Biology, 2(4):221–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/221.

Hodge:1956:FSS

[Hod56b] A. J. Hodge. The fine structure of striated muscle. Journal of Cell Biology, 2(4):131–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
131.

Holmes:1959:LTC

[Hol59] Richard Holmes. Long-term cultivation of human cells (Chang) in chemically defined medium and effect of added peptone fractions. Journal of Cell Biology, 6(3): 535–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/535.

Howatson:1956:SPB

[How56] Allan F. Howatson. The structure of pigeon breast muscle mitochondria. Journal of Cell Biology, 2(4): 363–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/363.

Hall:1959:FMS

[HS59] Cecil E. Hall and Henry S. Slayter. The Fibrinogen Molecule: Its Size, Shape, and Mode of Polymerization. Journal of Cell Biology, 5(1):11–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/
11.

Hartman:1959:EMH

[HSH^{+}59] Richard E. Hartman, Robert B. W. Smith, Roberta S. Hartman, Charles E. Butterworth, and Jack M. Molesworth. The Electron Microscopy of Human Intestinal Epithelium Obtained with the Crosby Intestinal Biopsy Capsule. Journal of Cell Biology, 5(1): 171–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/171.

Hess:1958:CLO

[HSP58] R. Hess, D. G. Scarpelli, and A. G. E. Pearse. The Cytochemical Localization of Oxidative Enzymes. Journal of Cell Biology, 4(6):753–??, November 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/6/
753.

Huxley:1957:DAF

[Hux57] H. E. Huxley. The double array of filaments in crossstriated muscle. Journal of Cell Biology, 3(5):631– ??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/631.

Inoue:1957:SDL

[IH57] Shinya Inoué and W. Lewis Hyde. Studies on depolarization of light at microscope lens surfaces. Journal of Cell Biology, 3(6):831–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/6/831.

Jakus:1956:SCI

[Jak56] Marie A. Jakus. Studies on the cornea. II. The fine structure of Descemet’s membrane. Journal of Cell Biology, 2(4):243–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/243.

Johnston:1957:ASU

[JC57] Perry M. Johnston and Cyril L. Comar. Autoradiographic studies of the utilization of s^{35}-sulfate by the chick embryo. Journal of Cell Biology, 3(2):231–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/2/231.

Jensen:1956:CLA

[JK56] William A. Jensen and Leroy G. Kavaljian. The cytochemical localization of ascorbic acid in root tip cells. Journal of Cell Biology, 2(1):87–??, January 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/1/

Jacobson:1957:DEC

[JK57] K. Bruce Jacobson and Nathan O. Kaplan. Distribution of enzymes cleaving pyridine nucleotides in animal tissues. Journal of Cell Biology, 3(1):31–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/1/31.

Johnston:1958:ASU

[Joh58] Perry M. Johnston. Autoradiographic Studies of the Utilization of Ca^{45} by the Chick Embryo. Journal of Cell Biology, 4(2):163–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/2/163.

Jackson:1957:SBC

[JS57] Sylvia Fitton Jackson and R. H. Smith. Studies on the biosynthesis of collagen. Journal of Cell Biology, 3(6): 897–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/897.

Johnston:1959:INU

[JSS59] F. B. Johnston, G. Setterfield, and Herbert Stern. The Isolation of Nucleoli from Ungerminated Pea Embryos. Journal of Cell Biology, 6(1):53–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
53.

Karrer:1956:UMLb

[Kar56a] H. E. Karrer. The ultrastructure of mouse lung. Journal of Cell Biology, 2(4):115–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
115.

Karrer:1956:UMLc

[Kar56b] H. E. Karrer. The ultrastructure of mouse lung. Journal of Cell Biology, 2(4):287–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/

Karrer:1956:UMLa

[Kar56c] H. E. Karrer. The ultrastructure of mouse lung general architecture of capillary and alveolar walls. Journal of Cell Biology, 2(3):241–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/3/241.

Karrer:1958:UML

[Kar58] H. E. Karrer. The Ultrastructure of Mouse Lung: The Alveolar Macrophage. Journal of Cell Biology, 4(6): 693–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/693.

Karrer:1959:SMB

[Kar59] H. E. Karrer. The Striated Musculature of Blood Vessels. Journal of Cell Biology, 6(3):383–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/3/383.

Kallman:1959:FSC

[KAWI59] Frances Kallman, John M. Adams, Robley C. Williams, and David T. Imagawa. Fine Structure of Cellular Inclusions in Measles Virus Infections. Journal of Cell Biology, 6(3):379–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/3/379.

Knappeis:1956:EMS

[KC56] G. G. Knappeis and F. Carlsen. Electron microscopical study of skeletal muscle during isotonic (afterload) and isometric contraction. Journal of Cell Biology, 2 (2):201–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/2/201.

Kuwabara:1959:ELP

[KC59] Toichiro Kuwabara and David G. Cogan. Experimental lipogenesis by Purkinje fibers. Journal of Cell Biology, 6(3):519–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/3/519.

Kaufmann:1956:FSC

[KD56] Berwind P. Kaufmann and Deepesh N. De. Fine structure of chromosomes. Journal of Cell Biology, 2(4): 419–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/419.

Kawamura:1958:CSS

[KD58] Naoko Kawamura and Katsuma Dan. A Cytochemical Study of the Sulfhydryl Groups of Sea Urchin Eggs during the First Cleavage. Journal of Cell Biology, 4(5): 615–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/615.

Kemp:1956:DCC

[Kem56a] Norman E. Kemp. Differentiation of the cortical cytoplasm and inclusions in oocytes of the frog. Journal of Cell Biology, 2(4):187–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/187.

Kemp:1956:EMG

[Kem56b] Norman E. Kemp. Electron microscopy of growing oocytes of Rana Pipiens. Journal of Cell Biology, 2 (3):281–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/3/281.

Kleinfeld:1956:EMI

[KGF56] Ruth G. Kleinfeld, Marie H. Greider, and Walter J. Frajola. Electron microscopy of intranuclear inclusions found in human and rat liver parenchymal cells. Journal of Cell Biology, 2(4):435–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
435.

Kuff:1956:CBE

[KHD56] Edward L. Kuff, George H. Hogeboom, and Albert J. Dalton. Centrifugal, biochemical, and electron microscopic analysis of cytoplasmic particulates in liver homogenates. Journal of Cell Biology, 2(1):33–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/1/33.

Kihlman:1956:FAP

[Kih56] B. A. Kihlman. Factors affecting the production of chromosome aberrations by chemicals. Journal of Cell Biology, 2(5):543–??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/5/543.

Kihlman:1957:EICa

[Kih57] B. A. Kihlman. Experimentally induced chromosome aberrations in plants. Journal of Cell Biology, 3(3): 363–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/3/363.

Kihlman:1959:ERI

[Kih59a] B. A. Kihlman. The Effect of Respiratory Inhibitors and Chelating Agents on the Frequencies of Chromosomal Aberrations Produced by X–Rays in Vicia. Journal of Cell Biology, 5(3):479–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/3/479.

Kihlman:1959:REC

[Kih59b] B. A. Kihlman. On the Radiomimetic Effects of Cupferron and Potassium Cyanide. Journal of Cell Biology, 5 (2):351–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/2/351.

Kisch:1956:SH

[Kis56] Bruno Kisch. The sarcosomes of the heart. Journal of Cell Biology, 2(4):361–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/361.

Karnovsky:1955:CHS

[KJTD55] Manfred L. Karnovsky, Susan Skilling Jeffrey, Michaela Smits Thompson, and Helen Wendler Deane. A chemical and histochemical study of the lipides of the pyloric cecum of the starfish, asterias forbesi. Journal of Cell Biology, 1(2):173–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/2/173.

Kihlman:1957:EICb

[KMS57] B. A. Kihlman, T. Merz, and C. P. Swanson. Experimentally induced chromosome aberrations in plants. Journal of Cell Biology, 3(3):381–??, May 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/3/
381.

Koechlin:1955:CCA

[Koe55] Bernard A. Koechlin. On the chemical composition of the axoplasm of squid giant nerve fibers with particular reference to its ion pattern. Journal of Cell Biology, 1 (6):511–??, November 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/6/511.

Koenig:1958:ASN

[Koe58a] Harold Koenig. An Autoradiographic Study of Nucleic Acid and Protein Turnover in the Mammalian Neuraxis. Journal of Cell Biology, 4(6):785–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/6/785.

Koenig:1958:UAC

[Koe58b] Harold Koenig. Uptake of adenine-8-C^{14} and orotic-6-C^{14} acid into nuclear DNA of non-dividing cells in the adult feline neuraxis. Journal of Cell Biology, 4(5):664– ??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/664.

Koenig:1955:SVS

[KP55] Virgil L. Koenig and J. D. Perrings. Sedimentation and viscosity studies on the capsular and somatic polysaccharides of Pneumococcus type III. Journal of Cell Biology, 1(2):93–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/2/93.

Kellenberger:1958:CWC

[KR58] Edouard Kellenberger and Antoinette Ryter. Cell Wall and Cytoplasmic Membrane of Escherichia coli. Journal of Cell Biology, 4(3):323–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/3/
323.

Kellenberger:1958:EMS

[KRS58] Edouard Kellenberger, Antoinette Ryter, and Janine Séchaud. Electron Microscope Study of DNA– Containing Plasms. Journal of Cell Biology, 4(6):671– ??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/671.

King:1959:RBP

[KSB59] Donald W. King, Edward L. Socolow, and Klaus G. Bensch. The Relation between Protein Synthesis and Lipide Accumulation in L Strain Cells and Ehrlich Ascites Cells. Journal of Cell Biology, 5(3):421–??, May 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/3/
421.

Kleinfeld:1959:ETR

[KvH59] R. G. Kleinfeld and E. von Haam. Effect of Thioacetamide on Rat Liver Regeneration. Journal of Cell Biology, 6(3):393–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/3/393.

Kallman:1958:FSC

[KWDV58] Frances Kallman, Robley C. Williams, Renato Dulbecco, and Marguerite Vogt. Fine Structure of Changes Produced in Cultured Cells Sampled at Specified Intervals During a Single Growth Cycle of Polio Virus. Journal of Cell Biology, 4(3):301–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/3/301.

Lacy:1957:GAN

[Lac57] Dennis Lacy. The Golgi apparatus in neurons and epithelial cells of the common limpet patella vulgata. Journal of Cell Biology, 3(5):779–??, September 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/5/
779.

Ladman:1958:FSR

[Lad58] Aaron J. Ladman. The Fine Structure of the Rod– Bipolar Cell Synapse in the Retina of the Albino Rat. Journal of Cell Biology, 4(4):459–??, July 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/4/
459.

Lafontaine:1958:PCF

[Laf58a] J. G. Lafontaine. A particulate component found in nucleoli of Allium cepa and Vicia faba. Journal of Cell Biology, 4(2):229–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/2/229.

Lafontaine:1958:SMF

[Laf58b] J. G. Lafontaine. Structure and Mode of Formation of the Nucleolus in Meristematic Cells of Vicia faba and Allium cepa. Journal of Cell Biology, 4(6):777– ??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/777.

Latta:1959:CRA

[Lat59] Harrison Latta. A Cellular Reaction to Antibody in Tissue Culture Studied with Electron Microscopy. Journal of Cell Biology, 5(3):405–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/3/
405.

Leuchtenberger:1957:OIC

[LB57] Cecile Leuchtenberger and Georgiana S. Boyer. The Occurrence of Intranuclear Crystals in Living HeLa Cells Infected with Adenoviruses. Journal of Cell Biology, 3 (2):323–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/2/323.

LaBella:1958:DHA

[LB58] Frank S. LaBella and J. H. U. Brown. Distribution of hydrolases among anterior pituitary cell fractions (anterior pituitary hydrolases). Journal of Cell Biology, 4 (6):833–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/6/833.

Labella:1959:CFA

[LB59] Frank S. Labella and J. H. U. Brown. Cell Fractionation of Anterior Pituitary Glands from Beef and Pig. Journal of Cell Biology, 5(1):17–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/
17.

Lowe:1959:CSD

[LBVS59] Charles U. Lowe, Harold Box, P. R. Venkataraman, and D. S. Sarkaria. Cytophotometric Study of Deoxyribonucleic Acid in Cortisone–Treated Rat Hepatocytes. Journal of Cell Biology, 5(2):251–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
251.

Lacy:1956:SGA

[LC56] Dennis Lacy and Cyril E. Challice. Studies on the Golgi apparatus by electron microscopy with particular reference to Aoyama’s technique. Journal of Cell Biology, 2 (4):395–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/395.

Lever:1958:MIR

[LC58] J. D. Lever and J. B. Chappell. Mitochondria Isolated from Rat Brown Adipose Tissue and Liver. Journal of Cell Biology, 4(3):287–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/3/287.

Lasansky:1959:SCV

[LD59] Arnaldo Lasansky and Eduardo De Robertis. Submicroscopic Changes in Visual Cells of the Rabbit Induced by Iodoacetate. Journal of Cell Biology, 5(2): 245–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/2/245.

Leuchtenberger:1955:BOT

[LDK55] Cecilie Leuchtenberger, Paul F. Doolin, and Anne H. Kutsakis. Buffered osmium tetroxide (OsO_{4}) fixation for cytological and Feulgen microspectrophotometric studies of human rectal polyps. Journal of Cell Biology, 1(5): 385–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/5/385.

Lever:1956:PIC

[Lev56a] J. D. Lever. Physiologically induced changes in adrenocortical mitochondria. Journal of Cell Biology, 2(4): 313–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/313.

Lever:1956:SSC

[Lev56b] J. D. Lever. The subendothelial space in certain endocrine tissues. Journal of Cell Biology, 2(4):293–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/293.

Low:1956:SEC

[LF56] Frank N. Low and James A. Freeman. Some Experiments with Chromium Compounds as Fixers for Electron Microscopy. Journal of Cell Biology, 2(5):629– ??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/5/629.

Luft:1957:EMC

[LH57] John Luft and Oscar Hechter. An electron microscopic correlation of structure with function in the isolated perfused cow adrenal, preliminary observations. Journal of Cell Biology, 3(4):615–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/4/615.

Littau:1959:CEB

[Lit59] Virginia C. Littau. Cytological Evidence that Both RNA and DNA May Form a Complex with the Same Protein. Journal of Cell Biology, 5(2):231–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
5/2/231.

Lowe:1955:OPL

[LL55] Charles Upton Lowe and Albert L. Lehninger. Oxidation and Phosphorylation in Liver Mitochondria Lacking “Polymerized” Ribonucleic Acid. Journal of Cell Biology, 1(1):89–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/1/89.

Lasfargues:1959:PMA

[LMM^{+}59] E. Y. Lasfargues, Dan H. Moore, Margaret R. Murray, Cushman D. Haagensen, and E. C. Pollard. Production of the Milk Agent in Cultures of Mouse Mammary Carcinoma. Journal of Cell Biology, 5(1):93–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
5/1/93.

Lowe:1955:ECR

[LMS55] Charles Upton Lowe, Deloris MacKinney, and Daljit Sarkaria. Effects of cortisone on rat liver mitochondria. Journal of Cell Biology, 1(3):237–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/3/237.

Lehrer:1959:DCM

[LO59] Gerard M. Lehrer and Leonard Ornstein. A Diazo Coupling Method for the Electron Microscopic Localization of Cholinesterase. Journal of Cell Biology, 6(3):399– ??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/399.

Locker:1959:SPO

[Loc59] Ronald H. Locker. Striation patterns of ox muscle in rigor mortis. Journal of Cell Biology, 6(3):419– ??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/419.

Love:1956:CCP

[LOK56] Robert Love, Ernest V. Orsi, and Hilary Koprowski. Cytological and cytochemical properties of ten ascites tumors. Journal of Cell Biology, 2(1):1–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/1/1.

Love:1957:NPA

[Lov57] Robert Love. The Nucleolus and Parachromatin of the Ascites Tumor Cell. Journal of Cell Biology, 3(1): 123–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/123.

Low:1956:MS

[Low56] Frank N. Low. Mitochondrial structure. Journal of Cell Biology, 2(4):337–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/337.

Low:1956:PSC

[LP56] Barbara W. Low and Peter R. Pinnock. The Packing Structure of Crystalline RI–APC Virus. Journal of Cell Biology, 2(4):483–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/483.

Lowe:1956:ECD

[LR56a] Charles U. Lowe and Royden N. Rand. The effect of cortisone on DNA content of rat hepatocytes. Journal of Cell Biology, 2(6):711–??, November 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/6/

Lowe:1956:ABC

[LR56b] Charles Upton Lowe and Royden N. Rand. Alterations in biochemical composition and ribonucleic acid metabolism induced in rat liver by cortisone. Journal of Cell Biology, 2(3):331–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/3/331.

Lafontaine:1958:EMS

[LR58] Jean G. Lafontaine and Hans Ris. An Electron Microscope Study of Lampbrush Chromosomes. Journal of Cell Biology, 4(1):99–??, January 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/1/
99.

Lehninger:1959:SRL

[LRS59] Albert L. Lehninger, Betty Lou Ray, and Marion Schneider. The Swelling of Rat Liver Mitochondria by Thyroxine and its Reversal. Journal of Cell Biology, 5(1): 97–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/97.

Leuchtenberger:1955:EDN

[LS55] Cecilie Leuchtenberger and Franz Schrader. Exceptional Desoxyribose Nucleic Acid (DNA) Findings in a Sterile Dwarf Bull. Journal of Cell Biology, 1(6):615– ??, November 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/6/615.

Locker:1957:SCS

[LS57] Ronald H. Locker and Francis O. Schmitt. Some chemical and structural properties of paramyosin. Journal of Cell Biology, 3(6):889–??, November 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/6/
889.

Lehninger:1959:MSI

[LS59] Albert L. Lehninger and Marion Schneider. Mitochondrial Swelling Induced by Glutathione. Journal of Cell Biology, 5(1):109–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/1/109.

Langridge:1957:MSD

[LSW^{+}57] R. Langridge, W. E. Seeds, H. R. Wilson, C. W. Hooper, M. H. F. Wilkins, and L. D. Hamilton. Molecular structure of deoxyribonucleic acid (DNA). Journal of Cell Biology, 3(5):767–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/5/767.

Luft:1956:FSE

[Luf56a] John H. Luft. The fine structure of the electric organ of the electric eel and torpedo ray. Journal of Cell Biology, 2(4):229–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/229.

Luft:1956:PNF

[Luf56b] John H. Luft. Permanganate — a new fixative for electron microscopy. Journal of Cell Biology, 2(6):799– ??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/6/799.

Luse:1956:EMO

[Lus56a] Sarah A. Luse. Electron microscopic observations of the central nervous system. Journal of Cell Biology, 2(5): 531–??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/5/531.

Luse:1956:FMC

[Lus56b] Sarah A. Luse. Formation of myelin in the central nervous system of mice and rats, as studied with the electron microscope. Journal of Cell Biology, 2(6):777– ??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/6/777.

Lund:1958:BCC

[LVH58] H. A. Lund, A. E. Vatter, and J. B. Hanson. Biochemical and Cytological Changes Accompanying Growth and Differentiation in the Roots of Zea mays. Journal of Cell Biology, 4(1):87–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/1/87.

Leduc:1959:EMS

[LW59] Elizabeth H. Leduc and J. Walter Wilson. An electron microscope study of intranuclear inclusions in mouse liver and hepatoma. Journal of Cell Biology, 6(3):427– ??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/427.

Ladman:1958:EMS

[LY58] Aaron J. Ladman and William C. Young. An Electron Microscopic Study of the Ductuli Efferentes and Rete Testis of the Guinea Pig. Journal of Cell Biology, 4(2): 219–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/2/219.

Manton:1959:OMS

[Man59] Irene Manton. Observations on the Microanatomy of the spermatozoid of the bracken fern (Pteridium aquilinum). Journal of Cell Biology, 6(3):413–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/3/413.

Matoltsy:1956:SSE

[Mat56] A. Gedeon Matoltsy. Sedimentation studies of epidermal keratins. Keratin A and keratin B. Journal of Cell Biology, 2(3):361–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/3/361.

Matsui:1958:PCT

[Mat58] Chiaki Matsui. Pathological cytology of tobacco leaf infected with tobacco mosaic virus III. Journal of Cell Biology, 4(6):831–??, November 1958. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/831.

Matoltsy:1955:SCC

[MB55] A. Gedeon Matoltsy and Constance A. Balsamo. A study of the components of the cornified epithelium of human skin. Journal of Cell Biology, 1(4):339–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/4/339.

Mercer:1959:EML

[MB59] E. H. Mercer and P. C. J. Brunet. The Electron Microscopy of the Left Colleterial Gland of the Cockroach. Journal of Cell Biology, 5(2):257–??, March 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/2/
257.

Morgan:1955:MPL

[MBMR55] Councilman Morgan, G. H. Bergold, Dan H. Moore, and Harry M. Rose. The macromolecular paracrystalline lattice of insect viral polyhedral bodies demonstrated in ultrathin sections examined in the electron microscope. Journal of Cell Biology, 1(3):187–??, May 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/3/
187.

Morgan:1956:USS

[MBR56] Councilman Morgan, G. H. Bergold, and Harry M. Rose. Use of serial sections to delineate the structure of porthetria dispar virus in the electron microscope. Journal of Cell Biology, 2(1):23–??, January 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/1/
23.

Mende:1958:SST

[MC58] Thomas J. Mende and Edward L. Chambers. Studies on Solute Transfer in the Vascular Endothelium. Journal of Cell Biology, 4(3):319–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/3/319.

Mendelsohn:1958:TWMa

[Men58a] Mortimer L. Mendelsohn. The Two–Wavelength Method of Microspectrophotometry. Journal of Cell Biology, 4 (4):407–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/4/407.

Mendelsohn:1958:TWMb

[Men58b] Mortimer L. Mendelsohn. The Two–Wavelength Method of Microspectrophotometry. Journal of Cell Biology, 4 (4):415–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/4/415.

Mendelsohn:1958:TWMc

[Men58c] Mortimer L. Mendelsohn. The Two–Wavelength Method of Microspectrophotometry. Journal of Cell Biology, 4 (4):425–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/4/425.

Merriam:1958:CLO

[Mer58] R. W. Merriam. The Contribution of Lower Oxides of Osmium to the Density of Biological Specimens in Electron Microscopy. Journal of Cell Biology, 4(5):579– ??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/579.

Merriam:1959:OFA

[Mer59a] R. W. Merriam. The Origin and Fate of Annulate Lamellae in Maturing Sand Dollar Eggs. Journal of Cell Biology, 5(1):117–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/1/117.

Merriam:1959:PSC

[Mer59b] R. W. Merriam. Permeability and Structural Characteristics of Isolated Nuclei from Chaetopterus Eggs. Journal of Cell Biology, 6(3):353–??, December 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/3/

Merz:1959:EEA

[Mer59c] Timothy Merz. The Effect of Extended Anaerobic Treatments on the Chromosomes of Vicia faba. Journal of Cell Biology, 5(1):135–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/1/135.

Muhlethaler:1959:ESP

[MFW59] Kurt Mühlethaler and Albert Frey-Wyssling. Entwicklung und Struktur der Proplastiden. (German) [Development and structure of proplastids]. Journal of Cell Biology, 6(3):507–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/3/507.

Moore:1957:PME

[MG57] Dan H. Moore and Philip M. Grimley. Problems in methacrylate embedding for electron microscopy. Journal of Cell Biology, 3(2):255–??, March 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/2/
255.

Morgan:1957:EMH

[MGR^{+}57] Councilman Morgan, Gabriel C. Godman, Harry M. Rose, Calderon Howe, and Joseph S. Huang. Electron Microscopic and Histochemical Studies of an Unusual Crystalline Protein Occurring in Cells Infected by Type 5 Adenovirus. Preliminary Observations. Journal of Cell Biology, 3(3):505–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/3/505.

Maxfield:1955:TPP

[MH55] Myles Maxfield and Robert W. Hartley. Two proteins purified from lobster nerve extract. Journal of Cell Biology, 1(4):279–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/4/279.

Morgan:1956:SDV

[MHRM56] Councilman Morgan, Calderon Howe, Harry M. Rose, and Dan H. Moore. Structure and development of viruses observed in the electron microscope. Journal of Cell Biology, 2(3):351–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/3/351.

Miller:1957:MOC

[Mil57] William H. Miller. Morphology of the ommatidia of the compound eye of limulus. Journal of Cell Biology, 3 (3):421–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/3/421.

Miller:1958:DCD

[Mil58] William H. Miller. Derivatives of Cilia in the Distal Sense Cells of the Retina of Pecten. Journal of Cell Biology, 4 (2):227–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/2/227.

McMaster-Kaye:1958:ETM

[MKT58] Rachel McMaster-Kaye and J. Herbert Taylor. Evidence for Two Metabolically Distinct Types of Ribonucleic Acid in Chromatin and Nucleoli. Journal of Cell Biology, 4(1):5–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/1/5.

McMaster-Kaye:1959:MCR

[MKT59] Rachel McMaster-Kaye and J. Herbert Taylor. The metabolism of chromosomal ribonucleic acid in Drosophila salivary glands and its relation to synthesis of desoxyribonucleic acid. Journal of Cell Biology, 5(3):461–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/3/461.

Monty:1956:IPL

[MLKD56] K. J. Monty, M. Litt, E. R. M. Kay, and A. L. Dounce. Isolation and properties of liver cell nucleoli. Journal of Cell Biology, 2(2):127–??, March 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/2/

Moore:1959:CPB

[MLM^{+}59] Dan H. Moore, E. Y. Lasfargues, Margaret R. Murray, Cushman D. Haagensen, and E. C. Pollard. Correlation of Physical and Biological Properties of Mouse Mammary Tumor Agent. Journal of Cell Biology, 5(1): 85–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/85.

Marshak:1955:TAS

[MM55] A. Marshak and C. Marshak. Thymine in the acidsoluble fraction of Arbacia eggs. Journal of Cell Biology, 1(2):167–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/2/167.

Morgan:1957:NAT

[MM57] Joseph F. Morgan and Helen J. Morton. The nutrition of animal tissues cultivated in vitro. Journal of Cell Biology, 3(2):141–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/2/141.

Morgan:1956:SEM

[MMR56] Councilman Morgan, Dan H. Moore, and Harry M. Rose. Some effects of the microtome knife and electron beam on methacrylate-embedded thin sections. Journal of Cell Biology, 2(4):21–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/21.

Matoltsy:1955:ISC

[MO55] A. Gedeon Matoltsy and George F. Odland. Investigation of the structure of the cornified epithelium of the human skin. Journal of Cell Biology, 1(3):191–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/1/3/191.

Mollenhauer:1959:PFP

[Mol59] Hilton H. Mollenhauer. Permanganate fixation of plant cells. Journal of Cell Biology, 6(3):431–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/3/431.

Montagna:1955:HCH

[Mon55] William Montagna. Histology and cytochemistry of human skin. Journal of Cell Biology, 1(1):13–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/1/1/13.

Montagna:1957:HCH

[Mon57] William Montagna. Histology and cytochemistry of human skin. Journal of Cell Biology, 3(3):343–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/3/343.

Moses:1956:CSC

[Mos56a] Montrose J. Moses. Chromosomal Structures in Crayfish Spermatocytes. Journal of Cell Biology, 2(2): 215–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/2/215.

Moses:1956:SNU

[Mos56b] Montrose J. Moses. Studies on nuclei using correlated cytochemical, light, and electron microscope techniques. Journal of Cell Biology, 2(4):397–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
397.

Moses:1958:RBA

[Mos58] Montrose J. Moses. The relation between the axial complex of meiotic prophase chromosomes and chromosome pairing in a salamander (Plethodon cinereus). Journal of Cell Biology, 4(5):633–??, September 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
633.

Maxwell:1956:EMC

[MP56] David S. Maxwell and Daniel C. Pease. The electron microscopy of the choroid Plexus. Journal of Cell Biology, 2(4):467–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/467.

Morton:1956:NAT

[MPM56] Helen J. Morton, Arthur E. Pasieka, and Joseph F. Morgan. The nutrition of animal tissues cultivated in vitro. Journal of Cell Biology, 2(5):589–??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/5/589.

Meyer:1959:EMS

[MQ59] H. Meyer and L. T. Queiroga. An Electron Microscope Study of Embryonic Heart Muscle Cells Grown in Tissue Cultures. Journal of Cell Biology, 5(1):169– ??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/169.

Moore:1957:EMS

[MR57a] Dan H. Moore and Helmut Ruska. Electron microscope study of mammalian cardiac muscle cells. Journal of Cell Biology, 3(2):261–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/2/261.

Moore:1957:FSC

[MR57b] Dan H. Moore and Helmut Ruska. The fine structure of capillaries and small arteries. Journal of Cell Biology, 3 (3):457–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/3/457.

Mendelsohn:1958:CST

[MR58] Mortimer L. Mendelsohn and Brian M. Richards. A Comparison of Scanning and Two–Wavelength Microspectrophotometry. Journal of Cell Biology, 4(6): 707–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/707.

Moore:1956:EMH

[MRC56] Dan H. Moore, Helmut Ruska, and Wilfred M. Copenhaver. Electron microscopic and histochemical observations of muscle degeneration after tourniquet. Journal of Cell Biology, 2(6):755–??, November 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/6/
755.

Muir:1957:EMS

[Mui57] Alan R. Muir. An electron microscope study of the embryology of the intercalated disc in the heart of the rabbit. Journal of Cell Biology, 3(2):193–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/2/193.

Munger:1958:POP

[Mun58] Bryce L. Munger. Polarization Optical Properties of the Pancreatic Acinar Cell of the Mouse. Journal of Cell Biology, 4(2):177–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/2/177.

Naora:1957:MCN

[Nao57] Hiroto Naora. Microspectrophotometry of cell nuclei stained with the Feulgen reaction. Journal of Cell Biology, 3(6):949–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/6/949.

Novikoff:1956:EML

[NBdD56] Alex B. Novikoff, H. Beaufay, and C. de Duve. Electron microscopy of lysosome-rich fractions from rat liver. Journal of Cell Biology, 2(4):179–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
179.

Napolitano:1958:FSB

[NF58] Leonard Napolitano and Don Fawcett. The Fine Structure of Brown Adipose Tissue in the Newborn Mouse and Rat. Journal of Cell Biology, 4(6):685–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/6/685.

Novikoff:1957:ERB

[NH57] Alex B. Novikoff and S. J. Holt. Esterase-rich Bodies in Osmium–Fixed Cells of Rat Kidney and Liver. Journal of Cell Biology, 3(1):127–??, January 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/1/
127.

Nebel:1956:EAC

[NM56] B. R. Nebel and O. T. Minick. Embedding aid for coverslip mounts in electron microscopy. Journal of Cell Biology, 2(4):61–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/61.

Novikoff:1956:PFS

[Nov56] Alex B. Novikoff. Preservation of the fine structure of isolated liver cell particulates with polyvinylpyrrollidonesucrose. Journal of Cell Biology, 2(4):65–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/65.

Novikoff:1959:PTC

[Nov59] Alex B. Novikoff. The Proximal Tubule Cell in Experimental Hydronephrosis. Journal of Cell Biology, 6(1): 136–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/136.

Nachmias:1958:HSN

[NP58] Vivianne T. Nachmias and Helen A. Padykula. A Histochemical Study of Normal and Denervated Red and White Muscles of the Rat. Journal of Cell Biology, 4 (1):47–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/1/47.

Nachlas:1956:QEL

[NPS56] Marvin M. Nachlas, William Prinn, and Arnold M. Seligman. Quantitative estimation of lyo- and desmoenzymes in tissue sections with and without fixation. Journal of Cell Biology, 2(5):487–??, September 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/5/
487.

Nasatir:1959:CAA

[NS59] Maimon Nasatir and Herbert Stern. Changes in the Activities of Aldolase and of D -Glyceraldehyde-3-Phosphate Dehydrogenase during the Mitotic Cycle in Microspores of Lilium longiflorum. Journal of Cell Biology, 6(2):189–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/189.

Nachlas:1958:HLT

[NWS58a] Marvin M. Nachlas, Donald G. Walker, and Arnold M. Seligman. The Histochemical Localization of Triphosphopyridine Nucleotide Diaphorase. Journal of Cell Biology, 4(4):467–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/4/467.

Nachlas:1958:HMD

[NWS58b] Marvin M. Nachlas, Donald G. Walker, and Arnold M. Seligman. A Histochemical Method for the Demonstration of Diphosphopyridine Nucleotide Diaphorase. Journal of Cell Biology, 4(1):29–??, January 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/1/
29.

Obland:1958:FSI

[Obl58] George F. Obland. The Fine Structure of the Interrelationship of Cells in the Human Epidermis. Journal of Cell Biology, 4(5):529–??, September 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/

Odor:1956:UTP

[Odo56] D. Louise Odor. Uptake and transfer of particulate matter from the peritoneal cavity of the rat. Journal of Cell Biology, 2(4):105–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/105.

Okada:1957:EGI

[OP57] Shigefumi Okada and Lee D. Peachey. Effect of gamma irradiation on the desoxyribonuclease ii activity of isolated mitochondria. Journal of Cell Biology, 3(2): 239–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/2/239.

Ornstein:1956:HTS

[Orn56a] Leonard Ornstein. How thin should a section be? Journal of Cell Biology, 2(4):53–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
53.

Ornstein:1956:MNI

[Orn56b] Leonard Ornstein. Mitochondrial and nuclear interaction. Journal of Cell Biology, 2(4):351–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/351.

Ornstein:1956:NCT

[Orn56c] Leonard Ornstein. A “new” connective tissue fibril. Journal of Cell Biology, 2(4):297–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
297.

Ornstein:1957:OFF

[Orn57] Leonard Ornstein. ”osmiophilia,” Fact or Fiction? Journal of Cell Biology, 3(5):809–??, September 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/5/

Palade:1955:SPC

[Pal55a] George E. Palade. A small particulate component of the cytoplasm. Journal of Cell Biology, 1(1):59–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/1/1/59.

Palade:1955:SER

[Pal55b] George E. Palade. Studies on the endoplasmic reticulum. Journal of Cell Biology, 1(6):567–??, November 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/6/567.

Palade:1956:ER

[Pal56a] George E. Palade. The endoplasmic reticulum. Journal of Cell Biology, 2(4):85–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
85.

Palade:1956:IGE

[Pal56b] George E. Palade. Intracisternal granules in the exocrine cells of the pancreas. Journal of Cell Biology, 2(4):417–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/417.

Palay:1956:SCN

[Pal56c] Sanford L. Palay. Synapses in the central nervous system. Journal of Cell Biology, 2(4):193–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/193.

Pappas:1956:FSN

[Pap56a] George D. Pappas. The fine structure of the nuclear envelope of Amoeba Proteus. Journal of Cell Biology, 2 (4):431–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/431.

Pappas:1956:HSN

[Pap56b] George D. Pappas. Helical Structures in the Nucleus of Amoeba proteus. Journal of Cell Biology, 2(2): 221–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/2/221.

Paul:1956:CDD

[Pau56] J. Paul. The Chemical Determination of Deoxyribonucleic Acid in Tissue Cultures. Journal of Cell Biology, 2(6):797–??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/6/797.

Pappas:1958:FSC

[PB58] George D. Pappas and Philip W. Brandt. The Fine Structure of the Contractile Vacuole in Ameba. Journal of Cell Biology, 4(4):485–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/4/
485.

Pappas:1959:M

[PB59] George D. Pappas and Philip W. Brandt. Mitochondria. Journal of Cell Biology, 6(1):85–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/1/85.

Pasteels:1958:CLU

[PCV58] J. J. Pasteels, P. Castiaux, and G. Vandermeerssche. Cytochemical Localisations and Ultrastructure in the Fertilized Unsegmented Egg of Paracentrotus lividus. Journal of Cell Biology, 4(5):575–??, September 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
575.

Pease:1956:IBP

[Pea56] Daniel C. Pease. Infolded basal plasma membranes found in epithelia noted for their water transport. Journal of Cell Biology, 2(4):203–??, July 1956. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/203.

Peachey:1958:TS

[Pea58] Lee D. Peachey. Thin Sections. Journal of Cell Biology, 4(3):233–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/233.

Peachey:1959:DST

[Pea59] Lee D. Peachey. A Device for Staining Tissue Sections for Electron Microscopy. Journal of Cell Biology, 5 (3):511–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/3/511.

Powers:1956:IOM

[PERM56] E. L. Powers, C. F. Ehret, L. E. Roth, and O. T. Minick. The internal organization of mitochondria. Journal of Cell Biology, 2(4):341–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/4/341.

Petrushka:1959:EMT

[PG59] Evelyn Petrushka and Antonio Giuditta. Electron Microscopy of Two Subcellular Fractions Isolated from Cerebral Cortex Homogenate. Journal of Cell Biology, 6 (1):129–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/129.

Petermann:1955:SFC

[PH55] Mary L. Petermann and Mary G. Hamilton. A Stabilizing Factor for Cytoplasmic Nucleoproteins. Journal of Cell Biology, 1(5):469–??, September 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/5/
469.

Petermann:1958:IAS

[PH58] Mary L. Petermann and Mary G. Hamilton. The Influence of Age, Sex, Pregnancy, Starvation, and Other Factors on the Cytoplasmic Ribonucleoproteins of Rat Liver.

Journal of Cell Biology, 4(6):771–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/6/771.

Pitelka:1956:EMS

[Pit56] Dorothy R. Pitelka. An electron microscope study of cortical structures of opalina obtrigonoidea. Journal of Cell Biology, 2(4):423–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/423.

Palay:1959:EMSa

[PK59a] Sanford L. Palay and Leonard J. Karlin. An Electron Microscopic Study of the Intestinal Villus. Journal of Cell Biology, 5(3):363–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/3/363.

Palay:1959:EMSb

[PK59b] Sanford L. Palay and Leonard J. Karlin. An Electron Microscopic Study of the Intestinal Villus. Journal of Cell Biology, 5(3):373–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/3/373.

Plaut:1956:DNS

[PM56] Walter Plaut and Daniel Mazia. The distribution of newly synthesized DNA in mitotic division. Journal of Cell Biology, 2(5):573–??, September 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/5/
573.

Porter:1956:SRM

[Por56] Keith R. Porter. The sarcoplasmic reticulum in muscle cells of amblystoma larvae. Journal of Cell Biology, 2 (4):163–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/163.

Palay:1955:FSN

[PP55] Sanford L. Palay and George E. Palade. The fine structure of neurons. Journal of Cell Biology, 1(1):

69–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/1/69.

Porter:1957:SER

[PP57] Keith R. Porter and George E. Palade. Studies on the endoplasmic reticulum. Journal of Cell Biology, 3(2): 269–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/2/269.

Porter:1959:CFF

[PP59] Keith R. Porter and George D. Pappas. Collagen Formation by Fibroblasts of the Chick Embryo Dermis. Journal of Cell Biology, 5(1):153–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/
153.

Pease:1957:EMP

[PQ57] Daniel C. Pease and T. Andrew Quilliam. Electron microscopy of the pacinian corpuscle. Journal of Cell Biology, 3(3):331–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/3/331.

Plaut:1957:CIR

[PR57a] Walter Plaut and Ronald C. Rustad. Cytoplasmic Incorporation of a Ribonucleic Acid Precursor in Amoeba proteus. Journal of Cell Biology, 3(4):625–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/4/625.

Poy:1957:EMA

[PR57b] R. K. F. Pak Poy and J. S. Robertson. Electron microscopy of the avian renal glomerulus. Journal of Cell Biology, 3(2):183–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/2/183.

Prescott:1959:NSC

[Pre59] David M. Prescott. Nuclear Synthesis of Cytoplasmic Ribonucleic Acid in Amoeba proteus. Journal of Cell Biology, 6(2):203–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/203.

Palade:1956:LM

[PS56a] G. E. Palade and P. Siekevitz. Liver microsomes. Journal of Cell Biology, 2(2):171–??, March 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/2/
171.

Palade:1956:PM

[PS56b] G. E. Palade and P. Siekevitz. Pancreatic microsomes. Journal of Cell Biology, 2(6):671–??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/6/671.

Plaut:1958:ITC

[PS58] W. Plaut and L. Alexander Sagan. Incorporation of Thymidine in the Cytoplasm of Amoeba proteus. Journal of Cell Biology, 4(6):843–??, November 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/6/
843.

Quilliam:1958:NPF

[Qui58] T. Andrew Quilliam. New Problems in the Functional Activity of the Pacinian Corpuscle. Journal of Cell Biology, 4(3):341–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/3/341.

Rouiller:1956:MPM

[RB56] C. Rouiller and W. Bernhard. ”microbodies” and the problem of mitochondrial regeneration in liver cells. Journal of Cell Biology, 2(4):355–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
355.

Rosenkranz:1958:NDA

[RB58] Herbert S. Rosenkranz and Aaron Bendich. On the Nature of the Deoxyribonucleic Acid–Methyl Green Reaction. Journal of Cell Biology, 4(5):663–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/5/663.

Robinson:1956:EMC

[RC56] Robert A. Robinson and D. A. Cameron. Electron microscopy of cartilage and bone matrix at the distal epiphyseal line of the femur in the newborn infant. Journal of Cell Biology, 2(4):253–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
253.

Rudkin:1957:PMM

[RC57] George T. Rudkin and Sally L. Corlette. A Photographic Method for Measuring Ultraviolet Radiation Scattered from Microscopic Objects. Journal of Cell Biology, 3 (5):821–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/5/821.

Rebhun:1956:EMBa

[Reb56a] Lionel I. Rebhun. Electron microscopy of basophilic structures of some invertebrate oocytes. Journal of Cell Biology, 2(1):93–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/1/93.

Rebhun:1956:EMBb

[Reb56b] Lionel I. Rebhun. Electron microscopy of basophilic structures of some invertebrate oocytes. Journal of Cell Biology, 2(2):159–??, March 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/2/2/159.

Rebhun:1957:NCD

[Reb57] Lionel I. Rebhun. Nuclear changes during spermiogenesis in a pulmonate snail. Journal of Cell Biology, 3 (4):509–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/4/509.

Redman:1959:PTM

[RH59] Colvin M. Redman and Lowell E. Hokin. Phospholipide Turnover in Microsomal Membranes of the Pancreas during Enzyme Secretion. Journal of Cell Biology, 6(2): 207–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/2/207.

Richter:1958:EMH

[Ric58] Goetz W. Richter. Electron Microscopy of Hemosiderin: Presence of Ferritin and Occurrence of Crystalline Lattices in Hemosiderin Deposits. Journal of Cell Biology, 4 (1):55–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/1/55.

Richter:1959:ISA

[Ric59] Goetz W. Richter. Internal structure of apoferritin as revealed by the “negative staining technique.”. Journal of Cell Biology, 6(3):531–??, December 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/3/
531.

Revel:1958:EMM

[RIF58] Jean Paul Revel, Susumu Ito, and Don W. Fawcett. Electron Micrographs of Myelin Figures of Phospholipide Simulating Intracellular Membranes. Journal of Cell Biology, 4(4):495–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/4/495.

Ris:1956:SCE

[Ris56] Hans Ris. A study of chromosomes with the electron microscope. Journal of Cell Biology, 2(4):385–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/385.

Randall:1958:FSF

[RJ58] J. T. Randall and Sylvia Fitton Jackson. Fine Structure and Function in Stentor polymorphus. Journal of Cell Biology, 4(6):807–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/807.

Rorem:1957:RNL

[RM57] Edward S. Rorem and Leonard Machlis. The ribonucleoprotein nature of large particles in the meiosporangia of allomyces. Journal of Cell Biology, 3(6):879– ??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/879.

Rowley:1959:RTM

[RMFW59] J. R. Rowley, K. Mühlethaler, and A. Frey-Wyssling. A route for the transfer of materials through the pollen grain wall. Journal of Cell Biology, 6(3):537–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/3/537.

Ruska:1957:BPC

[RMW57] Helmut Ruska, Dan H. Moore, and John Weinstock. The base of the proximal convoluted tubule cells of rat kidney. Journal of Cell Biology, 3(2):249–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/2/249.

Robertson:1955:UAV

[Rob55] J. David Robertson. The ultrastructure of adult vertebrate peripheral myelinated nerve fibers in relation to myelinogenesis. Journal of Cell Biology, 1(4):271–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/1/4/271.

Robertson:1956:URM

[Rob56a] J. David Robertson. The ultrastructure of a reptilian myoneural junction. Journal of Cell Biology, 2(4): 381–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/381.

Robinow:1956:OVS

[Rob56b] C. F. Robinow. Observations on vase-shaped, ironcontaining houses of two colorless flagellates of the family bicoecidae. Journal of Cell Biology, 2(4):233–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/233.

Robertson:1957:NOU

[Rob57] J. David Robertson. New Observations on the Ultrastructure of the Membranes of Frog Peripheral Nerve Fibers. Journal of Cell Biology, 3(6):1043–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/6/1043.

Robertson:1958:SAN

[Rob58a] J. David Robertson. Structural Alterations in Nerve Fibers Produced by Hypotonic and Hypertonic Solutions. Journal of Cell Biology, 4(4):349–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/4/349.

Robertson:1958:USL

[Rob58b] J. David Robertson. The Ultrastructure of Schmidt– Lanterman Clefts and Related Shearing Defects of the Myelin Sheath. Journal of Cell Biology, 4(1):39–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/1/39.

Rose:1957:MVS

[Ros57] George G. Rose. Microkinetospheres and vp satellites of pinocytic cells observed in tissue cultures of Gey’s strain HeLa with phase contrast cinematographic techniques. Journal of Cell Biology, 3(5):697–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/5/697.

Ross:1959:EMO

[Ros59] Leonard L. Ross. Electron Microscopic Observations of the Carotid Body of the Cat. Journal of Cell Biology, 6(2):253–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/253.

Roth:1956:ACF

[Rot56] L. E. Roth. Aspects of ciliary fine structure in euplotes patella. Journal of Cell Biology, 2(4):235–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/235.

Roth:1957:EMS

[Rot57a] L. E. Roth. An electron microscope study of the cytology of the protozoan euplotes patella. Journal of Cell Biology, 3(6):985–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/6/985.

Roth:1957:SCM

[Rot57b] L. E. Roth. The Spermatid Cell Membrane in Melanoplus differentialis. Journal of Cell Biology, 3(5): 816–??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/816.

Rothschild:1957:MCS

[Rot57c] Lord Rothschild. The membrane capacitance of the sea urchin egg. Journal of Cell Biology, 3(1):103– ??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/103.

Rudzinska:1955:OFS

[RP55] Maria A. Rudzinska and Keith R. Porter. Observations on the fine structure of the macronucleus of tokophrya infusionum. Journal of Cell Biology, 1(5): 421–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/5/421.

Ross:1958:EMO

[RPLL58] Michael H. Ross, George D. Pappas, Jonathan T. Lanman, and John Lind. Electron Microscope Observations on the Endoplasmic Reticulum in the Human Fetal Adrenal. Journal of Cell Biology, 4(5):659–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/5/659.

Rose:1958:CST

[RPST58] George G. Rose, C. M. Pomerat, T. O. Shindler, and J. B. Trunnell. A Cellophane–Strip Technique for Culturing Tissue in Multipurpose Culture Chambers. Journal of Cell Biology, 4(6):761–??, November 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/6/
761.

Rizki:1959:FSC

[RR59] M. T. M. Rizki and Rose M. Rizki. Functional significance of the crystal cells in the larva of Drosophila melanogaster. Journal of Cell Biology, 5(2):235–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/2/235.

Rasch:1959:NCP

[RSK59] Ellen Rasch, Hewson Swift, and Richard M. Klein. Nucleoprotein Changes in Plant Tumor Growth. Journal of Cell Biology, 6(1):11–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
11.

Rudzinska:1959:PTN

[RT59] Maria A. Rudzinska and William Trager. Phagotrophy and Two New Structures in the Malaria Parasite Plasmodium berghei. Journal of Cell Biology, 6(1): 103–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/103.

Rudzinska:1955:UOF

[Rud55] Maria A. Rudzinska. The use of OsO_{4} fixative for Feulgen-stained preparations. Journal of Cell Biology, 1(5):472–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/5/472.

Rudzinska:1956:FOF

[Rud56] Maria A. Rudzinska. Further observations on the fine structure of the macronucleus in tokophrya infusionum. Journal of Cell Biology, 2(4):425–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
425.

Rudzinska:1958:EMS

[Rud58] Maria A. Rudzinska. An Electron Microscope Study of the Contractile Vacuole in Tokophrya infusionum. Journal of Cell Biology, 4(2):195–??, March 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/2/
195.

Ruthmann:1958:BLS

[Rut58] August Ruthmann. Basophilic Lamellar Systems in the Crayfish Spermatocyte. Journal of Cell Biology, 4(3): 267–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/267.

Ruthmann:1959:FSM

[Rut59] August Ruthmann. The Fine Structure of the Meiotic Spindle of the Crayfish. Journal of Cell Biology, 5(1): 177–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/177.

Rabinovitch:1959:AOS

[RV59] M. Rabinovitch and I. Vugman. Autoradiographic Observations on the Silk Glands of Bombix mori. Journal of Cell Biology, 6(2):293–??, October 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/2/
293.

Runge:1958:SMS

[RVH58] Walter J. Runge, Robert L. Vernier, and J. Francis Hartmann. A Staining Method for Sections of Osmium– Fixed, Methacrylate–Embedded Tissue. Journal of Cell Biology, 4(3):327–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/3/327.

Rasch:1959:BPP

[RW59] Ellen Rasch and John W. Woodard. Basic Proteins of Plant Nuclei during Normal and Pathological Cell Growth. Journal of Cell Biology, 6(2):263–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/2/263.

Robbins:1955:SEA

[RWPP55] William C. Robbins, Robert F. Watson, George D. Pappas, and Keith R. Porter. Some effects of anticollagen serum on collagen formation in tissue culture: A preliminary report. Journal of Cell Biology, 1(4): 381–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/4/381.

Sacktor:1955:CSM

[Sac55] Bertram Sacktor. Cell structure and the metabolism of insect flight muscle. Journal of Cell Biology, 1(1): 29–??, January 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/1/29.

Selby:1956:SAR

[SB56] Cecily Cannan Selby and Richard S. Bear. The structure of actin-rich filaments of muscles according to xray diffraction. Journal of Cell Biology, 2(1):71–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/1/71.

Setterfield:1958:ACM

[SB58] G. Setterfield and S. T. Bayley. Arrangement of Cellulose Microfibrils in Walls of Elongating Parenchyma Cells. Journal of Cell Biology, 4(4):377–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
4/4/377.

Silvester:1959:QEU

[SB59a] N. R. Silvester and R. E. Burge. A Quantitative Estimation of the Uptake of Two New Electron Stains by the Cytoplasmic Membrane of Ram Sperm. Journal of Cell Biology, 6(2):179–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/2/179.

Steffensen:1959:APT

[SB59b] Dale Steffensen and John A. Bergeron. Autoradiographs of Pollen Tube Nuclei with Calcium-45. Journal of Cell Biology, 6(3):339–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/3/339.

Sampaio:1958:AUB

[SBF58] Mario M. Sampaio, Adolfo Brunner, and Benedito Oliveira Filho. Aspects of the Ultrastructure of the Brush Border of the Kidney of Normal Mouse. Journal of Cell Biology, 4(3):335–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/3/335.

Setterfield:1955:CSA

[SD55] George Setterfield and Robert E. Duncan. Cytological studies on the antimetabolite action of 2,6diaminopurine in Vicia Faba roots. Journal of Cell Biology, 1(5):399–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/1/5/399.

Schreiber:1958:CAA

[SD58] Richard W. Schreiber and Robert E. Duncan. A Cytological Analysis of the Antimetabolite Activity of 5-Hydroxyuracil in Vicia faba Roots. Journal of Cell Biology, 4(6):793–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/793.

Shelton:1959:EME

[SD59] Emma Shelton and A. J. Dalton. Electron microscopy of emperipolesis. Journal of Cell Biology, 6(3):513– ??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/513.

Selby:1955:EMS

[Sel55] Cecily Cannan Selby. An electron microscope study of the epidermis of mammalian skin in thin sections. Journal of Cell Biology, 1(5):429–??, September 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/5/
429.

Slautterback:1959:DCH

[SF59] David B. Slautterback and Don W. Fawcett. The Development of the Cnidoblasts of Hydra. Journal of Cell Biology, 5(3):441–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/3/441.

Sheldon:1956:EMS

[She56] Huntington Sheldon. An electron microscope study of the epithelium in the normal mature and immature mouse cornea. Journal of Cell Biology, 2(3):253–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/3/253.

Sheldon:1957:MEG

[She57] Huntington Sheldon. A method for evaluating glass knives. Journal of Cell Biology, 3(4):621–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/4/621.

Shoenberg:1958:EMS

[Sho58] C. F. Shoenberg. An Electron Microscope Study of Smooth Muscle in Pregnant Uterus of the Rabbit. Journal of Cell Biology, 4(5):609–??, September 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
609.

Scarpelli:1958:CLO

[SHP58] D. G. Scarpelli, R. Hess, and A. G. E. Pearse. The Cytochemical Localization of Oxidative Enzymes. Journal of Cell Biology, 4(6):747–??, November 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/6/
747.

Sidman:1957:SCS

[Sid57] Richard L. Sidman. The structure and concentration of solids in photoreceptor cells studied by refractometry and interference microscopy. Journal of Cell Biology, 3 (1):15–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/15.

Siekevitz:1955:DOA

[Sie55] Philip Siekevitz. The Distinctive Occurrence of Acid– Soluble Inosine in Rat Liver Microsomal Preparations. Journal of Cell Biology, 1(5):477–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/5/477.

Smith:1957:SBC

[SJ57] R. H. Smith and Sylvia Fitton Jackson. Studies on the biosynthesis of collagen. Journal of Cell Biology, 3(6): 913–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/913.

Stern:1959:SCP

[SJS59] Herbert Stern, F. B. Johnston, and G. Setterfield. Some Chemical Properties of Isolated Pea Nucleoli. Journal of Cell Biology, 6(1):57–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
57.

Smith:1957:EMT

[SL57] Douglas E. Smith and Yevette S. Lewis. Electron microscopy of the tissue mast cell. Journal of Cell Biology, 3(1):9–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/1/9.

Schultz:1958:BAF

[SM58] Richard L. Schultz and Roland K. Meyer. Biological Activity of Fractions Obtained by Differential Centrifugation of the Bovine Adrenal Cortex. Journal of Cell Biology, 4(1):23–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/1/23.

Smith:1956:SIV

[Smi56] Kenneth M. Smith. The structure of insect virus particles. Journal of Cell Biology, 2(3):301–??, May 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/3/301.

Smith:1959:RAN

[Smi59] Stuart W. Smith. ”reticular” and “Areticular” Nissl Bodies in Sympathetic Neurons of a Lizard. Journal of Cell Biology, 6(1):77–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
77.

Spater:1958:AAC

[SNM58] Herman W. Spater, Alex B. Novikoff, and Bertha Masek. Adenosinetriphosphatase Activity in the Cell Membranes of Kidney Tubule Cells. Journal of Cell Biology, 4(6):765–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/765.

Sedar:1955:FSC

[SP55] Albert W. Sedar and Keith R. Porter. The fine structure of cortical components of paramecium multimicronucleatum. Journal of Cell Biology, 1(6):583–??, November 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/1/6/583.

Sager:1957:SDC

[SP57] Ruth Sager and George E. Palade. Structure and development of the chloroplast in chlamydomonas. Journal of Cell Biology, 3(3):463–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/3/463.

Satir:1958:TS

[SP58a] Peter G. Satir and Lee D. Peachey. Thin Sections. Journal of Cell Biology, 4(3):345–??, May 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/3/
345.

Savage:1958:EHH

[SP58b] R. E. Savage and Walter Plaut. The Effect of HCl Hydrolysis on the Retention of Thymidine in DNA. Journal of Cell Biology, 4(6):701–??, November 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/6/
701.

Siekevitz:1958:CSPa

[SP58c] Philip Siekevitz and George E. Palade. A Cytochemical Study on the Pancreas of the Guinea Pig. Journal of Cell Biology, 4(2):203–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/2/203.

Siekevitz:1958:CSPb

[SP58d] Philip Siekevitz and George E. Palade. A Cytochemical Study on the Pancreas of the Guinea Pig. Journal of Cell Biology, 4(3):309–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/3/309.

Siekevitz:1958:CSPc

[SP58e] Philip Siekevitz and George E. Palade. A Cytochemical Study on the Pancreas of the Guinea Pig. Journal of Cell Biology, 4(5):557–??, September 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
557.

Stich:1958:ERP

[SP58f] H. Stich and W. Plaut. The Effect of Ribonuclease on Protein Synthesis in Nucleated and Enucleated Fragments of Acetabularia. Journal of Cell Biology, 4(1): 119–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/1/119.

Swim:1958:RCD

[SP58g] H. E. Swim and R. F. Parker. The Role of Carbon Dioxide as an Essential Nutrient for Six Permanent Strains of Fibroblasts. Journal of Cell Biology, 4(5): 525–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/525.

Siekevitz:1959:CSP

[SP59a] Philip Siekevitz and George E. Palade. A Cytochemical Study on the Pancreas of the Guinea Pig. Journal of Cell Biology, 5(1):1–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/1/1.

Sotelo:1959:EMS

[SP59b] J. Roberto Sotelo and Keith R. Porter. An Electron Microscope Study of the Rat Ovum. Journal of Cell Biology, 5(2):327–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/2/327.

Spencer:1959:PCF

[Spe59] M. Spencer. The Preparation of Carbon Films for Electron Microscopy. Journal of Cell Biology, 6(1):

125–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/125.

Spiro:1956:USM

[Spi56] David Spiro. The ultrastructure of striated muscle at various sarcomere lengths. Journal of Cell Biology, 2 (4):157–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/157.

Sedar:1956:MP

[SR56] Albert W. Sedar and Maria A. Rudzinska. Mitochondria of protozoa. Journal of Cell Biology, 2(4):331–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/331.

Sheldon:1957:EMS

[SR57] Huntington Sheldon and Robert A. Robinson. Electron microscope studies of crystal-collagen relationships in bone. Journal of Cell Biology, 3(6):1011–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/6/1011.

Sheldon:1958:SCE

[SR58] Huntington Sheldon and Robert A. Robinson. Studies on Cartilage: Electron Microscope Observations on Normal Rabbit Ear Cartilage. Journal of Cell Biology, 4 (4):401–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/4/401.

Sechaud:1959:CQC

[SRK59] Janine Séchaud, Antoinette Ryter, and Edouard Kellenberger. Considérations quantitatives sur des coupes ultraminces de bactéries infectées par un bactériophage. Journal of Cell Biology, 5(3):469–??, May 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/3/

Sacktor:1956:ETO

[SS56] Bertram Sacktor and Richard Sanborn. The Effect of Temperature on Oxidative Phosphorylation with Insect Flight Muscle Mitochondria. Journal of Cell Biology, 2(1):105–??, January 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/1/105.

Salpeter:1959:FSA

[SS59] Miriam M. Salpeter and Marcus Singer. The Fine Structure of the Adepidermal Reticulum in the Basal Membrane of the Skin of the Newt, Triturus. Journal of Cell Biology, 6(1):35–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/1/35.

Shatkin:1959:EMN

[ST59] A. J. Shatkin and E. L. Tatum. Electron microscopy of Neurospora crassa Mycelia. Journal of Cell Biology, 6 (3):423–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/3/423.

Sotelo:1957:EMS

[STC57] J. Roberto Sotelo and Omar Trujillo-Cenóz. Electron microscope study of the vitelline body of some spider oocytes. Journal of Cell Biology, 3(2):301–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/2/301.

Sotelo:1959:FSE

[STC59] J. Roberto Sotelo and O. Trujillo-Cenóz. The Fine Structure of an Elementary Contractile System. Journal of Cell Biology, 6(1):126–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
126.

Stephenson:1956:ICG

[Ste56] John L. Stephenson. Ice crystal growth during the rapid freezing of tissues. Journal of Cell Biology, 2 (4):45–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/45.

Steere:1957:EMS

[Ste57] Russell L. Steere. Electron microscopy of structural detail in frozen biological specimens. Journal of Cell Biology, 3(1):45–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/3/1/45.

Stern:1958:VSC

[Ste58] Herbert Stern. Variations in Sulfhydryl Concentration During Microsporocyte Meiosis in the Anthers of Lilium and Trillium. Journal of Cell Biology, 4(2): 157–??, March 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/2/157.

Stoeckenius:1959:EMS

[Sto59] Walther Stoeckenius. An Electron Microscope Study of Myelin Figures. Journal of Cell Biology, 5(3):491–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/3/491.

Straus:1956:CAP

[Str56] Werner Straus. Concentration of acid phosphatase, ribonuclease, desoxyribonuclease, β-glucuronidase, and cathepsin in “droplets” isolated from the kidney cells of normal rats. Journal of Cell Biology, 2(5):513– ??, September 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/5/513.

Straus:1957:CDF

[Str57a] Werner Straus. Changes in “droplet” fractions from rat kidney cells after intraperitoneal injection of egg white. Journal of Cell Biology, 3(6):933–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
3/6/933.

Straus:1957:SII

[Str57b] Werner Straus. Segregation of an Intravenously Injected Protein by “Droplets” of the Cells of Rat Kidneys. Journal of Cell Biology, 3(6):1037–??, November 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/6/
1037.

Straus:1958:CAD

[Str58] Werner Straus. Colorimetric analysis with N,Ndimethyl-p-phenylenediamine of the uptake of intravenously injected horseradish peroxidase by various tissues of the rat. Journal of Cell Biology, 4(5):541– ??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/541.

Straus:1959:RCI

[Str59a] Werner Straus. Rapid Cytochemical Identification of Phagosomes in Various Tissues of the Rat and their Differentiation from Mitochondria by the Peroxidase Method. Journal of Cell Biology, 5(2):193–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/5/2/193.

Strunk:1959:FIC

[Str59b] Stanley W. Strunk. The Formation of Intracellular Crystals in Midgut Glands of Limnoria lignorum. Journal of Cell Biology, 5(3):385–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/3/385.

Sulkin:1955:HSM

[Sul55] Norman M. Sulkin. Histochemical studies on mucoproteins in nerve cells of the dog. Journal of Cell Biology, 1(5):459–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/1/5/459.

Siekevitz:1956:CSM

[SW56] Philip Siekevitz and Michael L. Watson. Cytochemical studies of mitochondria. Journal of Cell Biology, 2(6):

653–??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/6/653.

Swift:1956:FSA

[Swi56] Hewson Swift. The fine structure of annulate lamellae. Journal of Cell Biology, 2(4):415–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
415.

Taylor:1959:DSF

[Tay59] Edwin W. Taylor. Dynamics of Spindle Formation and Its Inhibition by Chemicals. Journal of Cell Biology, 6(2):193–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/193.

Tobioka:1956:MLC

[TB56] Motohiko Tobioka and John J. Biesele. Mitochondria in living cells: an analysis of movements. Journal of Cell Biology, 2(4):319–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/319.

Tso:1956:MNP

[TBV56] Paul O. P. Ts’o, James Bonner, and Jerome Vinograd. Microsomal nucleoprotein particles from pea seedlings. Journal of Cell Biology, 2(4):451–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
451.

Tonna:1959:HAS

[TC59] Edgar A. Tonna and Eugene P. Cronkite. Histochemical and Autoradiographic Studies on the Effects of Aging on the Mucopolysaccharides of the Periosteum. Journal of Cell Biology, 6(2):171–??, October 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/2/

Trujillo-Cenoz:1959:ROS

[TCS59] O. Trujillo-Cenóz and J. Roberto Sotelo. Relationships of the Ovular Surface with Follicle Cells and Origin of the Zona Pellucida in Rabbit Oocytes. Journal of Cell Biology, 5(2):347–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/2/347.

Thornburg:1956:PEM

[TD56] Wayne Thornburg and Eduardo De Robertis. Polarization and electron microscope study of frog nerve axoplasm. Journal of Cell Biology, 2(4):475–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/475.

Tedeschi:1959:SMM

[Ted59] Henry Tedeschi. The Structure of the Mitochondrial Membrane: Inferences from Permeability Properties. Journal of Cell Biology, 6(2):241–??, October 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/2/
241.

Threadgold:1958:EMS

[TG58] L. T. Threadgold and R. A. R. Gresson. An Electron Microscope Study of the Cells Lining the Small Blood Vessels of Helix pomatia Linn. Journal of Cell Biology, 4(3):342–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/342.

Thaemert:1959:IBP

[Tha59] J. C. Thaemert. Intercellular Bridges as Protoplasmic Anastomoses between Smooth Muscle Cells. Journal of Cell Biology, 6(1):67–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
67.

Thornburg:1957:FBL

[Tho57] Wayne Thornburg. The form birefringence of lamellar systems containing three or more components. Journal of Cell Biology, 3(3):413–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/3/3/413.

Takahashi:1955:AAH

[TI55] Kohnosuke Takahashi and Shoji Iwase. On artifacts appearing in the histochemical fixation of glycogen. Journal of Cell Biology, 1(5):391–??, September 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/5/
391.

Tokuyasu:1959:NMM

[TO59] K. Tokuyasu and S. Okamura. A New Method for Making Glass Knives for Thin Sectioning. Journal of Cell Biology, 6(2):305–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/6/2/305.

Tahmisian:1956:LEM

[TPD56] T. N. Tahmisian, E. L. Powers, and R. L. Devine. Light and electron microscope studies of morphological changes of mitochondria during spermatogenesis in the grasshopper. Journal of Cell Biology, 2(4):325–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/325.

Trier:1958:FSP

[Tri58] Jerry Steven Trier. The Fine Structure of the Parathyroid Gland. Journal of Cell Biology, 4(1):13–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/4/1/13.

Trowell:1957:RUL

[Tro57] O. A. Trowell. Re-utilization of lymphocytes in lymphopoiesis. Journal of Cell Biology, 3(2):317–??, March 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/2/317.

Tso:1959:ILC

[TS59] Paul O. P. Ts’o and Clifford S. Sato. The incorporation of leucine-C^{14} into microsomal particles and other subcellular components of the pea epicotyl. Journal of Cell Biology, 5(1):59–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/5/1/59.

Tsujii:1958:SSF

[TSW58] Tadashi Tsujii, D. Gordon Sharp, and Karl M. Wilbur. Studies on Shell Formation. Journal of Cell Biology, 4 (3):275–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/275.

Tokuyasu:1959:FSBa

[TY59a] Kiyoteru Tokuyasu and Eichi Yamada. Fine Structure of Bacillus subtilis. Journal of Cell Biology, 5(1): 123–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/123.

Tokuyasu:1959:FSBb

[TY59b] Kiyoteru Tokuyasu and Eichi Yamada. Fine Structure of Bacillus subtilis. Journal of Cell Biology, 5(1): 129–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/129.

Tokuyasu:1959:FSR

[TY59c] Kiyoteru Tokuyasu and Eichi Yamada. The Fine Structure of the Retina Studied with the Electron Microscope. Journal of Cell Biology, 6(2):225–??, October 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/2/
225.

Underhay:1956:ILE

[UHBdD56] Eluned Underhay, S. J. Holt, H. Beaufay, and C. de Duve. Intracellular Localization of Esterase in Rat Liver. Journal of Cell Biology, 2(5):635–??, September 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/5/
635.

Uzman:1957:EMS

[UNG57] Betty Geren Uzman and Genevieve Nogueira-Graf. Electron microscope studies of the formation of nodes of ranvier in mouse sciatic nerves. Journal of Cell Biology, 3 (4):589–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/4/589.

Uzman:1956:FMP

[Uzm56] Betty Geren Uzman. The formation of myelin in the peripheral nerves of vertebrates. Journal of Cell Biology, 2(4):219–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/219.

vanBreemen:1955:SDC

[vBC55] V. L. van Breemen and C. D. Clemente. Silver deposition in the central nervous system and the hematoencephalic barrier studied with the electron microscope. Journal of Cell Biology, 1(2):161–??, March 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/2/
161.

vanBreemen:1956:OBM

[vBRC56] V. L. van Breemen, James F. Reger, and W. Gregory Cooper. Observations on the basement membranes in rat kidney. Journal of Cell Biology, 2(4):283–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
2/4/283.

Vial:1958:ECA

[Via58] Juan D. Vial. The Early Changes in the Axoplasm during Wallerian Degeneration. Journal of Cell Biology, 4 (5):551–??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/5/551.

Wolff:1958:FLN

[WARL58] Sheldon Wolff, K. C. Atwood, M. L. Randolph, and H. E. Luippold. Factors Limiting the Number of Radiation– Induced Chromosome Exchanges. Journal of Cell Biology, 4(4):365–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/4/365.

Watson:1955:NE

[Wat55a] Michael L. Watson. The nuclear envelope. Journal of Cell Biology, 1(3):257–??, May 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/1/3/257.

Watson:1955:UCF

[Wat55b] Michael L. Watson. The Use of Carbon Films to Support Tissue Sections for Electron Microscopy. Journal of Cell Biology, 1(2):183–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/1/2/183.

Watson:1956:CFS

[Wat56] Michael L. Watson. Carbon films and specimen stability. Journal of Cell Biology, 2(4):31–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
31.

Watson:1957:RHA

[Wat57] Michael L. Watson. Reduction of heating artifacts in thin sections examined in the electron microscope. Journal of Cell Biology, 3(6):1017–??, November 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/6/
1017.

Watson:1958:STSa

[Wat58a] Michael L. Watson. Staining of Tissue Sections for Electron Microscopy with Heavy Metals. Journal of Cell Biology, 4(4):475–??, July 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/4/475.

Watson:1958:STSb

[Wat58b] Michael L. Watson. Staining of Tissue Sections for Electron Microscopy with Heavy Metals. Journal of Cell Biology, 4(6):727–??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/4/6/727.

Watson:1959:FON

[Wat59] Michael L. Watson. Further Observations on the Nuclear Envelope of the Animal Cell. Journal of Cell Biology, 6(2):147–??, October 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/6/2/147.

Woodin:1955:PIS

[WB55] A. M. Woodin and S. A. Boruchoff. Particle interaction in solutions derived from ox vitreous humor. Journal of Cell Biology, 1(6):489–??, November 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/6/
489.

Wolken:1957:PS

[WCT57] J. J. Wolken, J. Capenos, and A. Turano. Photoreceptor structures. Journal of Cell Biology, 3(3):441–??, May 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/3/3/441.

Weiss:1957:SSS

[Wei57] Leon Weiss. A study of the structure of splenic sinuses in man and in the albino rat with the light microscope and the electron microscope. Journal of Cell Biology, 3 (4):599–??, July 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/4/599.

Weiner:1959:NMG

[Wei59] Stanley Weiner. A New Method of Glass Knife Preparation for Thin–Section Microtomy. Journal of Cell Biology, 5(1):175–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/1/175.

Went:1959:SIS

[Wen59a] Hans A. Went. Some Immunochemical Studies on the Mitotic Apparatus of the Sea Urchin. Journal of Cell Biology, 5(2):353–??, March 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/2/353.

Went:1959:SMA

[Wen59b] Hans A. Went. Studies on the mitotic apparatus of the sea urchin by means of antigen–antibody reactions in agar. Journal of Cell Biology, 6(3):447–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/6/3/447.

Weiss:1956:BLA

[WF56] Paul Weiss and Wayne Ferris. The basement lamella of amphibian skin. Journal of Cell Biology, 2(4):275–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/2/4/275.

Woessner:1957:CB

[WG57] J. Frederick Woessner and Bernard S. Gould. Collagen biosynthesis. Journal of Cell Biology, 3(5):685– ??, September 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/5/685.

Wanko:1959:EMS

[WG59] Theodor Wanko and Mary Ann Gavin. Electron Microscope Study of Lens Fibers. Journal of Cell Biology, 6 (1):97–??, August 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/1/97.

Wood:1959:UFG

[WH59] Richard L. Wood and Charles C. Howard. Use of Fine Grain Positive Sheet Film for Electron Microscopy. Journal of Cell Biology, 5(1):181–??, January 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/1/

Wischnitzer:1957:UYP

[Wis57] Saul Wischnitzer. The Ultrastructure of Yolk Platelets of Amphibian Oocytes. Journal of Cell Biology, 3(6): 1040–??, November 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/6/1040.

Williams:1955:IEM

[WK55] Robley C. Williams and Frances Kallman. Interpretations of electron micrographs of single and serial sections. Journal of Cell Biology, 1(4):301–??, July 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/4/301.

Wassermann:1956:OFC

[WK56] F. Wassermann and L. Kubota. Observations on fibrillogenesis in the connective tissue of the chick embryo with the aid of Silver impregnation. Journal of Cell Biology, 2(4):67–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/4/67.

Wislocki:1955:DBO

[WL55] George B. Wislocki and Aaron J. Ladman. The demonstration of a blood-ocular barrier in the albino rat by means of the intravitam deposition of Silver. Journal of Cell Biology, 1(6):501–??, November 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/6/
501.

Wassermann:1956:ECE

[WL56] F. Wassermann and Arthur Lindenbaum. Experiments concerning the effect of enzymes on the reconstitution of collagenous fibrils in vitro. Journal of Cell Biology, 2(4):299–??, July 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/4/299.

Williams:1959:NCT

[WLP59] William J. Williams, Edna Larson, and Theodore L. Phillips. A Neoplastic Connective Tissue Mast Cell Capable of Continuous Growth In Tissue Culture. Journal of Cell Biology, 6(3):361–??, December 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/3/
361.

Wachstein:1955:DHD

[WM55] M. Wachstein and E. Meisel. The distribution of histochemically demonstrable succinic dehydrogenase and of mitochondria in tongue and skeletal muscles. Journal of Cell Biology, 1(6):483–??, November 1955. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/1/6/
483.

Witter:1958:ESD

[WM58] Robert F. Witter and William Mink. Effect of Synthetic Detergents on the Swelling and the ATPase of Mitochondria Isolated from Rat Liver. Journal of Cell Biology, 4 (1):73–??, January 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/1/73.

Wachstein:1959:HDS

[WM59] Max Wachstein and Elizabeth Meisel. The Histochemical Demonstration of Secretory Capillaries in the Pancreas with the Aid of Substrate Specific Phosphatases. Journal of Cell Biology, 6(1):119–??, August 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/1/
119.

Whaley:1959:ERG

[WMK59a] W. Gordon Whaley, Hilton H. Mollenhauer, and Joyce E. Kephart. The Endoplasmic Reticulum and the Golgi Structures in Maize Root Cells. Journal of Cell Biology, 5(3):501–??, May 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/3/501.

Wilson:1959:SCM

[WMK59b] G. B. Wilson, J. H. Morrison, and N. Knobloch. Studies on the Control of Mitotic Activity in Excised Roots.

Journal of Cell Biology, 5(3):411–??, May 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/5/3/
411.

Wolken:1958:RSM

[Wol58] Jerome J. Wolken. Retinal Structure. Mollusc Cephalopods: Octopus, Sepia. Journal of Cell Biology, 4(6):835– ??, November 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/6/835.

Woodard:1956:DGE

[Woo56] John W. Woodard. DNA in gametogenesis and embryogeny in Tradescantia. Journal of Cell Biology, 2 (6):765–??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/2/6/765.

Woods:1957:CSH

[Woo57] Philip S. Woods. A chromatographic study of hydrolysis in the Feulgen nucleal reaction. Journal of Cell Biology, 3(1):71–??, January 1957. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/3/1/71.

Woodard:1958:IAN

[Woo58] John W. Woodard. Intracellular amounts of nucleic acids and protein during pollen grain growth in Tradescantia. Journal of Cell Biology, 4(4):383–??, July 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/4/
383.

Wood:1959:IAE

[Woo59] R. L. Wood. Intercellular Attachment in the Epithelium of Hydra As Revealed by Electron Microscopy. Journal of Cell Biology, 6(3):343–??, December 1959. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/6/3/

Westfall:1958:ARA

[WPE^{+}58] B. B. Westfall, E. V. Peppers, V. J. Evans, K. K. Sanford, N. M. Hawkins, M. C. Fioramonti, H. A. Kerr, G. L. Hobbs, and W. R. Earle. The Arginase and Rhodanese Activities of Certain Cell Strains after Long Cultivation in Vitro. Journal of Cell Biology, 4(5):567– ??, September 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/5/567.

Watson:1956:CSM

[WS56a] Michael L. Watson and Philip Siekevitz. Cytochemical studies of mitochondria. Journal of Cell Biology, 2(6): 639–??, November 1956. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/2/6/639.

Watson:1956:IAM

[WS56b] Michael L. Watson and Philip Siekevitz. The isolation and analysis of a mitochondrial membrane fraction. Journal of Cell Biology, 2(4):379–??, July 1956. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/2/4/
379.

Watabe:1958:SSF

[WSW58] Norimitsu Watabe, D. Gordon Sharp, and Karl M. Wilbur. Studies on Shell Formation. Journal of Cell Biology, 4(3):281–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/4/3/281.

Witter:1955:MAA

[WWC55] Robert F. Witter, Michael L. Watson, and Mary A. Cottone. Morphology and ATP-ase of isolated mitochondria. Journal of Cell Biology, 1(2):127–??, March 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
1/2/127.

Wilkins:1959:AHB

[WZ59] M. H. F. Wilkins and G. Zubay. The Absence of Histone in the Bacterium Escherichia coli. Journal of Cell Biology, 5(1):55–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/5/1/55.

Yamada:1955:FSG

[Yam55a] Eichi Yamada. The fine structure of the gall bladder epithelium of the mouse. Journal of Cell Biology, 1(5): 445–??, September 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/5/445.

Yamada:1955:FSR

[Yam55b] Eichi Yamada. The fine structure of the renal glomerulus of the mouse. Journal of Cell Biology, 1(6):551– ??, November 1955. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/1/6/551.

Yamada:1958:PLB

[Yam58] Eichi Yamada. A Peculiar Lamellated Body Observed in the Cells of the Pigment Epithelium of the Retina of the Bat, Pipistrellus abramus. Journal of Cell Biology, 4(3):329–??, May 1958. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/4/3/329.

Yamada:1959:CBF

[Yam59] Eichi Yamada. A crystalline body found in the rod inner segment of the frog’s eye. Journal of Cell Biology, 6(3): 517–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/6/3/517.

Young:1959:CMSa

[YFJ59a] I. Elizabeth Young and Philip C. Fitz-James. Chemical and morphological studies of bacterial spore formation. Journal of Cell Biology, 6(3):467–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/3/467.

Young:1959:CMSb

[YFJ59b] I. Elizabeth Young and Philip C. Fitz-James. Chemical and morphological studies of bacterial spore formation. Journal of Cell Biology, 6(3):483–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/3/483.

Young:1959:CMSc

[YFJ59c] I. Elizabeth Young and Philip C. Fitz-James. Chemical and morphological studies of bacterial spore formation. Journal of Cell Biology, 6(3):499–??, December 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
6/3/499.

Yasuzumi:1957:SAR

[YI57] G. Yasuzumi and Hiroshi Ishida. Spermatogenesis in animals as revealed by electron microscopy. Journal of Cell Biology, 3(5):663–??, September 1957. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/3/5/
663.

Yasuzumi:1958:SAR

[YT58] G. Yasuzumi and Hiroaki Tanaka. Spermatogenesis in Animals as Revealed by Electron Microscopy. Journal of Cell Biology, 4(5):621–??, September 1958. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/4/5/
621.

Zubay:1959:AHB

[ZW59] G. Zubay and M. R. Watson. The Absence of Histone in the Bacterium Escherichia coli. Journal of Cell Biology, 5(1):51–??, January 1959. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/5/1/51.